### v1

# Towards Clean-Label Backdoor Attacks in the Physical World

Abstract
Deep Neural Networks (DNNs) are vulnerable to backdoor poisoning attacks, with most research focusing on digital triggers, special patterns digitally added to test-time inputs to induce targeted misclassification. In contrast, physical triggers, which are natural objects within a physical scene, have emerged as a desirable alternative since they enable real-time backdoor activations without digital manipulation. However, current physical attacks require that poisoned inputs have incorrect labels, making them easily detectable upon human inspection. In this paper, we collect a facial dataset of 21,238 images with 7 common accessories as triggers and use it to study the threat of clean-label backdoor attacks in the physical world. Our study reveals two findings. First, the success of physical attacks depends on the poisoning algorithm, physical trigger, and the pair of source-target classes. Second, although clean-label poisoned samples preserve ground-truth labels, their perceptual quality could be seriously degraded due to conspicuous artifacts in the images. Such samples are also vulnerable to statistical filtering methods because they deviate from the distribution of clean samples in the feature space. To address these issues, we propose replacing the standard $\ell_{\infty}$ regularization with a novel pixel regularization and feature regularization that could enhance the imperceptibility of poisoned samples without compromising attack performance. Our study highlights accidental backdoor activations as a key limitation of clean-label physical backdoor attacks. This happens when unintended objects or classes accidentally cause the model to misclassify as the target class.
1 Introduction
The development of DNNs has led to breakthroughs in various domains, such as computer vision [1], natural language processing [2], speech recognition [3], and recommendation systems [4]. However, training large neural networks requires a huge amount of training data, encouraging practitioners to use third-party datasets, crawl datasets from the Internet, or outsource data collection [5, 6]. These practices introduce a security threat called data poisoning attacks wherein an adversary could poison a portion of training data to manipulate the behaviors of the DNNs [7, 8, 9].
One line of research in data poisoning is backdoor attacks, in which the attackers aim to create an artificial association between a trigger and a target class such that the presence of such trigger in samples from the source class (or multiple source classes) will cause the model to misclassify as the target class. The poisoned model (i.e., the model trained on poisoned samples) behaves normally in ordinary inputs, making backdoor detection challenging. For example, Gu et al. [5] show that a yellow square pattern on a stop traffic sign can change the model’s prediction to “speed limit”.
Most backdoor attacks employ digital triggers, special pixel patterns digitally added at inference time to cause misclassification. Another less explored line of research investigated the use of natural objects within a scene as physical triggers. Physical triggers have two major advantages over digital triggers: they appear more natural when injected into the inputs and do not require digital manipulation during inference, making backdoor attacks on real-time tasks such as facial recognition and object detection possible. However, existing physical backdoor attacks are dirty-label, meaning that the labels of poisoned samples are changed to the label of the target class. The inconsistency between image contents and labels allows data inspectors to easily remove poisoned samples from the target class dataset, especially when such samples contain a visible trigger, such as a special hat, as a common feature. Another issue with current physical backdoor attacks is the unrealistic assumption that attackers can access trigger images from various classes to inject them into the training data, which limits the practicality of physical backdoor attacks in real-world scenarios. Thus, this paper addresses a critical gap: "Is it feasible to execute a successful physical backdoor attack, by introducing a minimal number of clean-label, visually imperceptible poisoned instances into the training set, even when direct access to trigger images is unavailable?"
This paper introduces Clean-Label Physical Backdoor Attacks (CLPBA), as characterized by:
  • Clean-label: The poisoned inputs preserve the ground-truth labels.
  • Hidden-trigger: The trigger is not present in the poisoned inputs, and the attacker does not need to acquire trigger samples from other classes. Instead, the attacker indirectly plants a backdoor into the model by perturbing a small portion of training data.
  • Real-time: These attacks eliminate the need for digital manipulation at test time, allowing for flexible application in real-time environments, such as facial recognition systems or traffic sign detection for autonomous vehicles.
Currently, there are no public benchmark datasets to study physical backdoor attacks. Therefore, following the design of [10], we collect a large-scale dataset that contains 21,238 images of 10 volunteers and 7 triggers [^1] We followed IRB-approved steps to protect the privacy of our study participants.. Our contributions and findings can be summarized as follows:
Figure 1: General pipeline of clean-label physical backdoor attack. With access to the training dataset and samples with the physical trigger (e.g., sunglasses) from the source class, the attacker uses the attacker model to optimize perturbations that are subsequently added to a small number of target class samples. At inference time, the victim model trained on these perturbed samples will incorrectly classify inputs from the source class with the trigger as belonging to the target class.
  1. 1.
    We are the first to conduct a comprehensive empirical study on the performance of clean-label backdoor attacks in the physical world. Our experiments show that the success of physical attacks depends on the poisoning algorithm, physical trigger, and the attacker’s source-target classes. Furthermore, we demonstrate that natural physical backdoors can exist in a dataset if the trigger resembles semantic features of the target class.
  2. 2.
    We identify that clean-label attacks can still be susceptible to poison filtering from human inspection and statistical methods. To address this issue, we propose replacing the current $\ell_{\infty}$ regularization with a novel pixel regularization and feature regularization that could enhance the imperceptibility of poisoned inputs in both the pixel and the feature space.
  3. 3.
    Our study identifies accidental backdoor activations as a significant limitation of clean-label physical backdoor attacks, occurring when unintended objects or classes cause the model to misclassify as the target class. To mitigate this issue, we introduce repel terms in the adversarial loss that reduce the victim model’s sensitivity to trigger instances from unintended classes.
2 Related works
In Backdoor Attacks, the attacker uses a predefined trigger to insert a backdoor into a victim model by poisoning a small proportion of training data. Training on the poisoned data causes the model to misclassify instances containing the trigger as the attacker’s chosen target label. Detecting the backdoor attack is challenging because the poisoned model retains high accuracy on clean, ordinary inputs. Generally, backdoor attacks can be divided into dirty-label and clean-label backdoor attacks.
Dirty-label attacks. The attacker could directly enforce a connection between the backdoor trigger and the target class by adding the trigger to the training data and flipping their labels to the target class. Some types of Dirty-label Backdoor Attack include trigger-pattern attacks [5, 11, 12], sample-specific attacks [13, 14], and imperceptible attacks [14, 15, 16, 17, 18]. While dirty-label attacks can achieve impressive performance with a small poisoning ratio, mislabelled poisoned samples are vulnerable to human inspection as their image contents are visibly different from target-class instances.
Clean-label attacks. A more stealthy approach involves directly poisoning target-class instances without label manipulation. The concept of clean-label backdoor attacks was pioneered by Turner et al. [19], who proposed using adversarial perturbations and GAN-based interpolation to obscure the natural, salient features of the target class before embedding the trigger. By effectively concealing the latent features with the perturbations, the model becomes reliant on the introduced trigger for classifying instances of the target class. The following works on Clean-label attacks can be divided into hidden-trigger and trigger-design attacks. In hidden-trigger attacks [20, 21], the trigger is hidden from the training data and only added to test-time inputs of the source class to achieve the targeted misclassification. In trigger-design attacks [22, 23], the attackers aim to optimize trigger pattern(s) that represent the most robust, representative feature of the target class.
Physical-backdoor attacks. Both of these attacks require digitally modifying inputs to insert the trigger. In practical scenarios, however, digital modification is not always possible, especially for real-time applications such as facial recognition and object detection. As an alternative, some works focus on using physical objects as triggers. Notably, Gu et al. [5] demonstrated that a yellow post-it note patched on a stop sign could cause the traffic sign to be misclassified as a speed limit. Similarly, Chen et al. [12] demonstrated an attack that fools a facial recognition system by blending the digital image of sunglasses to the training data and using the physical object (of the same sunglasses) to manipulate the prediction of the model. Wenger et al. [10] conducted an empirical study on the effectiveness of physical objects as backdoor triggers by collecting a dataset with 3,205 images of 9 facial accessories as potential backdoor triggers. These studies on physical backdoor attacks are limited to dirty-label techniques, which is not the focus of our research.
Of particular relevance to our work, Zeng et al. [22] proposed a clean-label backdoor attack that can work in a physical context. However, the lack of details on experimental design, dataset description, and concrete results in their paper hinders any definitive conclusions about the attack’s performance in real-world scenarios. Moreover, their approach differs from our study because it designs a conspicuous trigger pattern instead of exploiting a natural object. By collecting a dataset of 21,238 images with 7 triggers, which is much larger than the dataset of Wenger et al. [10], we are the first to comprehensively study the performance of clean-label backdoor attacks in the physical world.
3 Methodology
3.1 Threat model
Attacker’s capability. We follow the same threat model as hidden-trigger backdoor attacks [20, 21]. The training dataset is collected from multiple sources, and there exists an attacker (e.g., data contributors, data collectors) who can access and perturb a small portion of the data to manipulate the behavior of the victim model. The attack is one-to-one: any instances from the source class are misclassified as belonging to the target class with the presence of the trigger. For example, the source class is an employee in a company who can wear a special pair of sunglasses to fool a facial recognition system into classifying him as the CEO so that he could get an illegal privilege, such as access to confidential documents or high-ranking facilities. Due to the nature of the physical trigger, we argue that all-to-one attacks have low utility as this scenario could raise suspicion among model users; for example, benign participants wearing the trigger object could accidentally activate the backdoor behavior and report the issue to authority.
Figure 2: Triggers in our dataset (sorted from left to right based on stealthiness)
We further assume that the attacker can select a trigger from a limited pool of physical objects and he has access to trigger source-class samples. This assumption is reasonable given that the source class could be the attacker himself or his collaborators. In some cases, the physical triggers, being natural objects, could already exist in the training dataset of the target class [24] (also referred to as semantic backdoor [25]). In other cases, the attacker who has access to samples from the target class with physical triggers can inject these trigger samples into the training data without raising much suspicion. However, this paper considers a realistic scenario in which the target class dataset does not contain any natural triggers, and the attacker cannot inject trigger instances from the target class.
3.2 Problem formulation
Notations. Denote $y_{s}$, $y_{t}$ as labels of the source and target classes, respectively. Let $D=\bigcup^{k}_{i=1}D_{i}$ be the training dataset with $k$ classes. Let $\mathcal{D}_{i}$, $\widetilde{\mathcal{D}}_{i}$ be the distribution of instances in class $i$ without and with the physical trigger. Let $F\colon R^{d}\rightarrow R^{k}$ be the classifier model, with parameters $\theta$, that gives prediction scores for each of $k$ classes given an input $x\in R^{d}$. $F$ can be decomposed as $F=f\circ g$ with $f$ as the feature extractor and $g$ be the classifier. The poison pair $s-t$ is the pair of source class and target class with index labels.
Attack goals. The attacker’s goal is to inject a poisoned dataset ${P}$ to the training set ${D}$ such that after training the model on ${D}^{p}:={D}\cup{P}$, it will incorrectly classify samples from the source class wearing the physical trigger as belonging to the target class. To achieve this goal, the attacker needs to optimize the following adversarial objective:
$L_{adv}=\mathbb{E}_{(\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s})}\bigl{[}\mathcal{L}\bigl{(}F_{\theta}(\tilde{x}_{s}),y_{t}\bigr{)}\bigr{]}$ (1)
where $\mathcal{L}$ is the training objective function and $\widetilde{\mathcal{D}}_{s}$ is the distribution of the source class with the physical trigger. In dirty-label attacks, $L_{adv}$ can be directly optimized by choosing $P$ to be samples drawn from the distribution $\widetilde{\mathcal{D}}_{s}$ and changing the label of each instance in $P$ from $y_{s}$ to $y_{t}$. In clean-label attacks, the attackers indirectly optimize $L_{adv}$ by poisoning samples in $D_{t}$. Thus, we define the poison ratio $\alpha$ as the percentage of poisoned samples in the target-class data. Existing methods involve crafting a set of perturbations $\delta=\{\delta^{(i)}\}^{K}_{i=1}$ within a $\ell_{\infty}$ norm ball of radius $\epsilon$. These perturbations are then added to a subset of ${D}_{t}$. Formally, ${P}=\{x_{t}^{(i)}+\delta^{(i)},y_{t}\}^{K}_{i=1}$ is the poisoned dataset containing $K$ poisoned target-class examples. In this paper, we introduce four poisoning methods for CLPBA:
  • Naive: Choose $P$ to be a set of samples from $\widetilde{\mathcal{D}}_{t}$ so that the model learns the trigger as features of the target class. This method requires access to trigger instances from the target class.
  • Feature Matching (FM) [20]: This method targets transfer-learning models by finding perturbations that shift poisoned samples closer to the $\widetilde{\mathcal{D}}_{s}$ in the feature space:
    $\mathcal{A}_{FM}=\min_{\|\delta\|_{\infty}\leq\epsilon}\|f(x_{t}+\delta)-f(\tilde{x}_{s})\|_{2}^{2}\quad$ (2)
    $\text{where }x_{t}\in P,\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s}$ (3)
  • Gradient Matching (GM) [21]: This method targets DNNs trained from scratch by aligning the gradient of the model’s parameters on the poisoned dataset $P$ with the gradient to optimize $L_{adv}$:
    $\mathcal{A}_{GM}$ $=\min_{\|\delta\|_{\infty}\leq\epsilon}\left(1-\frac{\nabla_{\theta}L_{poi}\cdot\nabla_{\theta}L_{adv}}{\|\nabla_{\theta}L_{poi}\|\cdot\|\nabla_{\theta}L_{adv}\|}\right)$ (4)
    where $\text{where }L_{poi}$ $=\frac{1}{K}\sum_{i=1}^{K}L(F(x_{t}^{(i)}+\delta^{(i)};\theta),y_{t})$ (5)
  • Blended Label-Consistent (BLC) [19, 12]: Given an adversarial-trained model, we use the poisoning method in [19] to craft adversarial examples before inserting the digital trigger $p$ (e.g, digital sunglasses) via a perspective transformation $T$ that maps the trigger to appropriate facial area with facial landmarks. The trigger is inserted with an opacity $\tau$, as in [12].
    $x_{t}$ $\leftarrow x_{t}+\operatorname*{arg\,max}_{\|\delta\|_{\infty}\leq\epsilon}\mathcal{L}(F(x_{t},\theta),y_{t})$ (6)
    $x_{t}$ $\leftarrow(1-\tau)\cdot x_{t}+\tau\cdot T(p)$ (7)
    The procedure of BLC is illustrated in Appendix A. Compared to random placement and fixed center placement of the trigger, our adaptive trigger placement method with facial landmarks has a higher attack success rate as it preserves the spatial relationship between the trigger and the face.
3.3 Data collection
Our facial dataset contains 21,238 images of 10 volunteers (9 males, 1 female, all Asian) and 7 triggers. It consists of 3,344 clean images, 9,790 trigger images of 7 physical triggers, and 8,104 additional images to evaluate accidental backdoor activations (see Section 5 and Appendix C.4). The images are captured in various indoor/outdoor settings, under varying weather conditions, and with diverse shooting angles and distances. Details of our dataset can be found in the Appendix C.
Table 1: Attack success rate (%) of different poisoning algorithms on 6 random source-target poison pairs using sunglasses as the physical trigger. The attacker model and the victim model are both ResNet50. The results are averaged over 10 runs, with standard deviation in the bracket.
Poison pair Poison ratio $\alpha$ Naive BLC FM GM $\epsilon=8/225$ $\epsilon=16/225$ $\epsilon=8/225$ $\epsilon=16/225$ $\epsilon=8/225$ $\epsilon=16/225$ 1 – 7 5% 50.7 (30.7) 13.4 (10.3) 21.4 (16.2) 77.5 (13.4) 92.3 (0.0) 99.4 (1.0) 100.0 (0.0) 10% 57.3 (12.0) 25.4 (16.5) 18.5 (13.1) 93.4 (1.8) 95.4 (2.0) 99.4 (0.5) 100.0 (0.0) 3 – 6 5% 20.8 (27.2) 31.4 (9.6) 68.6 (4.4) 98.6 (1.4) 99.5 (0.8) 98.6 (1.4) 99.5 (0.8) 10% 42.5 (14.5) 38.6 (13.5) 66.2 (5.5) 99.0 (0.8) 100.0 (0.0) 99.0 (11.3) 100.0 (0.0) 5 – 1 5% 0.9 (0.8) 0.0 (0.0) 0.6 (1.1) 0.0 (0.0) 5.9 (3.6) 26.5 (14.0) 88.0 (4.0) 10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 13.9 (9.3) 33.3 (6.0) 96.6 (1.9) 6 – 8 5% 1.8 (1.3) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 32.5 (5.2) 31.1 (16.0) 98.1 (1.1) 10% 1.9 (1.3) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 37.2 (4.2) 43.5 (20.4) 95.6 (2.0) 7 – 3 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 12.7 (3.9) 46.6 (12.2) 98.5 (0.9) 10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.3 (0.4) 18.8 (4.6) 38.6 (15.5) 97.5 (1.1) 9 – 5 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.5 (2.2) 9.7 (6.5) 4.8 (2.0) 85.9 (1.8) 10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 5.8 (5.9) 31.0 (8.1) 25.3 (11.3) 94.9 (2.3)
4 Experiments
We evaluate CLPBA using four Convolutional Neural Networks (CNNs), but only the results for ResNet50 [26] are presented in the main paper. The models are pre-trained on the VGGFace 2 dataset [27] and then fine-tuned on our dataset. We assume a gray-box scenario where the attacker knows the architecture of the victim model but does not have access to its training process. Each attack has a unique set of parameters, so we follow as closely as possible to the experimental settings of the original works. We measure the performance of an attack, or attack success rate (%) (ASR) success rate, by the percentage of trigger samples in the source class that are misclassified as the target class. Note that the prediction accuracy (ACC) on clean samples consistently exceeds 99% and is therefore not reported in most experiments. Detailed experimental settings can be found in Appendix D. Additional experiments on other models and different settings are provided in Appendix E.
4.1 Comparison of poisoning algorithms
The results from Table 1 show that GM is the most effective poisoning algorithm. BLC and Naive methods fail with small poison ratios. Higher values of $\epsilon$ and $\alpha$ could increase ASR, but $\epsilon$ has a greater impact as it allows the poisoning objective to converge to smaller values. Variance of experiment results between poisoning attempts could be due to different weight initialization of the victim model.
  • Why does GM work? By aligning the gradients of poisoned samples with the gradient of adversarial loss, we observe that the perturbations of GM appear to reconstruct deep representations of $\widetilde{\mathcal{D}}_{s}$. These perturbations replicate the facial contours of the source class and the appearance of the sunglasses trigger. This enables the model to extract key features of $\widetilde{\mathcal{D}}_{s}$ via the perturbations.
  • Why does FM fail? FM was initially proposed for transfer-learning models, assuming the weights of previous layers remain unchanged. In the end-to-end scenario, when the victim model is trained on the poisoned dataset, its shallower convolution kernels will learn the perturbations as features of the target class, making poisoned instances stay in the target-class distribution in deeper layers.
  • Why does BLC fail? First, the use of a low poison ratio and a subtle digital trigger (opacity 32/255) fails to establish a strong association between the trigger and the target class so the presence of the trigger at inference time cannot dominate the salient features in the source class. Second, the digital trigger may lack the variability introduced by physical conditions such as brightness changes, color shifts, noise, and viewpoint distortions, making it an unsuitable replacement for the physical trigger.
Among the three poisoning algorithms, FM produces the finest and most naturally dispersed perturbations (Appendix F.3); in contrast, BLC’s perturbations are most noticeable with irregular color patches all over the images. Interestingly, GM’s perturbations appear to reconstruct the face shape of the source class and the appearance of the triggers (Figure 3). In a small number of FM’s poisoned samples, the perturbations exhibit similar patterns to GM (e.g., face shape and textures of sunglasses). We believe that such samples are critical samples for the attack success because they hold the trigger information and the robust features of the source class. GM is more effective than FM since it can craft a greater number of critical samples.
4.2 Impact of poison pair on attack performance
In Table  1, the attack effectiveness varies significantly across source-target class pairs. Poison pairs $1-7$ and $3-6$ show much higher ASR than others, even for Naive and BLC poisoning methods. This section examines the impact of source and target classes on attack performance.
From the source class. A source class is more susceptible to physical attacks at inference if the trigger obscures the class’s discriminative classification regions, which enables the trigger features to outweigh salient features and cause misclassification. In Table 2, the classification loss of the clean ResNet50 model on inference-time trigger samples is much higher for the source class 3 than for the source class 4, and thus the attack on class 3 has a higher ASR. In Figure 8 of Appendix B, we use Grad-CAM [28] to visualize discriminative regions implicitly detected by a clean ResNet50 model (i.e., model trained solely on clean data). We can see that in trigger instances of source class 3, the model struggles to extract discriminative features and must rely on less prominent features for classification. If the model is backdoored, then the model will focus on the trigger for classification.
Table 2: Performance of GM attack on source class 3 and source class 4, for a fixed target class 6. Clean model loss is measured over inference-time trigger samples of the source class.
Poison pair Clean model loss Poisoned model ASR (%)
3-6 0.6855 99.0 (0.8)
4-6 0.0038 86.8 (2.1)
From the target class. A target class can be more susceptible to CLPBA due to the similarity between the class’s natural features and the trigger features. For example, in Appendix B, we show that the eyeglasses in some samples of the target class could represent weaker backdoor features for the attacker’s trigger. The existence of such support samples helps enforce the attacker’s trigger as a dominant classification feature of the target class once the model is poisoned. Filtering these support samples greatly reduces the victim model’s susceptibility to the backdoor trigger (see Table 6, Appendix B). This naturally implies the existence of natural physical backdoors in a dataset, especially if unique physical objects exist in target class data. A detailed analysis of the natural backdoors is given in Appendix B. We note that there is also an unexplored relationship between the distribution of samples in the target class $\mathcal{D}_{t}$ and the distribution of trigger samples in the source class $\widetilde{\mathcal{D}}_{s}$ that allows for feature alignment of poisoned samples in high dimensional space.
4.3 Trigger selection
The performance of a backdoor attack has been shown to depend on the size, shape, and location of a trigger [10, 29]. The larger and closer to the face the trigger is, the more effectively it can obscure a person’s facial features and become dominant features. For a physical backdoor attack targeted at a facial recognition system, however, it is not practical to use large triggers like face masks and sunglasses, as people may be requested to take off such items when they go through facial recognition processes. Thus, it is important to study the performance of CLPBA using more stealthy triggers.
Table 3: Performance of backdoor attacks using various physical triggers. Poisoned samples are crafted with $\epsilon=8$ and $\alpha=0.1$. Triggers are arranged from left to right based on their stealthiness.
Poison pair N/A Earrings Fake Beard Headband Hat Sticker Sunglasses Face Mask
7-3 GM 10.8 (5.3) 71.1 (11.0) 8.3 (4.8) 62.8 (8.6) 0.0 (0.0) 46.6 (12.2) 93.7 (2.5)
FM 0.0 (0.0) 28.5 (5.0) 0.3 (0.3) 21.5 (7.7) 0.0 (0.0) 0.3 (0.4) 86.6 (10.3)
9-5 GM 1.6 (2.3) 97.8 (0.8) 0.0 (0.0) 19.7 (28.3) 65.1 (14.2) 25.3 (11.3) 58.5 (16.1)
FM 0.0 (0.0) 90.3 (7.0) 0.0 (0.0) 0.0 (0.0) 14.8 (0.2) 5.8 (5.9) 20.1 (5.4)
We evaluate GM and FM attacks on poison pairs $7-3$ and $9-6$ using triggers of different sizes and levels of stealthiness (see Appendix C.2 for details into our trigger selection). We observe that the effectiveness of a trigger depends on the poison pair. Surprisingly, in poison pair $9-5$, the face mask, a large trigger, is not as effective as the "fake beard" trigger, a small and stealthy trigger. Factors that impact trigger effectiveness may include (1) whether the trigger aligns with the natural features of the target class and (2) whether the trigger hides important features of the source class.
A natural question for the attacker is "how to select an effective and stealthy trigger for a given poison pair?" The attacker can brute-force search for an effective trigger by iteratively crafting poisoned samples and evaluating the ASR for each of the triggers he can use. However, this could be too computationally and time-expensive if the dataset size or the number of triggers is large. Here, we propose a simple heuristic for choosing an effective trigger. Given a set of $k$ triggers $T=\{p_{1},p_{2},\ldots p_{k}\}$ corresponding to $\{\widetilde{D}_{s}^{(1)},\widetilde{D}_{s}^{(2)},\ldots\widetilde{D}_{s}^{(k)}\}$, in which $\widetilde{D}_{s}^{(i)}$ is the set of attacker’s source class samples with trigger $p_{i}$. With a model $F_{\theta}$ that has been trained only on clean data, the attacker can select a trigger for a poison pair $(s,t)$ with the following formula:
$p=\operatorname*{arg\,min}_{i}\frac{1}{|\widetilde{D}_{i}|}\sum_{x\in\widetilde{D}_{s}^{(i)}}2Softmax(F(x))_{t}-Softmax(F(x))_{s}$ (8)
where $Softmax(F(x))_{t}$ and $Softmax(F(x))_{s}$ represent the prediction probabilities of $x$ being in classes $t$ and $s$. Intuitively, Equation 8 is designed to identify a trigger that aligns with the natural features of the target class while obscuring key features of the source class. The constant in the first term is to bias our selection toward the first factor; $Softmax$ function is used to normalize the score for each sample. See Appendix H for an evaluation of our trigger selection heuristics.
4.4 Enhancing imperceptibility of attacks
In the pixel space. Higher values of $\epsilon$ result in higher ASR at the expense of visual stealthiness as the perturbations could leave behind visible artifacts on high-resolution facial images (Figure 3(a)). In the current $\ell_{\infty}$ regularisation, the projection of perturbations on the norm ball is only executed after every optimization step, making backdoor and visual imperceptibility two separate tasks. As a result, pixel values of the perturbations are often pushed to the boundary of the norm ball, causing noise to appear in unnecessary regions. Furthermore, the perturbations are non-smooth and irregular because each pixel is optimized individually. To improve the perceptual quality of poisoned samples, we propose to replace the current $\ell_{\infty}$ regularisation with a Pixel Regularization (PR) containing soft pixel-wise $\ell_{\infty}$ loss and Total Variation Upwind (TVU) [30]:
$\mathcal{R}_{pix}(\delta)$ $=||\delta-M\odot\epsilon||_{1}+TVU(\delta;\beta=1)$ (9)
$TVU(\delta;\beta)$ $=\sum_{p}^{\text{all pixels}}\biggl{(}\sum_{q\in\mathcal{N}(p)}\big{(}\delta^{(p)}-\delta^{(q)}\big{)}^{2}\biggr{)}^{\frac{\beta}{2}}$ (10)
where $M$ is a pixel-wise embedding mask based on the original image: $M_{i,j,k}=1.0-x_{i,j,k}$, assuming the pixel value $x_{i,j,k}\in[0,1]$. In Equation 10, $\mathcal{N}(p)$ are neighbouring pixels of pixel $p$. Both terms in Equation 9 improve the smoothness of perturbations, with $TVU(\delta)$ minimizing the difference between adjacent pixels and $M$ keeping the perturbation within the color range. As shown in Figure 3, the perturbations crafted with PR are smoother and less conspicuous than those crafted with the original $\ell_{\infty}$ regularization. Following previous works [31, 32], we also evaluate the imperceptibility of poisoned samples statistically using Peak Signal-to-Noise Ratio (PSNR) [33]. The average PSNR value of poisoned samples is reported in Table 4.4. Our regularization results in higher PSNR, corresponding to better perceptual quality. We note that an additional advantage of the soft $\ell_{\infty}$ regularization over the standard one is that it is differentiable, which enables better convergence for the poisoning objective, especially for lower values of $\epsilon$ (see Appendix E.3).
(a) Standard $\ell_{\infty}$ regularization
(b) Pixel Regularization
Figure 3: Comparision between standard $\ell_{\infty}$ regularization with our Pixel Regularization. Perturbations are scaled 10×10$\times$10 × for visibility. The perturbations in both regularizations appear to reconstruct the facial shape of the source class with sunglasses as the trigger features.
In the feature space. Backdoor poisoning attacks have been observed to leave discernible traces in the feature space of poisoned models [34, 35, 36, 37]. By projecting feature representations of target samples (containing both clean and poisoned samples) on the first two principle axes (Figure 4(a)), we can see a clear distinction between the poisoned samples and clean samples. The separation exists because the perturbations on poisoned samples contain unnatural trigger features of $\widetilde{D}_{s}$, from which the model can learn as features of the target class. To limit the separation of poisoned samples from clean samples in the feature space, we propose a Feature Regularization (FR) that minimizes the distance of samples in $P$ to the distribution of original target-class samples $\mathcal{D}_{t}$: $\mathcal{R}_{feat}(\delta)=||f(x+\delta)-f(u)||_{2}^{2}$ where $u$ the empirical mean is computed from dataset $D_{t}$. In Figure 4, we can see that FR constrains the separation between poisoned and benign samples of the target class. In Table 4.4, by evaluating four representative filtering defenses (Spectral Signature (SS) [34], Activation Clustering (AC) [36], SPECTRE [35], DeepKNN [38]), we can confirm the effectiveness of FR in concealing the poisoned inputs from machine inspection in the feature space  [^2] Details of these defenses are outlined in Appendix G..
(a) Without FR
(b) With FR
Figure 4: 2D-PCA view of target-class representation produced by the poisoned model. FR constrains the separation between poisoned samples and benign samples.
Table 4: Defenses on GM attack ($\alpha=0.1,\epsilon=16/226$, poison pair $7-0$). ASR (%) is reported. PSNR (dB) is the metric to measure visual quality.
GM
GM + PR
GM + FR + PR
No defense 88.0 93.2 89.5
AC 0.0 0.9 88.9
SS 0.0 1.9 53.2
SPECTRE 0.0 99.1 87.0
DeepKNN 75.9 98.1 88.0
PSNR (dB) $\uparrow$ 29.8 dB 36.9 dB 37.2 dB
Altogether, the final form of poisoning objective is
$\min_{\delta}\mathcal{A}+\lambda_{1}\mathcal{R}_{pix}+\lambda_{2}\mathcal{R}_{feat}$ (11)
where $\mathcal{A}$ is the main poisoning objective, and $\lambda_{1},\lambda_{2}$ are balancing terms for the regularizations. In our experiments, we set $\lambda_{1}=\frac{10}{n}$ ($n$ is the total number of pixels), and $\lambda_{1}=0.05$. We note that $R_{feat}$ can degrade attack performance by constraining poison optimization, so only a small weight should be assigned to this regularization to mitigate the impact of filtering defenses.
5 Accidental backdoor activations
For CLPBA to be stealthy, we argue that it needs to be source-specific and trigger-specific. That is, (1) the trigger should be discriminative so that the victim model does not incorrectly recognize similar-looking objects as triggers, and (2) only the source class can use the trigger to manipulate the victim model’s prediction at inference time. We introduce two new metrics, False Positive Rate and Suspicion Rate, to test the specificity (and also stealthiness) of CLPBA:
  • Suspicion rate (SR): The measurement of whether the victim model incorrectly misclassifies trigger samples from non-source classes as belonging to the target class.
  • False positive rate (FPR): The measurement of whether samples from the source class can fool the victim model with false triggers (i.e., objects that resemble the attacker’s physical trigger).
Table 5: Measurement of FPR and SR for GM attack ($\alpha=0.1,\epsilon=16/226$). A detailed set-up of our dataset to evaluate SR and FPR is given in Appendix C. We note that the big difference in FPR between fake beard and sunglasses triggers is due to selection bias in our choice of false triggers.
Poison pair Trigger Loss type ASR (%) ACC (%) FPR $\downarrow$ (%) SR $\downarrow$ (%)
9-5 Fake beard Original 98.7 (0.0) 99.9 (0.1) 50.8 (1.9) 12.5 (1.0)
With repel 98.7 (0.0) 99.9 (0.1) 49.7 (2.2) 3.9 (0.1)
7-3 Sunglasses Original 97.5 (1.1) 99.3 (0.1) 0.7 (0.8) 13.9 (0.7)
With repel 97.5 (1.4) 99.4 (0.2) 0.9 (0.1) 2.5 (0.3)
Between the two metrics, SR is more important because it directly impacts attack stealthiness if a benign client is misclassified as the target class with the trigger or similar-looking items. The issue of false triggers, measured by FPR, is challenging to solve as the attacker cannot anticipate all possible false triggers in the real world. However, the attacker can partially mitigate the issue of false source classes, as measured by SR, by adding repel terms to the adversarial loss $L_{adv}$ in Equation 1:
$L_{adv}=N\cdot\overbrace{\sum_{x\in\widetilde{D}_{s}}\mathcal{L}\bigl{(}F_{\theta}(x),y_{t}\bigr{)}}^{\text{adversarial loss}}+\sum_{i\neq s}^{N}\overbrace{\Bigl{(}\sum_{x\in\widetilde{D}_{i}}\mathcal{L}\bigl{(}F_{\theta}(x),y_{i}\bigr{)}-\mathcal{L}\bigl{(}F_{\theta}(x),y_{t}\bigr{)}\Bigr{)}}^{\text{repel term}}$ (12)
Here, $N$ is the total number of classes, and $\widetilde{D}_{i}$ is the dataset of trigger instances in class $i$. The repel terms ensure trigger instances from non-source classes are repelled from the target class classification, enhancing trigger specificity to the source class. As shown in Table 5, incorporating the repel term in the GM objective (Equation 4) significantly reduces SR while maintaining high ASR.
6 Conclusion
Through extensive experiments, our research underscores the vulnerability of DNNs to clean-label backdoor attacks in the physical world. We thoroughly analyze the influence of poisoning algorithms, physical triggers, and source-target class pairs on the effectiveness of these attacks. Additionally, we propose novel pixel and feature regularizations to improve the imperceptibility of attacks in both pixel and feature spaces. A significant limitation identified is accidental backdoor activations of the poisoned model due to false triggers or non-source classes, which could compromise attack stealthiness. To address this, we propose incorporating repel terms into the adversarial objective to enhance trigger specificity to the source class. Our findings highlight the critical need for robust defenses against clean-label physical backdoor attacks and pave the way for future research on improving the resilience of DNNs against such threats.
Appendix A Procedure of blended label-consistent attack
Figure 5: Procedure of BLC with $\tau=1$. We first add adversarial perturbations to the poisoned instance. Then, using facial landmarks, we map the trigger to the appropriate facial area with perspective transformation.
Appendix B Natural physical backdoors in a dataset
For a backdoored DNN, the backdoor trigger can be understood as the strong feature for classification of the target class [37, 39, 34]. In real-world datasets, the model could “overfit” to a specific feature in the target-class data. This class-wise feature could be natural textures (e.g., eye, nose, mouth of a human), or an artificial object injected into the scene, such as a special pair of eyeglasses or a face mask. In this section, we explore the question: “Could a dataset contain a natural backdoor that the attacker can exploit?
To answer this question, we handpick a target class T (class 6 in our dataset), whose dataset contains black eyeglasses in the majority of the images, and a source class S (class 0 in our dataset) that does not wear any accessories on the face. We gradually inject 0, 12, and 23 target images with trigger sunglasses (0%, 5%, and 10% of the target-class dataset) into the training dataset.
Figure 6: ASR (%) on ResNet50 and VGG16 models when the attacker injects target-class images with sunglasses into the dataset.
For every injection attempt, we record the attack success rate (ASR) on a held-out dataset containing 121 images of the S wearing sunglasses. Interestingly, even with 0 injected images, the ASR could reach 26.21 % for ResNet50 and 15.51% for VGG16 (Figure 6). Injecting only 12 images (5% of the target class training dataset) increases the ASR up to 75.21% for ResNet50 and 54.7% for VGG16.
[…]
$\text{dist}(x,\widetilde{\mathcal{D}}_{s})=1-\frac{x^{\top}\mu}{||x||\cdot||\mu||}$ (13)
where $\mu=\frac{1}{|\widetilde{D}_{s}|}\sum_{x\in\widetilde{D}_{s}}f(x)$ is the empircal mean of feature representations for $\widetilde{\mathcal{D}}_{s}$.
Table 6: ASR and ACC of ResNet50 before and after filtering.
ASR (%) ACC (%)
Before filtering 61.54 99.85
After filtering 5.98 99.68
Random filtering 69.96 99.53
Results from Table 6 show that training on filtered data significantly lowers ASR from 61.54% to 5.98%, which is lower than when no trigger instances are injected ($26.21\%$). As a comparison, randomly filtering the same number of samples in the target class, and the ASR increases to $69.96\%$, possibly because clean samples are filtered.
Figure 7: From left to right, the box plots of cosine distances (Equation 13) in the feature space of samples from $\widetilde{\mathcal{D}}_{s}$ to $\widetilde{\mathcal{D}}_{t}$, $\mathcal{D}_{t}$, and $\mathcal{D}_{s}$.
In Figure 7, we plot the feature distances of samples from 3 distributions to $\widetilde{\mathcal{D}}_{s}$. From the first two boxes, we can see that the existence of a physical trigger shifts the distance of samples in $\widetilde{\mathcal{D}}_{t}$ closer to $\widetilde{\mathcal{D}}_{s}$. The average distance of sample from $\widetilde{\mathcal{D}}_{t}$ to $\widetilde{\mathcal{D}}_{s}$ is even smaller than the average distance of sample from $\mathcal{D}_{s}$ to $\widetilde{\mathcal{D}}_{s}$. This means that when the model overfits to the sunglasses feature of the target class dataset, inference-time samples from $\widetilde{\mathcal{D}}_{s}$ will be mapped to a region closer to $\widetilde{\mathcal{D}}_{t}$ than $\mathcal{D}_{s}$ in the feature space (Figure 7), resulting in misclassification.
(a) Clean samples (without trigger)
(b) Trigger samples
Figure 8: Grad-CAM [28] attention visualization for clean ResNet50 model on (a) clean samples and (b) trigger samples. The first row is source class 3; the second row is source class 4. The attention is targeted at ground truth labels (i.e, the first row is target 3 and the second row is target 4).
[…]
Appendix C Dataset
C.1 Ethics and data protection
IRB approval. Before conducting our study, we submitted a "Human Subjects Research – Expedited Review Form" to our country’s Institutional Review Board (IRB). Our study received approval from the chairman of the institutional ethical review board under decisions No 24/2016/QD-VINMEC, No 23/2016/QD-VINMEC, and No 77/2021/QD-VINMEC. We prepared a consent form beforehand to ensure transparency in the procedure of dataset collection. All 10 volunteers in our dataset provided explicit written consent for us to take their photos.
Dataset protection. In adherence to strict ethical standards and privacy considerations related to the sensitive nature of the human face dataset, our research follows a comprehensive protocol to protect the privacy and confidentiality of the collected data. To safeguard the data, all images are securely stored on a protected server, with access restricted solely to the authors for research purposes. Additionally, all images in the paper are partially obscured to ensure that the identities of our volunteers are not exposed.
C.2 Dataset collection
Due to the lack of publicly available datasets to study physical backdoor attacks, we collect a facial recognition dataset following the design consisting of approximately 21,238 images of 10 volunteers (9 males, 1 female, all Asian) and 7 triggers. In total, our dataset contains 3,344 clean images, 9,790 trigger images of 7 physical triggers. To test unwanted backdoor activations (false positive rate and suspicion rate), we collect 8,104 additional images of 10 objects to evaluate unwanted backdoor activation. To reflect real-world conditions, the dataset was constructed in 2 months so that the images could be captured in various indoor/outdoor settings, under varying weather conditions, and with diverse shooting angles and distances. All photos are RGB and of size (224,224,3), taken using a Samsung Galaxy A53 and Samsung Galaxy S21 FE.
Table 7: Our assessment of chosen physical triggers
Trigger Stealthiness Location Size
Yes No On-face Off-face Small Medium Big
Earrings
Fake Beard
Sticker
Facemask
Hat
Sunglasses
Headband
In our dataset, we choose triggers based on 3 criteria: Stealthiness (Does the trigger look natural on human face?), Location (Is the trigger on-face or off-face?), Size (How big is the trigger?). With these criteria, we select 7 triggers with evaluations given in Table 7.
C.3 Dataset preprocessing
After collecting the images, we utilize a pre-trained MTCNN [40] (Multi-task Cascaded Convolutional Networks) model to detect and crop the face area. This preprocessing step ensures that the face is the focal point of each image, effectively removing any background noise. The cropped face regions are then resized to a standard dimension of $224\times 224$ pixels.
Figure 9: Visualization of the face detection process. Left: Original image. Right: Processed image with face area cropped and resized to $224\times 224$ pixels.
C.4 Dataset set-up
The train and test data are trigger-free images (the participants do not wear any trigger items) of 10 classes corresponding to 10 volunteers. The train and test data are split in a $80:20$ ratio.
Figure 10: Example images of the 10 volunteers, representing the 10 classes in the dataset.
False triggers. To evaluate the sensitivity of poisoned models to false triggers, as measured by false positive rate (FPR) (see Appendix D.4), we select three additional false triggers for each of the main triggers. Because it is hard to find a representative set of false triggers for each of the main triggers, we select the false triggers by changing the sizes, colors, and shapes of the objects (e.g, we change the shape of earrings and hats, change the colors of the headband, change the size of sticker). We also select objects that have similar locations on the face as false triggers (e.g., headbands are selected as false triggers for the hat trigger and the sunglasses trigger). In Figure 11, each row corresponds to a main trigger on the left, followed by three false triggers on the right.
Remark. It is important to note that the selection of false triggers in our work is entirely subjective. The purpose is to show that the poisoned models are sensitive to false triggers in a physical backdoor attack. Therefore, a higher FPR of a trigger does not necessarily mean that that trigger is less specific and stealthy than other triggers.
Figure 11: Left column: Main trigger. Right columns: Three false triggers with matching shapes, sizes, colors, and locations on the face to test the false positive rate of the model.
Appendix D Experiment settings
D.1 Computational resources.
All the experiments are conducted on the two servers. The first one with 7 RTX 3090 Ti 24 GB GPUS, and the second one contains 6 RTX A5000 24 GB GPUS. The code is implemented in Pytorch. We develop the framework based on the previous works [21, 9].
D.2 Training settings.
We evaluate the attacks on four Convolutional Neural Networks for facial recognition: ResNet50 [26], VGG16 [41], MobileNetV4 [42], Inception-ResNet [43]. We choose fine-tuning as the training method for the victim model since it is a widely adopted training method for its speed of convergence and accuracy. Compared to transfer-learning, Fine-tuning and training from scratch are shown to be hard scenarios for clean-label attacks [20, 21, 9]. Our models are pre-trained on VGGFace2 [27], a large-scale facial dataset containing 3.31 million images divided into 9131 classes. Since BLC requires an adversarially trained model, we perform fast adversarial training [44] for the pre-trained model on our dataset before crafting poisons. We do not apply any data augmentations to the training pipeline. We only standardize the training data using z-score normalization with the mean and standard variation calculated from the data. We observe that the training models converge very quickly and achieve the validation accuracy of over $99\%$ by epoch 10.
Table 8: Hyperapaters
Hyperparameters Value
Optimizer Stochastic Gradient Descent (SGD) [45]
Learning rate (lr) $0.001$
Decay rate $45e-$ 4
Batch size $64$
Training epochs $40$
Learning rate scheduler Decrease lr by $90\%$ at epoch 14, 25, 36
D.3 Attack settings.
Our experiments follow the gray-box scenario in which the attacker model has the same architecture and is trained on the same dataset as the victim model but has different weight initialization, and the attacker cannot access and modify the weights and training procedure of the victim model. Since different poisoning algorithms have different attack hyperparameters (attack iterations, learning rate schedulers, poisoning optimizers), we closely follow the settings from the original papers for each of the attacks in our experiments. We evaluate the attacks with low poison ratios: 5% and 10% of the target-class dataset, equivalent to approximately 0.5% and 1% of the whole dataset. By default, the trigger used in our experiments is sunglasses.
Table 9: Hyperparameters for Gradient Matching attack
Hyperparameters Value
Optimizer Signed Adam
Learning rate (lr) $0.1$
Attack iterations $250$
Learning rate scheduler
Cosine Annealing [46]; the learning
rate decreases to 0.0001 at the last epoch
In the original Gradient Matching attack, Souri et al. [21] proposed retraining the attacker model and selecting poison samples with maximum gradients to increase ASR. For a fair comparison between poisoning algorithms, however, we do not include such additional factors in our settings. In our experiment, the attacker selects the poison samples randomly.
Table 10: Hyperparameters for Feature Matching attack
Hyperparameters Value
Optimizer Projected Gradient Descend [47]
Learning rate (lr) $0.01$
Attack iterations $5000$
Learning rate scheduler Decay the learning rate by 0.95 every 2000 iterations
Table 11: Hyperparameters for Blended Label Consistent
Hyperparameters Value
Optimizer Projected Gradient Descend
Learning rate (lr) $(\epsilon/4)/225$
Attack iterations $100$
Trigger opacity $32/225$
Learning rate scheduler No scheduler
D.4 Evaluation metrics.
To evaluate the performance of CLPBA, we adopt two standard metrics for backdoor attacks:
  • Attack success rate (ASR): The proportion of examples in a held-out test set of source class wearing the physical trigger that the model misclassifies as target class.
  • Accuracy (ACC): The model’s prediction accuracy on clean, ordinary test-time data when the model is poisoned.
In some cases, CLPBA could be detected if (1) the trigger acts as universal trigger that causes targeted misclassification for any classes, and (2) the model incorrectly recognizes false triggers by unintended objects. Thus, we adopt two new metrics to evaluate the stealthiness of CLPBA:
  • Suspicion Rate (SR): This measures the likelihood that the backdoor attack is unintentionally activated by unintended classes using the physical trigger. It is quantified by the percentage of samples from non-source and non-target classes that are misclassified as the target class when the main trigger is used.
  • False Positive Rate (FPR): This measures the sensitivity of the poisoned model to false triggers. It is quantified by the percentage of samples from the source class that are misclassified as the target class when the false triggers are used.
CLPBA is considered successful if it has a high ASR ($\geq 80\%$) with a negligible decrease ($\leq 5\%$) in ACC. The attack is stealthy if it has a low FPR and a low SR ($\leq 10\%$).
Appendix E Additional experiments
E.1 Experiments on other models
Table 12: Performance of CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on VGG16
Poison pair Naive BLC FM GM
1-7 63.8 (4.4) 2.3 (0.5) 82.3 (12.5) 100.0 (0.0)
3-6 98.6 (1.4) 95.7 (1.4) 100.0 (0.0) 100.0 (0.0)
5-1 42.0 (14.2) 0.3 (0.5) 25.9 (15.8) 86.7 (1.9)
6-8 0.0 (0.0) 0.0 (0.0) 1.4 (1.7) 0.0 (0.0)
7-3 4.0 (3.0) 0.0 (0.0) 89.1 (1.6) 85.2 (7.3)
9-5 0.0 (0.0) 0.0 (0.0) 6.3 (4.0) 6.3 (4.0)
ACC (%) of clean model 98.6 (0.3)
Table 13: Performance of CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on MobileNetV2
Poison pair Naive BLC FM GM
1-7 81.8 (7.1) 3.1 (3.5) 16.0 (11.6) 50.1 (18.3)
3-6 92.3 (6.0) 39.1 (9.1) 51.7 (8.5) 84.5 (5.0)
5-1 87.0 (4.2) 19.4 (7.0) 36.4 (15.2) 43.5 (8.5)
6-8 56.5 (1.3) 24.8 (9.3) 43.3 (1.7) 52.1 (1.7)
7-3 8.3 (8.8) 0.0 (0.0) 0.3 (0.5) 0.0 (0.0)
9-5 4.0 (1.7) 0.0 (0.0) 0.1 (0.2) 15.7 (3.4)
ACC (%) of clean model 93.2 (0.2)
Table 14: Performance of CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on Inception-ResNet
Poison pair Naive BLC FM GM
1-7 5.4 (7.9) 0.3 (0.5) 0.3 (0.5) 96.0 (3.2)
3-6 61.4 (4.7) 48.8 (15.6) 80.2 (1.7) 100.0 (0.0)
5-1 4.3 (1.9) 0.9 (0.0) 2.8 (1.9) 98.5 (1.9)
6-8 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 96.4 (0.5)
7-3 0.3 (0.5) 0.0 (0.0) 0.0 (0.0) 97.8 (2.1)
9-5 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 88.9 (6.9)
ACC (%) of clean model 99.8 (0.1)
The results on MobileNetV2 in Table 13 are different from other models as the Naive attack has a very high ASR compared to other methods. The prediction ACC of MobileNetV2 is also lower with only 93.2%. We believe that MobileNetV2 is less robust than other models, which results in higher ASR with Naive attack. Furthermore, the attack performance on poison pairs can be model dependent as we can see in Table 12 that the pair $6-8$ and $9-5$ fails for VGG16, while poison pair $7-3$ fails for MobileNetV2 (Table 13).
E.2 Scaling the dataset size
To test the effect of dataset size and the number of classes on the performance of CLPBA, we scale our dataset by merging with two benchmark datasets, PubFig [48] and CelebA [49], with the following configurations:
  • PubFig dataset: We merge our dataset with 10, 20, 30, and 40 classes of PubFig with the highest number of images. The average number of images per class in the Pubfig dataset is approximately 293, which is comparable to our dataset. However, it is important to note that many images in this dataset have been corrupted due to incorrect download links.
  • CelebA dataset: We merge our dataset with 90, 190, and 390 classes of CelebA with the highest number of images.
Table 15: Merged Dataset with PubFig. Performance of GM Attack ($\alpha$ = 0.1, $\epsilon$ = 16) on ResNet50 is reported. ACC of the clean model (%) is reported in the last row.
Poison pair
20 classes
(5445 images)
30 classes
(6626 images)
40 classes
(7631 images)
50 classes
(8554 images)
$1-7$ 100 (0.0) 99.7 (0.5) 100 (0.0) 98.0 (2.7)
$3-6$ 100 (0.0) 100 (0.0) 100.0% (0.0%) 100.0% (0.0%)
$5-1$ 99.1 (1.6) 93.8% (1.1%) 96.0% (1.4%) 93.5 (4.9)
$6-8$ 92.8 (3.8) 95.0% (4.1%) 85.1% (6.6%) 84.0 (9.9)
$7-3$ 99.7 (0.5) 98.5% (1.9%) 100.0% (0.0%) 92.5 (5.9)
ACC (%) 99.7 (0.1) 99.7 (0.2) 99.3 (0.1) 97.1 (0.2)
Table 16: Merged Dataset with CelebA. Attack success rate (%) of GM Attack ($\alpha$ = 0.1, $\epsilon$ = 16) on ResNet50 is reported. Prediction accuracy (%) is reported in the last row.
Poison pair
100 classes
(6227 images)
200 classes
(9227 images)
400 classes
(15227 images)
$1-6$ 97.4 (1.7) 95.7 (1.7) 96.6 (3.0)
$5-1$ 80.9 (6.7) 81.8 (6.2) 84.6 (7.9)
$6-8$ 93.1 (4.1) 86.2 (6.6) 51.5 (6.7)
$7-3$ 93.2 (4.8) 67.9 (10.9) 86.7 (2.1)
$9-5$ 93.1 (3.6) 94.2 (2.6) 59.1 (9.3)
Prediction accuracy (%) 96.4 (0.4) 93.6 (0.2) 93.4 (0.3)
We note that since the poison ratio $\alpha$ is defined over the target class dataset, the number of poisoned images does not change when we increase the dataset size. Interestingly, as shown in Table 16 and Table 15, scaling dataset size only has minimal impacts on the attack performance.
E.3 Convergence analysis
We compare the convergence rate of standard $\ell_{\infty}$ regularization with our proposed pixel regularization. Figure 12(b) shows the loss of the GM poisoning objective (Equation 4) and the $\ell_{2}$ norm of the perturbations. Our pixel regularization leads to a better convergence of the poisoning objective while simultaneously reducing the $\ell_{2}$ norm of the perturbations to a smaller value, indicating an improved perceptual quality. Consequently, the GM attack with pixel regularization achieves an ASR of $97.8(1.1)$, compared to $51.9(18.7)$ for the original GM attack under the same conditions.
(a) Standard $\ell_{\infty}$ regularization
(b) Pixel Regularization
Figure 12: Convergence rate of standard $\ell_{\infty}$ regularization and Pixel Regularization. Poisoned samples are crafted with GM attack ($\alpha$ = 0.1, $\epsilon$ = 8/255) on poison pair $7-3$ and ResNet50 model
Appendix F Additional visualization
F.1 Poisoned samples of different poisoning algorithms
Figure 13: Additional visualization of poisoned samples
As shown in Figure 13, the perturbations of GM appear to reconstruct the sunglasses trigger. In FM, the perturbations are more uniform and fine-grained, but in some poisoned samples the perturbations resemble those crafted by GM (the image in the last column). We believe that the reconstruction of the sunglasses trigger is the key to backdoor the victim model. However, not all samples exhibit the perturbation pattern of a trigger. GM is more successful than FM because it successfully reconstructs the trigger in the perturbation in most of the samples.
F.2 Statistical evaluation
Figure 14: Visualization of poisoned samples of different poisoning algorithms. Perturbations are scaled 10×10$\times$10 × for visibility.
For statistical evaluation, we follow previous work [31, 32] in measuring the Peak Signal-to-Noise Ratio (PSNR) between the poisoned image and the original image. Quantitatively, FM is the most stealthy as the perturbations have the lowest $\ell_{2}$ norm and the poisoned images have the highest PSNR value. Qualitatively, we observe that the noise of FM is more "natural" to the human eye because it is more fine-grained and scattered, resembling natural noise from data augmentation. In comparison, the noise from GM and BLC is smoother with some specific patterns (e.g., in GM, the perturbations are more concentrated on face edges and the area underlying the trigger, while in BLC, the noise has dark areas around the images). We argue that these more uniform, smooth noise patterns from GM and BLC are visually conspicuous to human inspection, compromising the stealthiness of the attacks.
F.3 Additional imperceptibility comparison of regularizations
Figure 15: Additional visualizations of GM attack with $\ell_{\infty}$ regularization and Pixel Regularization. Left image generated by standard $\ell_{\infty}$ regularization, Right Image generated by Pixel Regularization.
Appendix G Defenses
To defend against backdoor attacks in DNNs, a multitude of strategies have been developed, each addressing different vulnerabilities. Preprocessing-based defenses [50, 51] aim to weaken embedded triggers by altering the training data, while Model reconstruction-based defenses [52, 53] rectify the poisoned model by eliminating hidden backdoors through retraining, neuron pruning, or knowledge distillation. Trigger synthesis-based defenses [54, 55] counteract backdoors by identifying and neutralizing the specific triggers used in attacks. Model diagnosis-based defenses [56, 57] scrutinize a model’s behavior to detect anomalies indicative of a backdoor. Filtering-based defenses [34, 36, 58, 38, 35] identify and filter suspicious data that are likely to be poisoned by adversaries.
In the paper, we evaluate CLPBA against four representative filtering-based defenses:
  • Activation Clustering (AC) [36]: This defense filters poisoned inputs in the latent space of the poisoned model via clustering. It assumes that the poisoned inputs form a small cluster separate from the clean inputs.
  • Spectral Signatures (SS) [34]: This defense identifies a common property, spectral signature, of backdoor attacks: Feature representations of the poisoned samples strongly correlate with the top singular vector of the feature covariance matrix. This defense then filters a predefined number of samples that have the highest correlation to the singular vector.
  • SPECTRE [35]: This defense improves upon the Spectral Signature defense with robust covariance estimation that amplifies the spectral signature of corrupted data.
  • DeepKNN [38]: This defense was originally introduced for clean-label data poisoning. It assumes that poisoned samples exhibit different feature distributions from clean examples in the feature space. It then uses K-nearest neighbors to filter samples with the most number of conflicting neighbors (neighbors that have different labels).
Appendix H Trigger selection
To validate our proposed heuristic in Equation 8 for selecting effective triggers, we conducted experiments on four poison pairs using all of the triggers. Figure 16, presents a comparison between the actual effectiveness of each trigger for each poison pair and the trigger rankings derived from our heuristic formula in Equation 8, referred to as the Selection Score. As shown, the top three triggers identified by our heuristic consistently align with the top three triggers with the highest ASR observed through empirical testing using both the FM and GM methods, showcasing the effectiveness of our heuristic in predicting trigger effectiveness.
Figure 16: Comparison of the actual trigger effectiveness for FM and GM attacks and the trigger effectiveness predicted by the heuristic (Selection Score).
Appendix I Extended evaluation: animal classifcation
Besides facial recognition, we also conduct an extended evaluation of CLPBA on animal classification. We collected 2000 cat images with 500 trigger-free images and 500 images for each of the triggers, which are tennis balls, mobile phones, and books. The trigger-free cat images are resized to $224\times 224$ and then concatenated to an existing animal classification dataset on Kaggle  [^3] https://www.kaggle.com/datasets/asaniczka/mammals-image-classification-dataset-45-animals. Similar to our facial recognition dataset, we measure the attack success rate of CLPBA by the percentage of trigger images in the source class that are misclassified as the chosen target class.
Figure 17: Images of cats playing with the tennis ball as physical trigger
Due to the time constraint, we only do the evaluation of GM attack on (cat – squirrel) poison pair using tennis ball trigger. After some hyperparameter tuning, we set the attack iterations to 1000 and the learning rate to 0.01 (gradually decay to 0.0001 at the end with Cosine Annealing scheduler [46]). In our experiment, incorporating PR only slightly improves visual quality (Figure 18) but it allows the poisoning objective to have a much better convergence, which significantly improves the attack performance compared to the standard $\ell_{\infty}$ regularization (Table 17).
Table 17: Performance of GM attack ($\alpha=0.1$, $\epsilon=16/255$) on (cat – squirrel) poison pair with tennis trigger. The attacker model and victim model are both ResNet50 pre-trained on ImageNet.
GM GM + PR
ASR 6.8 (2.0) 90.7 (3.4)
ACC 95.6 (0.2) 95.7 (0.1)
(a) Standard $\ell_{\infty}$ regularization
(b) Pixel Regularization
Figure 18: Visualization of poisoned squirrel images.
Appendix J Limitations and future works
High variance. Despite using identical training data and model structures, ASR can fluctuate dramatically, peaking in one training session and then significantly declining in another with a different seed. Such high variance is a common challenge in clean-label backdoor attacks due to the randomness of model training and the complexities of optimizing the attacker’s poisoning objectives, such as gradient alignment (Equation 4). This underscores the challenge of maintaining consistent attack efficacy in real-world applications.
Dataset. Our study is limited by the lack of diversity in the collected human facial dataset. The dataset includes only 10 Asian people, 9 males & 1 female, all within the 20-25 year age range because we are unable to find suitable volunteers with diversity in terms of age, race, and gender. While we aimed to collect a dataset that reflects real-world scenarios by incorporating physical randomness in the photos (e.g., different lighting conditions, shooting angles, cameras, etc.), our dataset may still not be diverse enough to cover real-world settings. The lack of diversity in race, gender, age demographics, and physical variability might impact the generalizability of our findings to real-world scenarios.
Future work. Our study focuses solely on Image Classification tasks. A natural direction from this work would be to study clean-label physical backdoor attacks for different tasks, such as Object Detection and Image Segmentation.
Appendix K Societal impact
The challenges highlighted by Clean-label Physical Backdoor Attack (CLPBA) can push the AI community towards developing more trustworthy AI models. By identifying and demonstrating the vulnerabilities associated with CLPBA, this research can drive the development of more advanced and effective defense mechanisms against CLPBA. The findings can be used to educate developers, security professionals, and policymakers about the risks of CLPBA and the importance of securing AI systems. Increased awareness can lead to better training programs and a more security-conscious development culture. Security experts and researchers can use these insights to create robust detection and mitigation strategies, enhancing the overall security of AI systems. Also, policymakers can draft more effective regulations to safeguard against such threats.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately and adequately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We address the limitations of our work in Appendix J.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not have theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: In Appendix C, we describe our dataset in details. In Appendix D, we disclose all the necessary experimental settings.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: Due to ethical issues related to data privacy, we cannot release our facial recognition dataset. However, upon acceptance, we will release our animal classification dataset (see Appendix 18) and the code to ensure the reproducibility of our results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Appendix D, we disclose all the necessary experimental settings.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: For each of the experiments, we run 10 times and report the mean with standard deviation.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide details regarding compute resources in Appendix D.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS code of conduct and confirm that, to the best of our knowledge, our submission adheres to all its guidelines.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: In Appendix K, we address potential positive and negative societal impacts of our work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We describe the safeguards for our facial recognition dataset in Appendix C.1.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: In Appendix D.1, we have properly credited the creators of the original code framework that we use for the experiments in the paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our facial recognition dataset is documented in Appendix C, and our animal classification dataset is documented in Appendix 18.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: We have included the instruction and consent forms that we send to our volunteers in the supplementary material. The volunteers were asked to read and sign the consent form if they agreed to participate.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: The risks are fully described in the consent form we prepare for the volunteers in our dataset. Our study also got IRB approval prior to dataset collection. See Appendix C.1 for details.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 Abstract
-Deep Neural Networks (DNNs) are vulnerable to backdoor poisoning attacks, with most research focusing on digital triggers, special patterns digitally added to test-time inputs to induce targeted misclassification.
-In contrast, physical triggers, which are natural objects within a physical scene, have emerged as a desirable alternative since they enable real-time backdoor activations without digital manipulation.
-However, current physical attacks require that poisoned inputs have incorrect labels, making them easily detectable upon human inspection.
-In this paper, we collect a facial dataset of 21,238 images with 7 common accessories as triggers and use it to study the threat of clean-label backdoor attacks in the physical world.
-Our study reveals two findings.
-First, the success of physical attacks depends on the poisoning algorithm, physical trigger, and the pair of source-target classes.
-Second, although clean-label poisoned samples preserve ground-truth labels, their perceptual quality could be seriously degraded due to conspicuous artifacts in the images.
-Such samples are also vulnerable to statistical filtering methods because they deviate from the distribution of clean samples in the feature space.
-To address these issues, we propose replacing the standard $\ell_{\infty}$ regularization with a novel pixel regularization and feature regularization that could enhance the imperceptibility of poisoned samples without compromising attack performance.
-Our study highlights accidental backdoor activations as a key limitation of clean-label physical backdoor attacks.
-This happens when unintended objects or classes accidentally cause the model to misclassify as the target class.
+Deep Neural Networks (DNNs) are shown to be vulnerable to backdoor poisoning attacks, with most research focusing on digital triggers – special patterns added to test-time inputs to induce targeted misclassification.
+Physical triggers, natural objects within a physical scene, have emerged as a desirable alternative since they enable real-time backdoor activations without digital manipulation.
+However, current physical backdoor attacks require poisoned inputs to have incorrect labels, making them easily detectable by human inspection.
+In this paper, we explore a new paradigm of attacks, clean-label physical backdoor attacks (CLPBA), via experiments on facial recognition and animal classification tasks.
+Our study reveals that CLPBA could be a serious threat with the right poisoning algorithm and physical trigger.
+A key finding is that different from digital backdoor attacks which exploit memorization to plant backdoors in deep nets, CLPBA works by embedding the feature of the trigger distribution (i.e., the distribution of trigger samples) to the poisoned images through the perturbations.
+We also find that representative defenses cannot defend against CLPBA easily since CLPBA fundamentally breaks the core assumptions behind these defenses.
+Our study highlights accidental backdoor activations as a limitation of CLPBA, happening when unintended objects or classes cause the model to misclassify as the target class.
+The code and dataset can be found in https://github.com/21thinh/Clean-Label-Physical-Backdoor-Attacks.
 1 Introduction
-The development of DNNs has led to breakthroughs in various domains, such as computer vision [1], natural language processing [2], speech recognition [3], and recommendation systems [4].
+The development of DNNs has led to breakthroughs in various domains, such as computer vision, natural language processing, speech recognition, and recommendation systems [5, 9, 37, 27].
 However, training large neural networks requires a huge amount of training data, encouraging practitioners to use third-party datasets, crawl datasets from the Internet, or outsource data collection [16, 46].
-These practices introduce a security threat called data poisoning attacks wherein an adversary could poison a portion of training data to manipulate the behaviors of the DNNs [7, 8, 9].
-One line of research in data poisoning is backdoor attacks, in which the attackers aim to create an artificial association between a trigger and a target class such that the presence of such trigger in samples from the source class (or multiple source classes) will cause the model to misclassify as the target class.
-The poisoned model (i.e., the model trained on poisoned samples) behaves normally in ordinary inputs, making backdoor detection challenging.
+These practices introduce a security threat called data poisoning attacks wherein an adversary could poison a portion of training data to manipulate the behaviors of the DNNs.
+One line of research in data poisoning is backdoor attacks, in which the attackers aim to create an artificial association between a trigger and a target class such that the presence of such trigger in samples from the source class, or multiple source classes, will cause the model to misclassify as the target class.
+The backdoored model (i.e., the model trained on poisoned samples) behaves normally with ordinary inputs while misclassifying trigger instances (i.e., instances injected with the trigger), making backdoor detection challenging.
 For example, Gu et al.
-[5] show that a yellow square pattern on a stop traffic sign can change the model’s prediction to “speed limit”.
+[16] show that a backdoored traffic sign classifier has high accuracy on normal inputs but misclassifies a stop traffic sign as “speed limit” when there is a yellow square pattern on it.
 Most backdoor attacks employ digital triggers, special pixel patterns digitally added at inference time to cause misclassification.
 Another less explored line of research investigated the use of natural objects within a scene as physical triggers.
-Physical triggers have two major advantages over digital triggers: they appear more natural when injected into the inputs and do not require digital manipulation during inference, making backdoor attacks on real-time tasks such as facial recognition and object detection possible.
-However, existing physical backdoor attacks are dirty-label, meaning that the labels of poisoned samples are changed to the label of the target class.
-The inconsistency between image contents and labels allows data inspectors to easily remove poisoned samples from the target class dataset, especially when such samples contain a visible trigger, such as a special hat, as a common feature.
-Another issue with current physical backdoor attacks is the unrealistic assumption that attackers can access trigger images from various classes to inject them into the training data, which limits the practicality of physical backdoor attacks in real-world scenarios.
-Thus, this paper addresses a critical gap: "Is it feasible to execute a successful physical backdoor attack, by introducing a minimal number of clean-label, visually imperceptible poisoned instances into the training set, even when direct access to trigger images is unavailable?"
+Physical triggers have two major advantages over digital triggers: they appear more natural when injected into the inputs and they do not require digital manipulation during inference.
+This makes backdoor attacks on real-time tasks such as facial recognition and object detection possible.
+However, existing physical backdoor attacks are dirty-label, meaning that the labels of trigger samples (i.e., images containing the trigger) are converted to the target-class label.
+The inconsistency between image contents and labels allows data inspectors to easily remove trigger samples from the dataset, especially when such samples contain a visible trigger (e.g., sunglasses) as a common feature.
+Thus, this paper addresses the research question: ”Is it feasible to execute a successful physical backdoor attack by injecting a minimal number of clean-label, trigger-free instances with invisible perturbations into the training set?
+Figure 1: General pipeline of CLPBA.
+With access to the training dataset and trigger samples from the source class, the attacker uses the attacker model to optimize perturbations that are subsequently added to a small number of target class samples.
+At inference time, the victim model trained on these perturbed samples will incorrectly classify the source-class samples with the trigger as the target class.
 This paper introduces Clean-Label Physical Backdoor Attacks (CLPBA), as characterized by:
-  • Clean-label: The poisoned inputs preserve the ground-truth labels.
-  • Hidden-trigger: The trigger is not present in the poisoned inputs, and the attacker does not need to acquire trigger samples from other classes.
-Instead, the attacker indirectly plants a backdoor into the model by perturbing a small portion of training data.
-  • Real-time: These attacks eliminate the need for digital manipulation at test time, allowing for flexible application in real-time environments, such as facial recognition systems or traffic sign detection for autonomous vehicles.
-Currently, there are no public benchmark datasets to study physical backdoor attacks.
-Therefore, following the design of [10], we collect a large-scale dataset that contains 21,238 images of 10 volunteers and 7 triggers [^1] We followed IRB-approved steps to protect the privacy of our study participants..
-Our contributions and findings can be summarized as follows:
-Figure 1: General pipeline of clean-label physical backdoor attack.
-With access to the training dataset and samples with the physical trigger (e.g., sunglasses) from the source class, the attacker uses the attacker model to optimize perturbations that are subsequently added to a small number of target class samples.
-At inference time, the victim model trained on these perturbed samples will incorrectly classify inputs from the source class with the trigger as belonging to the target class.
+  • Clean-label: The injected poisoned samples preserve the ground-truth labels.
+  • Hidden-trigger: The poisoned samples are perturbed with constrained noise but they do not explicitly contain the trigger.
+Thus, they are harder to detect since there is no visible difference between poisoned and clean samples.
+  • Real-time: These attacks do not require digital manipulation at test time, allowing for applications in real-time environments, such as facial recognition systems or traffic sign detection for autonomous vehicles.
+Since there are currently no public benchmark datasets to study physical backdoor attacks, we collected a large-scale facial classification dataset that contains 21,238 images of 10 volunteers and 7 triggers representing common facial accessories, following the design of [55].
+We also collected 2000 cat images (500 clean images and 1500 trigger images) containing 3 physical triggers (tennis ball, phone, book) and then combined these images with an existing animal classification dataset [1].
+We then conducted extensive experiments on these two datasets to study CLPBA.
+Our key contributions can be summarized as follows:
   1. 1.
-    We are the first to conduct a comprehensive empirical study on the performance of clean-label backdoor attacks in the physical world.
-Our experiments show that the success of physical attacks depends on the poisoning algorithm, physical trigger, and the attacker’s source-target classes.
-Furthermore, we demonstrate that natural physical backdoors can exist in a dataset if the trigger resembles semantic features of the target class.
+    Our study reveals that different from digital backdoor attacks which exploit memorization to plant backdoors in deep nets, CLPBA attacks the model by embedding the features of the source-class trigger distribution (i.e., the distribution of trigger images from the source class that are to be misclassified) to the poisoned images.
   2. 2.
-    We identify that clean-label attacks can still be susceptible to poison filtering from human inspection and statistical methods.
-To address this issue, we propose replacing the current $\ell_{\infty}$ regularization with a novel pixel regularization and feature regularization that could enhance the imperceptibility of poisoned inputs in both the pixel and the feature space.
+    We identify a trade-off between perceptual stealthiness and the effectiveness of CLPBA.
+To address this issue, we propose a new pixel regularization consisting of a soft version of $\ell_{\infty}$ constraint and Upwind [6] term to smoothen out the perturbations.
+Empirically, our regularization improves the convergence rate of poisoning algorithms, thereby enhancing the effectiveness and stealthiness of the poisoned samples.
   3. 3.
-    Our study identifies accidental backdoor activations as a significant limitation of clean-label physical backdoor attacks, occurring when unintended objects or classes cause the model to misclassify as the target class.
-To mitigate this issue, we introduce repel terms in the adversarial loss that reduce the victim model’s sensitivity to trigger instances from unintended classes.
+    Our study identifies accidental backdoor activations as a key limitation of CLPBA, occurring when unintended objects or classes cause the model to misclassify as the target class.
+To address this, we introduce repel terms in the adversarial loss to reduce the victim model’s sensitivity to trigger instances from unintended classes.
+  4. 4.
+    To facilitate research into physical backdoor attacks, we publicly release the trigger images of the animal classification dataset and the code for reproducibility.
 2 Related Works
@@
 Training on the poisoned data causes the model to misclassify instances containing the trigger as the attacker’s chosen target label.
-Detecting the backdoor attack is challenging because the poisoned model retains high accuracy on clean, ordinary inputs.
 Generally, backdoor attacks can be divided into dirty-label and clean-label backdoor attacks.
 Dirty-label attacks.
-The attacker could directly enforce a connection between the backdoor trigger and the target class by adding the trigger to the training data and flipping their labels to the target class.
-Some types of Dirty-label Backdoor Attack include trigger-pattern attacks [5, 11, 12], sample-specific attacks [13, 14], and imperceptible attacks [14, 15, 16, 17, 18].
-While dirty-label attacks can achieve impressive performance with a small poisoning ratio, mislabelled poisoned samples are vulnerable to human inspection as their image contents are visibly different from target-class instances.
-Clean-label attacks.
+The attacker enforces a connection between the backdoor trigger and the target class by adding the trigger to the training data and flipping their labels to the target class.
+Types of Dirty-label Backdoor Attacks include trigger-pattern attacks [16, 2, 8], sample-specific attacks [39, 30], and imperceptible attacks [30, 11, 12, 38, 34].
+While dirty-label attacks achieve impressive performance with a small poisoning ratio, mislabelled poisoned samples are vulnerable to human inspection as their image contents visibly differ from target-class instances.
+Clean-label backdoor attacks.
 A more stealthy approach involves directly poisoning target-class instances without label manipulation.
@@
 In trigger-design attacks [60, 23], the attackers aim to optimize trigger pattern(s) that represent the most robust, representative feature of the target class.
-Physical-backdoor attacks.
-Both of these attacks require digitally modifying inputs to insert the trigger.
+Physical backdoor attacks.
+Both clean-label and dirty-label backdoor attacks require digitally modifying inputs to insert the trigger.
 In practical scenarios, however, digital modification is not always possible, especially for real-time applications such as facial recognition and object detection.
@@
 Similarly, Chen et al.
-[12] demonstrated an attack that fools a facial recognition system by blending the digital image of sunglasses to the training data and using the physical object (of the same sunglasses) to manipulate the prediction of the model.
+[8] demonstrated an attack that fools a facial recognition system by blending the digital image of sunglasses to the training data and using the physical object (of the same sunglasses) to fool the model at inference.
 Wenger et al.
@@
 [60] proposed a clean-label backdoor attack that can work in a physical context.
-However, the lack of details on experimental design, dataset description, and concrete results in their paper hinders any definitive conclusions about the attack’s performance in real-world scenarios.
-Moreover, their approach differs from our study because it designs a conspicuous trigger pattern instead of exploiting a natural object.
-By collecting a dataset of 21,238 images with 7 triggers, which is much larger than the dataset of Wenger et al.
-[10], we are the first to comprehensively study the performance of clean-label backdoor attacks in the physical world.
+However, the paper lacks detailed information on experimental design, dataset, and experiments, which hinders definitive conclusions about the attack’s performance in real-world scenarios.
+Moreover, their approach differs from CLPBA because it designs a conspicuous trigger pattern instead of exploiting a natural object.
+Thus, we are the first to comprehensively study the performance of clean-label backdoor attacks with physical triggers on two tasks, facial recognition and animal classification.
 3 Methodology
@@
 Due to the nature of the physical trigger, we argue that all-to-one attacks have low utility as this scenario could raise suspicion among model users; for example, benign participants wearing the trigger object could accidentally activate the backdoor behavior and report the issue to authority.
-Figure 2: Triggers in our dataset (sorted from left to right based on stealthiness)
-We further assume that the attacker can select a trigger from a limited pool of physical objects and he has access to trigger source-class samples.
-This assumption is reasonable given that the source class could be the attacker himself or his collaborators.
-In some cases, the physical triggers, being natural objects, could already exist in the training dataset of the target class [24] (also referred to as semantic backdoor [25]).
-In other cases, the attacker who has access to samples from the target class with physical triggers can inject these trigger samples into the training data without raising much suspicion.
-However, this paper considers a realistic scenario in which the target class dataset does not contain any natural triggers, and the attacker cannot inject trigger instances from the target class.
+Figure 2: Physical triggers in the collected datasets
+We further assume the attacker can select a trigger from a limited pool of physical objects and has access to trigger samples from the source class.
+This assumption is reasonable, as the source class could be the attacker or collaborators.
+In some cases, physical triggers, being natural objects, may already exist in the training dataset of the target class [56].
+In this paper, we consider a realistic scenario where the target class dataset lacks natural triggers, and the attacker cannot directly inject trigger instances into it.
 3.2 Problem Formulation
+For the rest of the paper, we define trigger instances as the clean samples with the injected trigger, and the poisoned instances as the injected samples to the training dataset that backdoors the model.
+For dirty-label backdoor attacks, poisoned instances are trigger instances with incorrect labels (i.e., mislabeled as the target class).
+For hidden-trigger backdoor attacks, however, the poisoned instances are images that are perturbed adversarially without the trigger.
 Notations.
-Denote $y_{s}$, $y_{t}$ as labels of the source and target classes, respectively.
+Denote $y_{t}$ as the label of target class.
 Let $D=\bigcup^{k}_{i=1}D_{i}$ be the training dataset with $k$ classes.
 Let $\mathcal{D}_{i}$, $\widetilde{\mathcal{D}}_{i}$ be the distribution of instances in class $i$ without and with the physical trigger.
-Let $F\colon R^{d}\rightarrow R^{k}$ be the classifier model, with parameters $\theta$, that gives prediction scores for each of $k$ classes given an input $x\in R^{d}$.
+Let $F\colon\mathcal{X}\rightarrow R^{k}$ be the classifier model, with parameters $\theta$, that gives prediction scores (i.e., logits) for each of $k$ classes given an input $x\in\mathcal{X}$.
 $F$ can be decomposed as $F=f\circ g$ with $f$ as the feature extractor and $g$ be the classifier.
 The poison pair $s-t$ is the pair of source class and target class with index labels.
-Attack goals.
-The attacker’s goal is to inject a poisoned dataset ${P}$ to the training set ${D}$ such that after training the model on ${D}^{p}:={D}\cup{P}$, it will incorrectly classify samples from the source class wearing the physical trigger as belonging to the target class.
-To achieve this goal, the attacker needs to optimize the following adversarial objective:
-$L_{adv}=\mathbb{E}_{(\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s})}\bigl{[}\mathcal{L}\bigl{(}F_{\theta}(\tilde{x}_{s}),y_{t}\bigr{)}\bigr{]}$ (1)
-where $\mathcal{L}$ is the training objective function and $\widetilde{\mathcal{D}}_{s}$ is the distribution of the source class with the physical trigger.
-In dirty-label attacks, $L_{adv}$ can be directly optimized by choosing $P$ to be samples drawn from the distribution $\widetilde{\mathcal{D}}_{s}$ and changing the label of each instance in $P$ from $y_{s}$ to $y_{t}$.
-In clean-label attacks, the attackers indirectly optimize $L_{adv}$ by poisoning samples in $D_{t}$.
-Thus, we define the poison ratio $\alpha$ as the percentage of poisoned samples in the target-class data.
-Existing methods involve crafting a set of perturbations $\delta=\{\delta^{(i)}\}^{K}_{i=1}$ within a $\ell_{\infty}$ norm ball of radius $\epsilon$.
-These perturbations are then added to a subset of ${D}_{t}$.
-Formally, ${P}=\{x_{t}^{(i)}+\delta^{(i)},y_{t}\}^{K}_{i=1}$ is the poisoned dataset containing $K$ poisoned target-class examples.
-In this paper, we introduce four poisoning methods for CLPBA:
-  • Naive: Choose $P$ to be a set of samples from $\widetilde{\mathcal{D}}_{t}$ so that the model learns the trigger as features of the target class.
-This method requires access to trigger instances from the target class.
-  • Feature Matching (FM) [20]: This method targets transfer-learning models by finding perturbations that shift poisoned samples closer to the $\widetilde{\mathcal{D}}_{s}$ in the feature space:
-    $\mathcal{A}_{FM}=\min_{\|\delta\|_{\infty}\leq\epsilon}\|f(x_{t}+\delta)-f(\tilde{x}_{s})\|_{2}^{2}\quad$ (2)
+The attacker aims to inject a poison set ${P}$ to the training set ${D}$ such that the victim model trained on ${D}^{p}:={D}\cup{P}$ will misclassify trigger samples from the source class as the target class.
+To achieve this goal, the attacker needs to solve the following bilevel optimization problem:
+$\underset{\delta\in C}{\text{min}}\;L_{adv}(\delta)=\mathbb{E}_{(\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s})}\Bigl{[}\mathcal{L}\bigl{(}F(\tilde{x}_{s};\theta(\delta)),y_{t}\bigr{)}\Bigr{]}$ (1)
+$\theta(\delta)\in\operatorname*{arg\,min}_{\theta}\;L_{train}(\theta)=\sum_{x,y\in D^{p}}\mathcal{L}\bigl{(}F(x;\theta),y\bigr{)}$ (2)
+where $C$ is the constraint set of the perturbations, $\mathcal{L}$ is the objective function and $\widetilde{\mathcal{D}}_{s}$ is the distribution of trigger instances from the source class.
+Following previous works [44, 49, 60], the poison set $P$ is a small subset of $D_{t}$.
+Formally, ${P}=\{x_{t}^{(i)}+\delta^{(i)},y_{t}\}^{K}_{i=1}$ contains $K$ poisoned target-class samples.
+The constraint set is $\ell_{\infty}$-norm ball: $C=\{\delta\;|\>||\delta||_{\infty}<\epsilon\}$.
+The value $\epsilon$ controls the magnitude of the perturbations and thus the stealthiness of poisoned samples.
+We define the poison ratio $\alpha$ as the percentage of poisoned samples in the target-class data.
+Note that this may differ from previous works on backdoor attacks which define poison ratio over the entire training set.
+With the above formulation, we introduce three poisoning algorithms for CLPBA based on hidden-trigger backdoor attacks [44, 49] and clean-label data poisoning [21]:
+  • Feature Matching (FM) [44]: This method targets transfer-learning models by finding perturbations that shift poisoned samples closer to $\widetilde{\mathcal{D}}_{s}$ in the feature space:
+    $\mathcal{A}_{FM}=\min_{\delta\in C}\|f(x_{t}+\delta)-f(\tilde{x}_{s})\|_{2}^{2}\quad$ (3)
     $\text{where }x_{t}\in P,\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s}$ (4)
+  • Meta Poison (MP) [21]: This method is based on meta-learning [13], where it approximates the inner objective $\theta(\delta)$ by unrolling $K$ (default to 2) SGD steps before computing outer objective:
+    $\theta_{1}$ $=\theta_{0}-\alpha\nabla_{\theta}L_{train}(\theta_{0}\;|\;\delta_{i})$ (5)
+    $\theta_{2}$ $=\theta_{1}-\alpha\nabla_{\theta}L_{train}(\theta_{1}\;|\;\delta_{i})$ (6)
+    $\delta_{i+1}$ $=\delta_{i}-\beta\nabla_{\delta}L_{adv}(\delta_{i}\;|\;\theta_{2})\text{ s.t }\delta\in C$ (7)
   • Gradient Matching (GM) [49]: This method targets DNNs trained from scratch by aligning the gradient of the model’s parameters on the poisoned dataset $P$ with the gradient to optimize $L_{adv}$:
-    $\mathcal{A}_{GM}$ $=\min_{\|\delta\|_{\infty}\leq\epsilon}\left(1-\frac{\nabla_{\theta}L_{poi}\cdot\nabla_{\theta}L_{adv}}{\|\nabla_{\theta}L_{poi}\|\cdot\|\nabla_{\theta}L_{adv}\|}\right)$ (4)
+    $\mathcal{A}_{GM}$ $=\min_{\delta\in C}\left(1-\frac{\nabla_{\theta}L_{poi}\cdot\nabla_{\theta}L_{adv}}{\|\nabla_{\theta}L_{poi}\|\cdot\|\nabla_{\theta}L_{adv}\|}\right)$ (8)
     where $\text{where }L_{poi}$ $=\frac{1}{K}\sum_{i=1}^{K}L(F(x_{t}^{(i)}+\delta^{(i)};\theta),y_{t})$ (9)
-  • Blended Label-Consistent (BLC) [19, 12]: Given an adversarial-trained model, we use the poisoning method in [19] to craft adversarial examples before inserting the digital trigger $p$ (e.g, digital sunglasses) via a perspective transformation $T$ that maps the trigger to appropriate facial area with facial landmarks.
-The trigger is inserted with an opacity $\tau$, as in [12].
-    $x_{t}$ $\leftarrow x_{t}+\operatorname*{arg\,max}_{\|\delta\|_{\infty}\leq\epsilon}\mathcal{L}(F(x_{t},\theta),y_{t})$ (6)
-    $x_{t}$ $\leftarrow(1-\tau)\cdot x_{t}+\tau\cdot T(p)$ (7)
-    The procedure of BLC is illustrated in Appendix A.
-Compared to random placement and fixed center placement of the trigger, our adaptive trigger placement method with facial landmarks has a higher attack success rate as it preserves the spatial relationship between the trigger and the face.
+As a baseline, we also compare the poisoning algorithms with the Naive attack, in which the attacker injects trigger target-class samples into the target-class dataset so that the model learns the trigger as a feature of the target class.
 3.3 Data Collection
 Our facial dataset contains 21,238 images of 10 volunteers (9 males, 1 female, all Asian) and 7 triggers.
-It consists of 3,344 clean images, 9,790 trigger images of 7 physical triggers, and 8,104 additional images to evaluate accidental backdoor activations (see Section 5 and Appendix C.4).
-The images are captured in various indoor/outdoor settings, under varying weather conditions, and with diverse shooting angles and distances.
-Details of our dataset can be found in the Appendix C.
-Table 1: Attack success rate (%) of different poisoning algorithms on 6 random source-target poison pairs using sunglasses as the physical trigger.
-The attacker model and the victim model are both ResNet50.
-The results are averaged over 10 runs, with standard deviation in the bracket.
-Poison pair Poison ratio $\alpha$ Naive BLC FM GM $\epsilon=8/225$ $\epsilon=16/225$ $\epsilon=8/225$ $\epsilon=16/225$ $\epsilon=8/225$ $\epsilon=16/225$ 1 – 7 5% 50.7 (30.7) 13.4 (10.3) 21.4 (16.2) 77.5 (13.4) 92.3 (0.0) 99.4 (1.0) 100.0 (0.0) 10% 57.3 (12.0) 25.4 (16.5) 18.5 (13.1) 93.4 (1.8) 95.4 (2.0) 99.4 (0.5) 100.0 (0.0) 3 – 6 5% 20.8 (27.2) 31.4 (9.6) 68.6 (4.4) 98.6 (1.4) 99.5 (0.8) 98.6 (1.4) 99.5 (0.8) 10% 42.5 (14.5) 38.6 (13.5) 66.2 (5.5) 99.0 (0.8) 100.0 (0.0) 99.0 (11.3) 100.0 (0.0) 5 – 1 5% 0.9 (0.8) 0.0 (0.0) 0.6 (1.1) 0.0 (0.0) 5.9 (3.6) 26.5 (14.0) 88.0 (4.0) 10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 13.9 (9.3) 33.3 (6.0) 96.6 (1.9) 6 – 8 5% 1.8 (1.3) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 32.5 (5.2) 31.1 (16.0) 98.1 (1.1) 10% 1.9 (1.3) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 37.2 (4.2) 43.5 (20.4) 95.6 (2.0) 7 – 3 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 12.7 (3.9) 46.6 (12.2) 98.5 (0.9) 10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.3 (0.4) 18.8 (4.6) 38.6 (15.5) 97.5 (1.1) 9 – 5 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.5 (2.2) 9.7 (6.5) 4.8 (2.0) 85.9 (1.8) 10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 5.8 (5.9) 31.0 (8.1) 25.3 (11.3) 94.9 (2.3)
+It consists of 3,344 clean images, 9,790 trigger images of 7 physical triggers, and 8,104 additional images to evaluate accidental backdoor activations (see Section 7).
+Our animal dataset is created by combining a public Kaggle dataset [1], containing 45 classes of mammals, with a new cat class containing 500 cat images.
+For this dataset, we collected 1,500 cat images with 3 physical triggers (tennis ball, phone, book) to study CLPBA.
+The images are captured in various indoor and outdoor settings, under varying weather conditions, and with diverse shooting angles and distances.
+Details of our dataset can be found in the Appendix B.
+Table 1: ASR (%) of poisoning algorithms on 5 random source-target pairs, averaged over 10 runs with standard deviation in brackets.
+Poison pair Poison ratio $\alpha$ Naive MP FM GM
+$\epsilon=8/225$ $\epsilon=16/255$ $\epsilon=8/225$ $\epsilon=16/255$ $\epsilon=8/225$ $\epsilon=16/255$
+(a) Facial Recognition (sunglasses trigger)
+1 – 7 5% 35.0 (18.9) 14.2 (2.6) 10.5 (3.6) 84.0 (3.9) 80.9 (7.4) 97.2 (0.5) 99.4 (0.5)
+10% 81.5 (3.5) 13.7 (2.3) 13.4 (2.6) 82.3 (0.5) 86.9 (2.5) 98.0 (0.5) 100.0 (0.0)
+2 – 0 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 85.0 (4.4)
+10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 99.4 (1.0)
+4 – 3 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.4 (1.2) 15.8 (16.4) 96.4 (1.2)
+10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 12.8 (11.6) 48.9 (16.2) 98.6 (0.3)
+6 – 8 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.7 (2.9) 18.2 (12.3) 7.2 (11.0) 97.8 (2.4)
+10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 4.7 (8.1) 22.9 (8.3) 21.8 (19.2) 98.3 (0.8)
+9 – 5 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.6 (0.5) 5.8 (3.2) 22.7 (10.9) 95.6 (1.8)
+10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 3.4 (0.5) 30.4 (1.5) 51.4 (17.8) 97.8 (0.2)
+(b) Animal Classification (tennis trigger)
+cat – fox 5% N/A 0.0 (0.0) 0.0 (0.0) 0.3 (0.5) 0.2 (0.3) 8.2 (4.4) 9.6 (4.1)
+10% N/A 0.0 (0.0) 0.0 (0.0) 1.0 (0.8) 0.2 (0.4) 20.5 (8.9) 17.8 (7.7)
+cat – kangaroo 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.9 (0.1) 8.4 (0.8)
+10% N/A 0.0 (0.0) 0.0 (0.0) 0.1 (0.1) 0.1 (0.1) 9.4 (2.1) 18.0 (2.0)
+cat – mongoose 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 2.7 (0.3) 5.3 (0.4)
+10% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 6.7 (0.1) 10.3 (2.9)
+cat – sea-lion 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 7.4 (0.6) 13.4 (2.9)
+10% N/A 0.0 (0.0) 0.0 (0.0) 0.3 (0.1) 0.0 (0.0) 22.1 (5.3) 23.6 (5.8)
+cat – tapir 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.8 (0.8) 7.8 (5.0)
+10% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 29.8 (8.6) 36.3 (7.6)
 4 Experiments
-We evaluate CLPBA using four Convolutional Neural Networks (CNNs), but only the results for ResNet50 [26] are presented in the main paper.
-The models are pre-trained on the VGGFace 2 dataset [27] and then fine-tuned on our dataset.
+For Facial recognition, we choose the victim model to be ResNet50 [18] pre-trained on the VGGFace2 dataset [4].
+For Animal classification, the victim model is ResNet18 pre-trained on Imagenet [43].
 We assume a gray-box scenario where the attacker knows the architecture of the victim model but does not have access to its training process.
-Each attack has a unique set of parameters, so we follow as closely as possible to the experimental settings of the original works.
-We measure the performance of an attack, or attack success rate (%) (ASR) success rate, by the percentage of trigger samples in the source class that are misclassified as the target class.
-Note that the prediction accuracy (ACC) on clean samples consistently exceeds 99% and is therefore not reported in most experiments.
-Detailed experimental settings can be found in Appendix D.
-Additional experiments on other models and different settings are provided in Appendix E.
+To enable a fair comparison across different poisoning algorithms, we standardize the poisoning hyperparameters, including the number of iterations, optimizer, batch size, learning rate, and other relevant parameters.
+The performance of attacks is measured by attack success rate (%) (ASR), the percentage of trigger source-class samples that are misclassified as the target class.
+Note that the prediction accuracy (ACC) on clean samples consistently exceeds 99% for facial recognition and 93% for animal classification across different attacks, so it is not reported in the paper.
+Detailed experimental configurations can be found in Appendix C.
+Appendix D provides additional experiments on other architectures, hyperparameters, and physical triggers.
 4.1 Comparison of Poisoning Algorithms
 The results from Table 1 show that GM is the most effective poisoning algorithm.
-BLC and Naive methods fail with small poison ratios.
-Higher values of $\epsilon$ and $\alpha$ could increase ASR, but $\epsilon$ has a greater impact as it allows the poisoning objective to converge to smaller values.
-Variance of experiment results between poisoning attempts could be due to different weight initialization of the victim model.
-  • Why does GM work?
-By aligning the gradients of poisoned samples with the gradient of adversarial loss, we observe that the perturbations of GM appear to reconstruct deep representations of $\widetilde{\mathcal{D}}_{s}$.
-These perturbations replicate the facial contours of the source class and the appearance of the sunglasses trigger.
-This enables the model to extract key features of $\widetilde{\mathcal{D}}_{s}$ via the perturbations.
-  • Why does FM fail?
-FM was initially proposed for transfer-learning models, assuming the weights of previous layers remain unchanged.
-In the end-to-end scenario, when the victim model is trained on the poisoned dataset, its shallower convolution kernels will learn the perturbations as features of the target class, making poisoned instances stay in the target-class distribution in deeper layers.
-  • Why does BLC fail?
-First, the use of a low poison ratio and a subtle digital trigger (opacity 32/255) fails to establish a strong association between the trigger and the target class so the presence of the trigger at inference time cannot dominate the salient features in the source class.
-Second, the digital trigger may lack the variability introduced by physical conditions such as brightness changes, color shifts, noise, and viewpoint distortions, making it an unsuitable replacement for the physical trigger.
-Among the three poisoning algorithms, FM produces the finest and most naturally dispersed perturbations (Appendix F.3); in contrast, BLC’s perturbations are most noticeable with irregular color patches all over the images.
-Interestingly, GM’s perturbations appear to reconstruct the face shape of the source class and the appearance of the triggers (Figure 3).
-In a small number of FM’s poisoned samples, the perturbations exhibit similar patterns to GM (e.g., face shape and textures of sunglasses).
-We believe that such samples are critical samples for the attack success because they hold the trigger information and the robust features of the source class.
-GM is more effective than FM since it can craft a greater number of critical samples.
-4.2 Impact of poison pair on attack performance
-In Table  1, the attack effectiveness varies significantly across source-target class pairs.
-Poison pairs $1-7$ and $3-6$ show much higher ASR than others, even for Naive and BLC poisoning methods.
-This section examines the impact of source and target classes on attack performance.
-From the source class.
-A source class is more susceptible to physical attacks at inference if the trigger obscures the class’s discriminative classification regions, which enables the trigger features to outweigh salient features and cause misclassification.
-In Table 2, the classification loss of the clean ResNet50 model on inference-time trigger samples is much higher for the source class 3 than for the source class 4, and thus the attack on class 3 has a higher ASR.
-In Figure 8 of Appendix B, we use Grad-CAM [28] to visualize discriminative regions implicitly detected by a clean ResNet50 model (i.e., model trained solely on clean data).
-We can see that in trigger instances of source class 3, the model struggles to extract discriminative features and must rely on less prominent features for classification.
-If the model is backdoored, then the model will focus on the trigger for classification.
-Table 2: Performance of GM attack on source class 3 and source class 4, for a fixed target class 6. Clean model loss is measured over inference-time trigger samples of the source class.
-Poison pair Clean model loss Poisoned model ASR (%)
+Higher values of $\epsilon$ and $\alpha$ could increase ASR, but $\epsilon$ has a greater impact as it allows the poisoning objective to converge to smaller values and the poisoned samples to contain more adversarial features of the source-class trigger distribution.
+Additionally, CLPBA is more effective for facial recognition than animal classification, probably because the ”sunglasses” is a very dominant feature that hides important classification features (i.e., the eyes and the nose) for face recognition, while the ”tennis ball” trigger is a relatively minor feature so it cannot outweigh the natural features of the cat.
+Dataset bias.
+Interestingly, we can observe that the poison pair 1-7 in Table 1a has an unusually high ASR compared to other pairs.
+Even with no trigger images injected, ASR reaches 26.21 % for the Naive method.
+We believe this is because certain samples in the target class contain similar features to the physical trigger, reinforcing the trigger as a dominant feature once trigger images are added.
+In this case, it may be the eyeglasses of the target-class that have a similar shape and texture to the sunglasses.
+While this correlation may be accidental due to our dataset collection and trained classifier, it still raises an important issue of dataset bias that a backdoor attacker can exploit to select an effective trigger aligning with the natural features of the target class.
+In Appendix A, we show that the removal of training samples with strong trigger features drastically reduces the ASR and thus prevents the dataset bias.
+4.2 Analysis of CLPBA
+Two unique attributes of CLPBA are that it is hidden-trigger and source-specific.
+The first attribute means the trigger is not present in the poisoned samples, while the second ensures only source-class samples can cause the model to misclassify with the trigger.
+CLPBA achieves these attributes by embedding features of the source-class trigger distribution, $\widetilde{\mathcal{D}}_{s}$, into the perturbations in the poisoned target samples.
+Specifically, the embedded features consist of adversarial features and trigger features.
+We define adversarial features as the non-robust features [25] that shift the representations of target instances toward the source-class distribution in the feature space, and trigger features as the high-level features of the physical trigger extracted by the neural network.
+The existence of adversarial features ensures the source-specific attribute, while reconstructed trigger features allow the model to indirectly learn the physical trigger as a feature of the target class.
+Figure 3: Experiment to show the existence of adversarial features. We set a high $\epsilon$ of 32, $\alpha=2$, poison pair is $9-5$.
+Existence of adversarial features.
+To justify the existence of adversarial features, we take inspiration from [25] to test whether the model trained on poisoned target samples could successfully predict source-class samples.
+We design the experiment as illustrated in Figure 3.
+First, we craft the GM-poisoned samples with the original dataset and the attacker model.
+Then, we filter out all source-class samples, before relabelling all GM-poisoned samples to the source class.
+We then train the victim model from scratch on this modified dataset before evaluating the recall (true positive rate) of predicting source-class samples that have been filtered previously.
+To avoid randomness, we set a prediction threshold of 85% after the softmax function.
+Notably, the model trained on this modified dataset has a prediction accuracy of 75% on source-class samples.
+This shows that the perturbations on poisoned target instances contain adversarial features representative of source-class distribution.
+To make sure that the trained model is not biased towards predicting out-of-distribution samples as the source class, we also evaluate the false positive rate of classifying 200 random face images from the CelebA dataset [35] as the source class, and the accuracy is 0%.
+Existence of trigger features.
+By scaling the perturbations on GM-poisoned samples, we observe that the perturbations appear to reconstruct the shape of the physical trigger.
+Interestingly, setting a high value of $\epsilon$ even reconstructs the samples in $\widetilde{\mathcal{D}}_{s}$, as can be observed in Figure 5.
+(a) Clean model
+(b) Poisoned model
+Figure 4: Feature space visualization.
+Feature space analysis.
+In Figure 4, we project source-class samples (green), target-class samples (blue), trigger source samples (yellow), and poisoned target samples (red) in 2-dimension feature space of both clean and poisoned models where the first axis is $\mu_{s}-\mu_{t}$ and the second axis is $\theta_{s}-\theta_{t}$, the normal vector to the decision boundary between the source and target class.
+In both feature spaces, we can see that some poisoned target instances are close to the samples of $\widetilde{\mathcal{D}}_{s}$.
+For the clean model, $\widetilde{\mathcal{D}}_{s}$ overlaps with ${\mathcal{D}}_{s}$ in the feature space.
+For the poisoned model, however, $\widetilde{\mathcal{D}}_{s}$ is shifted towards $\widetilde{\mathcal{D}}_{t}$ in the feature space.
+From this observation, we can say that since poisoned target instances contain representative features of $\widetilde{\mathcal{D}}_{s}$, especially the trigger feature which is not a part of ${\mathcal{D}}_{s}$, training the model on these poisoned instances cause $\widetilde{\mathcal{D}}_{s}$ to move closer to the classification region of the target class in the feature space.
+Figure 5: From left to right: tennis, fake-beard, and sunglasses triggers. First row: a sample in $\widetilde{D}_{s}$. Second row: GM perturbation.
+4.3 Defenses
+We evaluate the performance of CLPBA against representative filtering defenses (filter poisoned samples during training), and firewall defenses (filter trigger samples during inference).
+For the filtering defenses, we evaluate Elimination Rate (ER), the percentage of poisoned samples that are correctly filtered, and Sacrifice Rate (SR), the percentage of clean samples that are incorrectly filtered.
+For firewall defenses, we evaluate the True Positive Rate (TPR), the percentage of trigger source-class samples that are correctly filtered, and the False Positive Rate (FPR), the percentage of non-trigger samples that are incorrectly filtered.
+For details on these defenses, we refer readers to Appendix F.
+Table 2: Defense evaluation on GM attack ($\alpha=0.1$, $\epsilon=16/226$, poison pair $7-0$). ASR is 99.7% without defenses.
+Filtering Defenses
+Metrics AC[3] SS[52] SPECTRE[17] SCAN[51]
+ER (%) 62.16 97.30 72.97 0
+SR (%) 15.77 19.77 19.50 0
+ASR (%) 99.7 0.0 99.7 99.7
+Firewall Defenses
+Metrics STRIP[14] CD[20] IBD-PSC[19] Frequency[59]
+TPR (%) 0.0 79.6 0.0 0.0
+FPR (%) 6.7 65.3 16.7 0.9
+For filtering defenses, we evaluate Spectral Signature (SS) [52], Activation Clustering (AC) [3], SPECTRE [17], and SCAN [51].
+Table 2 shows that only SS is effective against CLPBA.
+Although we can see that AC and SPECTRE have high elimination rates (62.16% and 72.97%), the attack still achieves 99.7 % ASR.
+This shows that the GM poisoning objective (Equation 8) focuses on a small subset of effective poison instances to match the gradient of adversarial loss and distill the features from the trigger distribution.
+This is supported in Figure 4 where only a subset of poisoned samples overlap with the trigger distribution.
+Detecting such effective instances is generally difficult since they correlate strongly with each other and they transform the model’s decision boundary of the target class.
+Table 2 also shows that firewall defenses are not effective for CLPBAs because they either miss trigger samples (low TPR) or incorrectly filter out too many benign samples (high FPR).
+This behavior is expected since these defenses are built on the memorization property of backdoor attacks, which differs from the generalization property of CLPBA.
+5 Improving the Effectiveness of CLPBA
+Empirically, higher $\epsilon$ lead to a higher ASR since the poisoning objective can converge to a smaller value.
+However, the perceptual quality of poisoned samples is compromised since the poisoned images have high-frequency artifacts that are clearly visible (Figure 5).
+The drawbacks of the $\ell_{\infty}$ regularization on perturbations are that it is non-differentiable, and the projection on $\ell_{\infty}$-norm ball is only executed at the end of an optimization step, making backdoor and poison stealthiness two separate tasks.
+As a consequence, the perceptual quality of poisoned images is compromised because (1) the pixel values of perturbation are often pushed to the boundary of the norm ball, (2) the noise can appear in unnecessary regions of the image, and (3) the perturbations contain high-frequency artifacts.
+To improve the visual quality and convergence of the poisoning objective, we propose to replace the current $\ell_{\infty}$ regularization with a Pixel regularization containing a soft pixel-wise $\ell_{\infty}$ loss and discretized Total Variation Upwind (TVU) [6]:
+$\mathcal{R}_{pix}(\delta)$ $=||\delta^{+}-M\odot\epsilon||_{1}+\alpha TVU(\delta;\beta=1)$ (10)
+$TVU(\delta;\beta)$ $=\sum_{p}^{\text{all pixels}}\biggl{(}\sum_{q\in\mathcal{N}(p)}\big{(}\delta^{(p)}-\delta^{(q)}\big{)}^{2}\biggr{)}^{\frac{\beta}{2}}$ (11)
+where $\delta^{+}$ takes the absolute value of every element in $\delta$, $M$ is a mask of pixel-wise embedding ratio: $M_{i,j,k}=\frac{x_{i,j,k}}{\max_{k}{x_{i,j,k}}}$, assuming $x_{i,j,k}\in[0,255]$.
+$M$ encourages the perturbation to align with the underlying image content.
+In Equation 11, $\mathcal{N}(p)$ are neighbouring pixels of pixel $p$.
+The weighting factor $\alpha$ is set to 5.
+Both terms in Equation 10 improve the smoothness of perturbations, with $TVU(\delta)$ minimizing the difference between adjacent pixels.
+The side-by-side comparison between standard $\ell_{\infty}$ regularization and our pixel regularization is given in Figure 6.
+Our pixel regularization leads to a better convergence of the poisoning objective (Equation 8) while simultaneously reducing the $\ell_{2}$ norm of the perturbations to a smaller value, indicating an improved perceptual quality.
+Consequently, the GM attack with pixel regularization achieves an ASR of $97.8(1.1)$, compared to $51.9(18.7)$ for the original GM attack under the same settings.
+We also evaluate the perceptual quality of the poisoned samples statistically using Peak Signal-to-Noise Ratio (PSNR) [24], following previous works [29, 31].
+Our pixel regularization results in higher values of PSNR (36.9dB vs 29.8dB), corresponding to a better perceptual quality.
+(a) Standard $\ell_{\infty}$ regularization
+(b) Pixel Regularization
+(c) Poisoning loss
+(d) Perturbation norm
+Figure 6: Comparison between standard $\ell_{\infty}$-regularization and our Pixel Regularization.
+Top row: visual comparison; Bottom row: convergence rate.
+Poisoned samples are crafted with GM attack ($\alpha=0.1$, $\epsilon=8/255$) on poison pair 77773333.
+6 Analysis of Attack Factors
+The results in Table 1 show that ASR varies between different poison pairs for a given trigger, while the results in Table 4 show that different triggers may have different performances for a given poison pair.
+This section examines the impact of poison pairs and the physical trigger on CLPBA.
+Poison pair Clean model loss Backdoored model ASR (%)
 3-6 0.6855 99.0 (0.8)
 4-6 0.0038 86.8 (2.1)
-From the target class.
-A target class can be more susceptible to CLPBA due to the similarity between the class’s natural features and the trigger features.
-For example, in Appendix B, we show that the eyeglasses in some samples of the target class could represent weaker backdoor features for the attacker’s trigger.
-The existence of such support samples helps enforce the attacker’s trigger as a dominant classification feature of the target class once the model is poisoned.
-Filtering these support samples greatly reduces the victim model’s susceptibility to the backdoor trigger (see Table 6, Appendix B).
-This naturally implies the existence of natural physical backdoors in a dataset, especially if unique physical objects exist in target class data.
-A detailed analysis of the natural backdoors is given in Appendix B.
-We note that there is also an unexplored relationship between the distribution of samples in the target class $\mathcal{D}_{t}$ and the distribution of trigger samples in the source class $\widetilde{\mathcal{D}}_{s}$ that allows for feature alignment of poisoned samples in high dimensional space.
-4.3 Trigger selection
-The performance of a backdoor attack has been shown to depend on the size, shape, and location of a trigger [10, 29].
-The larger and closer to the face the trigger is, the more effectively it can obscure a person’s facial features and become dominant features.
-For a physical backdoor attack targeted at a facial recognition system, however, it is not practical to use large triggers like face masks and sunglasses, as people may be requested to take off such items when they go through facial recognition processes.
-Thus, it is important to study the performance of CLPBA using more stealthy triggers.
-Table 3: Performance of backdoor attacks using various physical triggers.
-Poisoned samples are crafted with $\epsilon=8$ and $\alpha=0.1$.
-Triggers are arranged from left to right based on their stealthiness.
+Table 3: Performance of GM attack on source class 3 and source class 4, for a fixed target class 6. Clean model loss is measured over inference-time trigger samples of the source class.
+Source class.
+A source class is more susceptible to physical attacks at inference if the trigger obscures the class’s discriminative classification regions, enabling the trigger features to outweigh salient features and cause misclassification.
+In Table 3, the classification loss of the clean ResNet50 model on inference-time trigger samples is much higher for source class 3 than for source class 4, resulting in a higher ASR for class 3.
+In Appendix A, we use Grad-CAM [47] to visualize discriminative regions detected by a clean and poisoned ResNet50 model (i.e., trained on clean and poisoned data).
+We observe that in trigger instances of source class 3, the clean model struggles to extract meaningful facial features and relies on less prominent features for classification.
+Conversely, the poisoned model focuses on the physical trigger for classification.
 Poison pair N/A Earrings Fake Beard Headband Hat Sticker Sunglasses Face Mask
@@
 FM 0.0 (0.0) 90.3 (7.0) 0.0 (0.0) 0.0 (0.0) 14.8 (0.2) 5.8 (5.9) 20.1 (5.4)
-We evaluate GM and FM attacks on poison pairs $7-3$ and $9-6$ using triggers of different sizes and levels of stealthiness (see Appendix C.2 for details into our trigger selection).
+Table 4: Performance of GM attack using various physical triggers. Poisoned samples are crafted with $\epsilon=8$ and $\alpha=0.1$.
+Target class.
+A target class can be more susceptible to CLPBA due to the similarity between the class’s natural features and the trigger features.
+For example, in Appendix A, we show that the eyeglasses in some samples of the target class could represent weaker backdoor features for the attacker’s trigger.
+We note that there is also an unexplored relationship between the distribution of samples in the target class $\mathcal{D}_{t}$ and the distribution of trigger samples in the source class $\widetilde{\mathcal{D}}_{s}$ that allows for feature alignment of poisoned samples in high dimensional space.
+Physical trigger.
+We evaluate GM and FM attacks on poison pairs $7-3$ and $9-5$ using triggers of different sizes and levels of stealthiness (see Appendix B for details on our trigger selection).
 We observe that the effectiveness of a trigger depends on the poison pair.
-Surprisingly, in poison pair $9-5$, the face mask, a large trigger, is not as effective as the "fake beard" trigger, a small and stealthy trigger.
-Factors that impact trigger effectiveness may include (1) whether the trigger aligns with the natural features of the target class and (2) whether the trigger hides important features of the source class.
-A natural question for the attacker is "how to select an effective and stealthy trigger for a given poison pair?" The attacker can brute-force search for an effective trigger by iteratively crafting poisoned samples and evaluating the ASR for each of the triggers he can use.
-However, this could be too computationally and time-expensive if the dataset size or the number of triggers is large.
-Here, we propose a simple heuristic for choosing an effective trigger.
-Given a set of $k$ triggers $T=\{p_{1},p_{2},\ldots p_{k}\}$ corresponding to $\{\widetilde{D}_{s}^{(1)},\widetilde{D}_{s}^{(2)},\ldots\widetilde{D}_{s}^{(k)}\}$, in which $\widetilde{D}_{s}^{(i)}$ is the set of attacker’s source class samples with trigger $p_{i}$.
-With a model $F_{\theta}$ that has been trained only on clean data, the attacker can select a trigger for a poison pair $(s,t)$ with the following formula:
-$p=\operatorname*{arg\,min}_{i}\frac{1}{|\widetilde{D}_{i}|}\sum_{x\in\widetilde{D}_{s}^{(i)}}2Softmax(F(x))_{t}-Softmax(F(x))_{s}$ (8)
-where $Softmax(F(x))_{t}$ and $Softmax(F(x))_{s}$ represent the prediction probabilities of $x$ being in classes $t$ and $s$.
-Intuitively, Equation 8 is designed to identify a trigger that aligns with the natural features of the target class while obscuring key features of the source class.
-The constant in the first term is to bias our selection toward the first factor; $Softmax$ function is used to normalize the score for each sample.
-See Appendix H for an evaluation of our trigger selection heuristics.
-4.4 Enhancing imperceptibility of attacks
-In the pixel space.
-Higher values of $\epsilon$ result in higher ASR at the expense of visual stealthiness as the perturbations could leave behind visible artifacts on high-resolution facial images (Figure 3(a)).
-In the current $\ell_{\infty}$ regularisation, the projection of perturbations on the norm ball is only executed after every optimization step, making backdoor and visual imperceptibility two separate tasks.
-As a result, pixel values of the perturbations are often pushed to the boundary of the norm ball, causing noise to appear in unnecessary regions.
-Furthermore, the perturbations are non-smooth and irregular because each pixel is optimized individually.
-To improve the perceptual quality of poisoned samples, we propose to replace the current $\ell_{\infty}$ regularisation with a Pixel Regularization (PR) containing soft pixel-wise $\ell_{\infty}$ loss and Total Variation Upwind (TVU) [30]:
-$\mathcal{R}_{pix}(\delta)$ $=||\delta-M\odot\epsilon||_{1}+TVU(\delta;\beta=1)$ (9)
-$TVU(\delta;\beta)$ $=\sum_{p}^{\text{all pixels}}\biggl{(}\sum_{q\in\mathcal{N}(p)}\big{(}\delta^{(p)}-\delta^{(q)}\big{)}^{2}\biggr{)}^{\frac{\beta}{2}}$ (10)
-where $M$ is a pixel-wise embedding mask based on the original image: $M_{i,j,k}=1.0-x_{i,j,k}$, assuming the pixel value $x_{i,j,k}\in[0,1]$.
-In Equation 10, $\mathcal{N}(p)$ are neighbouring pixels of pixel $p$.
-Both terms in Equation 9 improve the smoothness of perturbations, with $TVU(\delta)$ minimizing the difference between adjacent pixels and $M$ keeping the perturbation within the color range.
-As shown in Figure 3, the perturbations crafted with PR are smoother and less conspicuous than those crafted with the original $\ell_{\infty}$ regularization.
-Following previous works [31, 32], we also evaluate the imperceptibility of poisoned samples statistically using Peak Signal-to-Noise Ratio (PSNR) [33].
-The average PSNR value of poisoned samples is reported in Table 4.4.
-Our regularization results in higher PSNR, corresponding to better perceptual quality.
-We note that an additional advantage of the soft $\ell_{\infty}$ regularization over the standard one is that it is differentiable, which enables better convergence for the poisoning objective, especially for lower values of $\epsilon$ (see Appendix E.3).
-(a) Standard $\ell_{\infty}$ regularization
-(b) Pixel Regularization
-Figure 3: Comparision between standard $\ell_{\infty}$ regularization with our Pixel Regularization.
-Perturbations are scaled 10×10$\times$10 × for visibility.
-The perturbations in both regularizations appear to reconstruct the facial shape of the source class with sunglasses as the trigger features.
-In the feature space.
-Backdoor poisoning attacks have been observed to leave discernible traces in the feature space of poisoned models [34, 35, 36, 37].
-By projecting feature representations of target samples (containing both clean and poisoned samples) on the first two principle axes (Figure 4(a)), we can see a clear distinction between the poisoned samples and clean samples.
-The separation exists because the perturbations on poisoned samples contain unnatural trigger features of $\widetilde{D}_{s}$, from which the model can learn as features of the target class.
-To limit the separation of poisoned samples from clean samples in the feature space, we propose a Feature Regularization (FR) that minimizes the distance of samples in $P$ to the distribution of original target-class samples $\mathcal{D}_{t}$: $\mathcal{R}_{feat}(\delta)=||f(x+\delta)-f(u)||_{2}^{2}$ where $u$ the empirical mean is computed from dataset $D_{t}$.
-In Figure 4, we can see that FR constrains the separation between poisoned and benign samples of the target class.
-In Table 4.4, by evaluating four representative filtering defenses (Spectral Signature (SS) [34], Activation Clustering (AC) [36], SPECTRE [35], DeepKNN [38]), we can confirm the effectiveness of FR in concealing the poisoned inputs from machine inspection in the feature space  [^2] Details of these defenses are outlined in Appendix G..
-(a) Without FR
-(b) With FR
-Figure 4: 2D-PCA view of target-class representation produced by the poisoned model. FR constrains the separation between poisoned samples and benign samples.
-Table 4: Defenses on GM attack ($\alpha=0.1,\epsilon=16/226$, poison pair $7-0$). ASR (%) is reported. PSNR (dB) is the metric to measure visual quality.
-GM
-GM + PR
-GM + FR + PR
-No defense 88.0 93.2 89.5
-AC 0.0 0.9 88.9
-SS 0.0 1.9 53.2
-SPECTRE 0.0 99.1 87.0
-DeepKNN 75.9 98.1 88.0
-PSNR (dB) $\uparrow$ 29.8 dB 36.9 dB 37.2 dB
-Altogether, the final form of poisoning objective is
-$\min_{\delta}\mathcal{A}+\lambda_{1}\mathcal{R}_{pix}+\lambda_{2}\mathcal{R}_{feat}$ (11)
-where $\mathcal{A}$ is the main poisoning objective, and $\lambda_{1},\lambda_{2}$ are balancing terms for the regularizations.
-In our experiments, we set $\lambda_{1}=\frac{10}{n}$ ($n$ is the total number of pixels), and $\lambda_{1}=0.05$.
-We note that $R_{feat}$ can degrade attack performance by constraining poison optimization, so only a small weight should be assigned to this regularization to mitigate the impact of filtering defenses.
+Surprisingly, in poison pair $9-5$, the ”face mask”, a large trigger, is not as effective as the ”fake beard” trigger, a small and stealthy trigger.
+From our previous analysis, there are two factors impacting trigger effectiveness: whether the trigger aligns with the natural features of the target class, and whether the trigger hides important features of the source class.
 7 Accidental Backdoor Activations
 For CLPBA to be stealthy, we argue that it needs to be source-specific and trigger-specific.
-That is, (1) the trigger should be discriminative so that the victim model does not incorrectly recognize similar-looking objects as triggers, and (2) only the source class can use the trigger to manipulate the victim model’s prediction at inference time.
-We introduce two new metrics, False Positive Rate and Suspicion Rate, to test the specificity (and also stealthiness) of CLPBA:
+That is, (1) only the source class can use the trigger to manipulate the victim model’s prediction at inference time; (2) the trigger should be discriminative so that the victim model does not incorrectly recognize similar-looking objects as triggers.
+We introduce two new metrics, False Positive Rate, and Suspicion Rate, to test accidental backdoor activations:
   • Suspicion rate (SR): The measurement of whether the victim model incorrectly misclassifies trigger samples from non-source classes as belonging to the target class.
   • False positive rate (FPR): The measurement of whether samples from the source class can fool the victim model with false triggers (i.e., objects that resemble the attacker’s physical trigger).
-Table 5: Measurement of FPR and SR for GM attack ($\alpha=0.1,\epsilon=16/226$).
-A detailed set-up of our dataset to evaluate SR and FPR is given in Appendix C.
-We note that the big difference in FPR between fake beard and sunglasses triggers is due to selection bias in our choice of false triggers.
-Poison pair Trigger Loss type ASR (%) ACC (%) FPR $\downarrow$ (%) SR $\downarrow$ (%)
-9-5 Fake beard Original 98.7 (0.0) 99.9 (0.1) 50.8 (1.9) 12.5 (1.0)
-With repel 98.7 (0.0) 99.9 (0.1) 49.7 (2.2) 3.9 (0.1)
-7-3 Sunglasses Original 97.5 (1.1) 99.3 (0.1) 0.7 (0.8) 13.9 (0.7)
-With repel 97.5 (1.4) 99.4 (0.2) 0.9 (0.1) 2.5 (0.3)
+Poison pair Trigger Loss type ASR ACC FPR SR
+9-5 Fake beard Original 98.7 99.9 50.8 12.5
+With repel 98.7 99.9 49.7 3.9
+7-3 Sunglasses Original 97.5 99.3 0.7 13.9
+With repel 97.5 99.4 0.9 2.5
+Table 5: FPR and SR for GM attack. The set-up of our dataset to evaluate SR and FPR is given in Appendix B.
 Between the two metrics, SR is more important because it directly impacts attack stealthiness if a benign client is misclassified as the target class with the trigger or similar-looking items.
@@
 However, the attacker can partially mitigate the issue of false source classes, as measured by SR, by adding repel terms to the adversarial loss $L_{adv}$ in Equation 2:
-$L_{adv}=N\cdot\overbrace{\sum_{x\in\widetilde{D}_{s}}\mathcal{L}\bigl{(}F_{\theta}(x),y_{t}\bigr{)}}^{\text{adversarial loss}}+\sum_{i\neq s}^{N}\overbrace{\Bigl{(}\sum_{x\in\widetilde{D}_{i}}\mathcal{L}\bigl{(}F_{\theta}(x),y_{i}\bigr{)}-\mathcal{L}\bigl{(}F_{\theta}(x),y_{t}\bigr{)}\Bigr{)}}^{\text{repel term}}$ (12)
+$L_{adv}=N\cdot\overbrace{\sum_{x\in\widetilde{D}_{s}}\mathcal{L}\bigl{(}F_{\theta}(x),y_{t}\bigr{)}}^{\text{adversarial loss}}$
+[…]
 Here, $N$ is the total number of classes, and $\widetilde{D}_{i}$ is the dataset of trigger instances in class $i$.
@@
 8 Conclusion
-Through extensive experiments, our research underscores the vulnerability of DNNs to clean-label backdoor attacks in the physical world.
-We thoroughly analyze the influence of poisoning algorithms, physical triggers, and source-target class pairs on the effectiveness of these attacks.
-Additionally, we propose novel pixel and feature regularizations to improve the imperceptibility of attacks in both pixel and feature spaces.
-A significant limitation identified is accidental backdoor activations of the poisoned model due to false triggers or non-source classes, which could compromise attack stealthiness.
+Through extensive experiments, our research underscores the vulnerability of DNNs to clean-label physical backdoor attacks.
+We then show that CLPBA works by embedding adversarial features and trigger features in the poisoned instances.
+Additionally, we propose a novel pixel regularization to improve the effectiveness and stealthiness of poisoned samples.
+A significant limitation identified is accidental backdoor activations of the backdoored model due to false triggers or non-source classes, which could compromise attack stealthiness.
 To address this, we propose incorporating repel terms into the adversarial objective to enhance trigger specificity to the source class.
-Our findings highlight the critical need for robust defenses against clean-label physical backdoor attacks and pave the way for future research on improving the resilience of DNNs against such threats.
-Appendix A Procedure of blended label-consistent attack
-Figure 5: Procedure of BLC with $\tau=1$.
-We first add adversarial perturbations to the poisoned instance.
-Then, using facial landmarks, we map the trigger to the appropriate facial area with perspective transformation.
-Appendix B Natural physical backdoors in a dataset
+Our findings highlight the critical need for robust defenses against CLPBA.
+By releasing the dataset and code, we hope to facilitate future research on improving the resilience of DNNs against such threats.
+Supplementary Material
+Appendix A Natural Physical Backdoors
 For a backdoored DNN, the backdoor trigger can be understood as the strong feature for classification of the target class [41, 26, 52].
@@
 This means that when the model overfits to the sunglasses feature of the target class dataset, inference-time samples from $\widetilde{\mathcal{D}}_{s}$ will be mapped to a region closer to $\widetilde{\mathcal{D}}_{t}$ than $\mathcal{D}_{s}$ in the feature space (Figure 8), resulting in misclassification.
-(a) Clean samples (without trigger)
+(a) Clean samples (w/o trigger)
 (b) Trigger samples
@@
 […]
-Appendix C Dataset
-C.1 Ethics and data protection
+Appendix B Dataset
+B.1 Ethics & Data Protection
 IRB approval.
-Before conducting our study, we submitted a "Human Subjects Research – Expedited Review Form" to our country’s Institutional Review Board (IRB).
+Before conducting our study, we submitted a ”Human Subjects Research – Expedited Review Form” to our country’s Institutional Review Board (IRB).
 Our study received approval from the chairman of the institutional ethical review board under decisions No 24/2016/QD-VINMEC, No 23/2016/QD-VINMEC, and No 77/2021/QD-VINMEC.
 We prepared a consent form beforehand to ensure transparency in the procedure of dataset collection.
-All 10 volunteers in our dataset provided explicit written consent for us to take their photos.
+All 10 volunteers in our dataset provided explicit written consent for us to collect the dataset and use the images for research purposes, including permission to use the captured images in the research paper.
 Dataset protection.
@@
 Additionally, all images in the paper are partially obscured to ensure that the identities of our volunteers are not exposed.
-C.2 Dataset collection
+B.2 Dataset Collection
 Due to the lack of publicly available datasets to study physical backdoor attacks, we collect a facial recognition dataset following the design consisting of approximately 21,238 images of 10 volunteers (9 males, 1 female, all Asian) and 7 triggers.
@@
 With these criteria, we select 7 triggers with evaluations given in Table 7.
-C.3 Dataset preprocessing
+B.3 Dataset Preprocessing
 After collecting the images, we utilize a pre-trained MTCNN [61] (Multi-task Cascaded Convolutional Networks) model to detect and crop the face area.
@@
 Figure 10: Visualization of the face detection process. Left: Original image. Right: Processed image with face area cropped and resized to $224\times 224$ pixels.
-C.4 Dataset set-up
+B.4 Dataset Structure
 The train and test data are trigger-free images (the participants do not wear any trigger items) of 10 classes corresponding to 10 volunteers. The train and test data are split in a $80:20$ ratio.
-Figure 10: Example images of the 10 volunteers, representing the 10 classes in the dataset.
 False triggers.
-To evaluate the sensitivity of poisoned models to false triggers, as measured by false positive rate (FPR) (see Appendix D.4), we select three additional false triggers for each of the main triggers.
+To evaluate the sensitivity of poisoned models to false triggers, as measured by false positive rate (FPR) (see Appendix C.4), we select three additional false triggers for each of the main triggers.
 Because it is hard to find a representative set of false triggers for each of the main triggers, we select the false triggers by changing the sizes, colors, and shapes of the objects (e.g, we change the shape of earrings and hats, change the colors of the headband, change the size of sticker).
@@
 Therefore, a higher FPR of a trigger does not necessarily mean that that trigger is less specific and stealthy than other triggers.
-Figure 11: Left column: Main trigger. Right columns: Three false triggers with matching shapes, sizes, colors, and locations on the face to test the false positive rate of the model.
-Appendix D Experiment settings
-D.1 Computational resources.
+B.5 Animal Classification Dataset
+Besides facial recognition, we also conduct an extended evaluation of CLPBA on animal classification.
+We collected 2000 cat images with 500 trigger-free images and 500 images for each of the triggers, which are tennis balls, mobile phones, and books.
+The trigger-free cat images are resized to $224\times 224$ and then concatenated to an existing animal classification dataset on Kaggle.
+Visualization of trigger images for this dataset is given in Figure 16  [^1] https://www.kaggle.com/datasets/asaniczka/mammals-image-classification-dataset-45-animals.
+Appendix C Experiment Settings
+C.1 Computational Resources.
 All the experiments are conducted on the two servers.
@@
 We develop the framework based on the previous works [49, 15].
-D.2 Training settings.
+We use Pytorch as the framework for our model training.
+C.2 Training Settings.
 We evaluate the attacks on four Convolutional Neural Networks for facial recognition: ResNet50 [18], VGG16 [48], MobileNetV4 [45], Inception-ResNet [50].
@@
 We only standardize the training data using z-score normalization with the mean and standard variation calculated from the data.
-We observe that the training models converge very quickly and achieve the validation accuracy of over $99\%$ by epoch 10.
-Table 8: Hyperapaters
+We observe that the training models converge very quickly and achieve the validation accuracy of over $99\%$ and $93\%$ by epoch 10 for Facial Recognition and Animal Classification tasks.
+Table 8: Training Hyperparameters
 Hyperparameters Value
@@
 Learning rate scheduler Decrease lr by $90\%$ at epoch 14, 25, 36
-D.3 Attack settings.
-Our experiments follow the gray-box scenario in which the attacker model has the same architecture and is trained on the same dataset as the victim model but has different weight initialization, and the attacker cannot access and modify the weights and training procedure of the victim model.
-Since different poisoning algorithms have different attack hyperparameters (attack iterations, learning rate schedulers, poisoning optimizers), we closely follow the settings from the original papers for each of the attacks in our experiments.
-We evaluate the attacks with low poison ratios: 5% and 10% of the target-class dataset, equivalent to approximately 0.5% and 1% of the whole dataset.
-By default, the trigger used in our experiments is sunglasses.
-Table 9: Hyperparameters for Gradient Matching attack
+Table 9: Hyperparameters for Poisoning Algorithms
 Hyperparameters Value
@@
 rate decreases to 0.0001 at the last epoch
+Trigger opacity (BLC) $32/225$
+Inner SGD steps (MP) $2$
+C.3 Attack Settings.
+Our experiments follow the gray-box scenario in which the attacker model has the same architecture and is trained on the same dataset as the victim model but has different weight initialization, and the attacker cannot access and modify the weights and training procedure of the victim model.
+The dataset of trigger samples is split 50-50 into train and test sets.
+The train set is used for crafting poisons while the test set is used to validate the attacks.
+Table 9 gives the hyperparameters for poisoning algorithms.
 In the original Gradient Matching attack, Souri et al.
 [49] proposed retraining the attacker model and selecting poison samples with maximum gradients to increase ASR.
-For a fair comparison between poisoning algorithms, however, we do not include such additional factors in our settings.
+For a fair comparison between poisoning algorithms, however, we do not include such factors in our settings.
 In our experiment, the attacker selects the poison samples randomly.
-Table 10: Hyperparameters for Feature Matching attack
-Hyperparameters Value
-Optimizer Projected Gradient Descend [47]
-Learning rate (lr) $0.01$
-Attack iterations $5000$
-Learning rate scheduler Decay the learning rate by 0.95 every 2000 iterations
-Table 11: Hyperparameters for Blended Label Consistent
-Hyperparameters Value
-Optimizer Projected Gradient Descend
-Learning rate (lr) $(\epsilon/4)/225$
-Attack iterations $100$
-Trigger opacity $32/225$
-Learning rate scheduler No scheduler
-D.4 Evaluation metrics.
+C.4 Evaluation Metrics.
 To evaluate the performance of CLPBA, we adopt two standard metrics for backdoor attacks:
@@
   • Accuracy (ACC): The model’s prediction accuracy on clean, ordinary test-time data when the model is poisoned.
-In some cases, CLPBA could be detected if (1) the trigger acts as universal trigger that causes targeted misclassification for any classes, and (2) the model incorrectly recognizes false triggers by unintended objects.
-Thus, we adopt two new metrics to evaluate the stealthiness of CLPBA:
-  • Suspicion Rate (SR): This measures the likelihood that the backdoor attack is unintentionally activated by unintended classes using the physical trigger.
-It is quantified by the percentage of samples from non-source and non-target classes that are misclassified as the target class when the main trigger is used.
-  • False Positive Rate (FPR): This measures the sensitivity of the poisoned model to false triggers.
-It is quantified by the percentage of samples from the source class that are misclassified as the target class when the false triggers are used.
-CLPBA is considered successful if it has a high ASR ($\geq 80\%$) with a negligible decrease ($\leq 5\%$) in ACC. The attack is stealthy if it has a low FPR and a low SR ($\leq 10\%$).
-Appendix E Additional experiments
-E.1 Experiments on other models
-Table 12: Performance of CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on VGG16
-Poison pair Naive BLC FM GM
-1-7 63.8 (4.4) 2.3 (0.5) 82.3 (12.5) 100.0 (0.0)
-3-6 98.6 (1.4) 95.7 (1.4) 100.0 (0.0) 100.0 (0.0)
-5-1 42.0 (14.2) 0.3 (0.5) 25.9 (15.8) 86.7 (1.9)
-6-8 0.0 (0.0) 0.0 (0.0) 1.4 (1.7) 0.0 (0.0)
-7-3 4.0 (3.0) 0.0 (0.0) 89.1 (1.6) 85.2 (7.3)
-9-5 0.0 (0.0) 0.0 (0.0) 6.3 (4.0) 6.3 (4.0)
+Appendix D Additional Experiments
+D.1 Blended Label-Consistent Attack
+For completeness, in our evaluation, we also evaluate Blended Label-Consistent attack [53, 8] which works by perturbing the poisoned samples with adversarial noise to make the salient features of the samples harder to learn before injecting the digital trigger, forcing the model’s classification to depend on the trigger.
+Following [53], we first train an adversarial-trained model, before using the poisoning method in [53] to craft adversarial examples.
+Then we insert the digital trigger $p$ (e.g., digital sunglasses) via a perspective transformation $T$ that maps the trigger to the appropriate facial area using facial landmarks.
+The trigger is inserted with an opacity $\tau$.
+$x_{t}$ $\leftarrow x_{t}+\operatorname*{arg\,max}_{\|\delta\|_{\infty}\leq\epsilon}\mathcal{L}(F(x_{t},\theta),y_{t})$ (14)
+$x_{t}$ $\leftarrow(1-\tau)\cdot x_{t}+\tau\cdot T(p)$ (15)
+The procedure of BLC is illustrated in Figure 11.
+Compared to random placement and fixed center placement of the trigger, our adaptive trigger placement method with facial landmarks has a higher attack success rate as it preserves the spatial relationship between the trigger and the face.
+Figure 11: Procedure of BLC with $\tau=1$.
+We first add adversarial perturbations to the poisoned instance.
+Then, using facial landmarks, we map the trigger to the appropriate facial area with perspective transformation.
+D.2 Experiments on Different Models
+Table 10: CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on VGG16
+Poison pair Naive BLC MP FM GM
+1-7 63.8 (4.4) 2.3 (0.5) 0.3 (0.5) 82.3 (12.5) 100.0 (0.0)
+3-6 98.6 (1.4) 95.7 (1.4) 39.6 (9.9) 100.0 (0.0) 100.0 (0.0)
+5-1 42.0 (14.2) 0.3 (0.5) 0.9 (0.0) 25.9 (15.8) 86.7 (1.9)
+6-8 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.4 (1.7) 0.0 (0.0)
+7-3 4.0 (3.0) 0.0 (0.0) 0.0 (0.0) 89.1 (1.6) 85.2 (7.3)
+9-5 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 6.3 (4.0) 6.3 (4.0)
 ACC (%) of clean model 98.6 (0.3)
-Table 13: Performance of CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on MobileNetV2
-Poison pair Naive BLC FM GM
-1-7 81.8 (7.1) 3.1 (3.5) 16.0 (11.6) 50.1 (18.3)
-3-6 92.3 (6.0) 39.1 (9.1) 51.7 (8.5) 84.5 (5.0)
-5-1 87.0 (4.2) 19.4 (7.0) 36.4 (15.2) 43.5 (8.5)
-6-8 56.5 (1.3) 24.8 (9.3) 43.3 (1.7) 52.1 (1.7)
-7-3 8.3 (8.8) 0.0 (0.0) 0.3 (0.5) 0.0 (0.0)
-9-5 4.0 (1.7) 0.0 (0.0) 0.1 (0.2) 15.7 (3.4)
-ACC (%) of clean model 93.2 (0.2)
-Table 14: Performance of CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on Inception-ResNet
-Poison pair Naive BLC FM GM
-1-7 5.4 (7.9) 0.3 (0.5) 0.3 (0.5) 96.0 (3.2)
-3-6 61.4 (4.7) 48.8 (15.6) 80.2 (1.7) 100.0 (0.0)
-5-1 4.3 (1.9) 0.9 (0.0) 2.8 (1.9) 98.5 (1.9)
-6-8 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 96.4 (0.5)
-7-3 0.3 (0.5) 0.0 (0.0) 0.0 (0.0) 97.8 (2.1)
-9-5 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 88.9 (6.9)
+Table 11: CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on Inception-ResNet
+Poison pair Naive BLC MP FM GM
+1-7 5.4 (7.9) 0.3 (0.5) 13.7 (9.9) 0.3 (0.5) 96.0 (3.2)
+3-6 61.4 (4.7) 48.8 (15.6) 96.1 (2.2) 80.2 (1.7) 100.0 (0.0)
+5-1 4.3 (1.9) 0.9 (0.0) 0.0 (0.0) 2.8 (1.9) 98.5 (1.9)
+6-8 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 96.4 (0.5)
+7-3 0.3 (0.5) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 97.8 (2.1)
+9-5 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 88.9 (6.9)
 ACC (%) of clean model 99.8 (0.1)
-The results on MobileNetV2 in Table 13 are different from other models as the Naive attack has a very high ASR compared to other methods.
-The prediction ACC of MobileNetV2 is also lower with only 93.2%.
-We believe that MobileNetV2 is less robust than other models, which results in higher ASR with Naive attack.
-Furthermore, the attack performance on poison pairs can be model dependent as we can see in Table 12 that the pair $6-8$ and $9-5$ fails for VGG16, while poison pair $7-3$ fails for MobileNetV2 (Table 13).
-E.2 Scaling the dataset size
+We further evaluate the performance of CLPBA on 6 random poison pairs, on VGG16 and Inception-ResNet model architecture for the Facial Recognition task in Table 10 and Table 11.
+D.3 Scaling the Dataset Size
 To test the effect of dataset size and the number of classes on the performance of CLPBA, we scale our dataset by merging with two benchmark datasets, PubFig [28] and CelebA [35], with the following configurations:
@@
 (8554 images)
-$1-7$ 100 (0.0) 99.7 (0.5) 100 (0.0) 98.0 (2.7)
-$3-6$ 100 (0.0) 100 (0.0) 100.0% (0.0%) 100.0% (0.0%)
-$5-1$ 99.1 (1.6) 93.8% (1.1%) 96.0% (1.4%) 93.5 (4.9)
-$6-8$ 92.8 (3.8) 95.0% (4.1%) 85.1% (6.6%) 84.0 (9.9)
-$7-3$ 99.7 (0.5) 98.5% (1.9%) 100.0% (0.0%) 92.5 (5.9)
+$1-7$ 100 (0.0) 99.7 (0.5) 100.0 (0.0) 98.0 (2.7)
+$3-6$ 100 (0.0) 100 (0.0) 100.0 (0.0) 100.0 (0.0)
+$5-1$ 99.1 (1.6) 93.8 (1.1) 96.0 (1.4) 93.5 (4.9)
+$6-8$ 92.8 (3.8) 95.0 (4.1) 85.1 (6.6) 84.0 (9.9)
+$7-3$ 99.7 (0.5) 98.5 (1.9) 100.0 (0.0) 92.5 (5.9)
 ACC (%) 99.7 (0.1) 99.7 (0.2) 99.3 (0.1) 97.1 (0.2)
@@
 Prediction accuracy (%) 96.4 (0.4) 93.6 (0.2) 93.4 (0.3)
-We note that since the poison ratio $\alpha$ is defined over the target class dataset, the number of poisoned images does not change when we increase the dataset size.
-Interestingly, as shown in Table 16 and Table 15, scaling dataset size only has minimal impacts on the attack performance.
-E.3 Convergence analysis
-We compare the convergence rate of standard $\ell_{\infty}$ regularization with our proposed pixel regularization.
-Figure 12(b) shows the loss of the GM poisoning objective (Equation 4) and the $\ell_{2}$ norm of the perturbations.
-Our pixel regularization leads to a better convergence of the poisoning objective while simultaneously reducing the $\ell_{2}$ norm of the perturbations to a smaller value, indicating an improved perceptual quality.
-Consequently, the GM attack with pixel regularization achieves an ASR of $97.8(1.1)$, compared to $51.9(18.7)$ for the original GM attack under the same conditions.
-(a) Standard $\ell_{\infty}$ regularization
-(b) Pixel Regularization
-Figure 12: Convergence rate of standard $\ell_{\infty}$ regularization and Pixel Regularization.
-Poisoned samples are crafted with GM attack ($\alpha$ = 0.1, $\epsilon$ = 8/255) on poison pair $7-3$ and ResNet50 model
-Appendix F Additional visualization
-F.1 Poisoned samples of different poisoning algorithms
+Since the poison ratio $\alpha$ is defined over the target class dataset, the number of poisoned images does not change when we increase the dataset size.
+Interestingly, as shown in Table 13 and Table 12, scaling dataset size only minimally impacts the attack performance.
+D.4 Evaluating Digital Trigger for CLPBA
+In CLPBA, we assume that the attackers hold trigger samples from the source class to perform poisoning.
+However, this may not be practical in real-world scenarios when the attacker can access clean images from the source class but not the trigger images.
+To mitigate this, the attacker can inject a digital image of the physical trigger into the source-class images as an alternative.
+In this section, we aim to evaluate the performance of CLPBA when the poisoned samples are crafted with photoshopped trigger images.
+As in Table 14, using photoshopped images has varying effectiveness for different poison pairs.
+Table 14: Evaluating the use of digitally-modified trigger images for GM attack. The metric reported is ASR (%)
+Poison pairs 1-7 2-0 6-8 9-5
+GM-digital 98.0 (1.0) 0.0 (0.0) 68.6 (8.6) 66.3 (21.5)
+GM-physical 100.0 (0.0) 99.4 (1.0) 98.3 (0.8) 97.8 (0.2)
+D.5 Evaluating transferability of CLPBA
+We evaluate the black-box settings of CLPBA for two poison pairs $7-1$ and $3-6$ in Table 15 and Table 16 Generally, we find that CLPBA under black-box settings have a worse performance, which means that the semantic features extracted by different model architectures may differ greatly.
+The same attack on different poison pairs may have vastly different results.
+For example, the poison pair $3-6$ with sunglasses trigger has a much better transferability than the poison pair $7-1$.
+This shows that the effectiveness of the trigger depends on the alignment between the trigger and natural features of the target class in the physical space.
+Table 15: Evaluate transferability of CLPBA on poison pair $1-7$
+Attacker Victim Inception-ResNetV1 ResNet50 VGG16
+GM FM GM FM GM FM
+Inception-ResNetV1 96.9 0.3 46.4 16.2 20.8 5.4
+ResNet50 1.7 0.3 99.7 92.0 35.3 6.3
+VGG16 3.7 0.3 91.5 24.2 100.0 83.2
+Table 16: Evaluate transferability of CLPBA on poison pair $3-6$
+Attacker Victim Inception-ResNetV1 ResNet50 VGG16
+GM FM GM FM GM FM
+Inception-ResNetV1 100.0 79.2 30.4 18.8 99.5 97.6
+ResNet50 72.9 55.6 100.0 92.3 98.6 98.1
+VGG16 72.9 49.3 71.5 22.2 100.0 100.0
+Appendix E Additional Visualization
+E.1 Poisoned samples of different poisoning algorithms
 Figure 12: Additional visualization of poisoned samples
@@
 GM is more successful than FM because it successfully reconstructs the trigger in the perturbation in most of the samples.
-F.2 Statistical evaluation
-Figure 14: Visualization of poisoned samples of different poisoning algorithms. Perturbations are scaled 10×10$\times$10 × for visibility.
-For statistical evaluation, we follow previous work [31, 32] in measuring the Peak Signal-to-Noise Ratio (PSNR) between the poisoned image and the original image.
-Quantitatively, FM is the most stealthy as the perturbations have the lowest $\ell_{2}$ norm and the poisoned images have the highest PSNR value.
-Qualitatively, we observe that the noise of FM is more "natural" to the human eye because it is more fine-grained and scattered, resembling natural noise from data augmentation.
-In comparison, the noise from GM and BLC is smoother with some specific patterns (e.g., in GM, the perturbations are more concentrated on face edges and the area underlying the trigger, while in BLC, the noise has dark areas around the images).
-We argue that these more uniform, smooth noise patterns from GM and BLC are visually conspicuous to human inspection, compromising the stealthiness of the attacks.
-F.3 Additional imperceptibility comparison of regularizations
+E.2 Additional imperceptibility comparison of regularizations
 Figure 13: Additional visualizations of GM attack with $\ell_{\infty}$ regularization and Pixel Regularization.
 Left image generated by standard $\ell_{\infty}$ regularization, Right Image generated by Pixel Regularization.
-Appendix G Defenses
+Appendix F Defenses
 To defend against backdoor attacks in DNNs, a multitude of strategies have been developed, each addressing different vulnerabilities.
@@
 Filtering-based defenses [52, 3, 14, 40, 17] identify and filter suspicious data that are likely to be poisoned by adversaries.
-In the paper, we evaluate CLPBA against four representative filtering-based defenses:
+Firewall-based defenses aim to detect malicious inputs at inference time and discard them so that the trained backdoored model will not behave maliciously.
+In the paper, we evaluate CLPBA against four representative filtering-based defenses and four representative firewall defenses:
   • Activation Clustering (AC) [3]: This defense filters poisoned inputs in the latent space of the poisoned model via clustering.
@@
 It then uses K-nearest neighbors to filter samples with the most number of conflicting neighbors (neighbors that have different labels).
-Appendix H Trigger selection
-To validate our proposed heuristic in Equation 8 for selecting effective triggers, we conducted experiments on four poison pairs using all of the triggers.
-Figure 16, presents a comparison between the actual effectiveness of each trigger for each poison pair and the trigger rankings derived from our heuristic formula in Equation 8, referred to as the Selection Score.
-As shown, the top three triggers identified by our heuristic consistently align with the top three triggers with the highest ASR observed through empirical testing using both the FM and GM methods, showcasing the effectiveness of our heuristic in predicting trigger effectiveness.
-Figure 16: Comparison of the actual trigger effectiveness for FM and GM attacks and the trigger effectiveness predicted by the heuristic (Selection Score).
-Appendix I Extended evaluation: animal classifcation
-Besides facial recognition, we also conduct an extended evaluation of CLPBA on animal classification.
-We collected 2000 cat images with 500 trigger-free images and 500 images for each of the triggers, which are tennis balls, mobile phones, and books.
-The trigger-free cat images are resized to $224\times 224$ and then concatenated to an existing animal classification dataset on Kaggle  [^3] https://www.kaggle.com/datasets/asaniczka/mammals-image-classification-dataset-45-animals.
-Similar to our facial recognition dataset, we measure the attack success rate of CLPBA by the percentage of trigger images in the source class that are misclassified as the chosen target class.
-Figure 17: Images of cats playing with the tennis ball as physical trigger
-Due to the time constraint, we only do the evaluation of GM attack on (cat – squirrel) poison pair using tennis ball trigger.
-After some hyperparameter tuning, we set the attack iterations to 1000 and the learning rate to 0.01 (gradually decay to 0.0001 at the end with Cosine Annealing scheduler [46]).
-In our experiment, incorporating PR only slightly improves visual quality (Figure 18) but it allows the poisoning objective to have a much better convergence, which significantly improves the attack performance compared to the standard $\ell_{\infty}$ regularization (Table 17).
-Table 17: Performance of GM attack ($\alpha=0.1$, $\epsilon=16/255$) on (cat – squirrel) poison pair with tennis trigger. The attacker model and victim model are both ResNet50 pre-trained on ImageNet.
-GM GM + PR
-ASR 6.8 (2.0) 90.7 (3.4)
-ACC 95.6 (0.2) 95.7 (0.1)
-(a) Standard $\ell_{\infty}$ regularization
-(b) Pixel Regularization
-Figure 18: Visualization of poisoned squirrel images.
-Appendix J Limitations and future works
+  • STRIP [14]: This defense detects poisoned inputs by applying random perturbations and observing the model’s prediction entropy.
+Poisoned inputs typically produce more consistent (lower entropy) predictions under perturbations compared to clean inputs.
+  • IBD-PSC [19]: This defense clusters inputs based on their feature representations and identifies poisoned samples as distinct clusters in the feature space, separate from clean samples.
+  • Frequency-based Detection [59]: This defense identifies backdoor triggers by analyzing frequency patterns in the input data.
+Trigger artifacts often show statistically distinct patterns that can be isolated through frequency domain analysis.
+  • Cognitive Distillation [20]: This method detects backdoor patterns by isolating minimal features, called Cognitive Patterns (CPs), that trigger the same model output.
+Backdoor samples consistently yield unusually small CPs, making them easy to identify.
+F.1 Evaluation on filtering defenses.
+Backdoor poisoning attacks have been observed to leave discernible traces in the feature space of poisoned models.
+[52, 17, 3, 41].
+By projecting feature representations of target samples (containing both clean and poisoned samples) on the first two principle axes (Figure 14), we can see a clear distinction between the poisoned samples and clean samples.
+The separation exists because the perturbations on poisoned samples contain unnatural trigger features of $\widetilde{D}_{s}$, from which the model can learn as features of the target class.
+However, from Table 2 in the main paper, we can still observe that the GM attack is still effective despite the filtering defenses having high TPR.
+This shows that only a few effective instances are meaningful for the attack.
+Figure 14: 2D-PCA view of target-class representation produced by the poisoned model
+F.2 Evaluation on firewall defenses.
+As demonstrated in Table 2 in the main paper, input-level detection methods are not effective for CLPBAs because they either miss trigger samples or incorrectly filter out too many benign samples.
+This behavior is expected, as CLPBAs challenge the main assumptions underlying these defenses:
+  • STRIP and IBD-PSC: These methods assume that backdoor correlation (trigger and target label prediction) is more consistent than the classification of benign samples, and thus find ways to unlearn normal classification tasks to highlight trigger samples.
+However, since CLPBAs work by synthesizing natural features from the target class with the trigger, unlearning normal classification tasks also unlearns the backdoor correlation between the trigger and the target class.
+  • Frequency-based Detection: This method assumes that poisoned samples exhibit high-frequency artifacts that differ from benign ones.
+This assumption holds true for digital triggers, where there is no inherent correlation between the trigger and the natural image content in the pixel space.
+However, physical triggers, which are integrated naturally into the image, do not produce such high-frequency artifacts, making this defense less effective.
+  • Cognitive Distillation: This approach assumes that a backdoored model focuses on much smaller regions for classifying trigger samples than for classifying clean samples.
+However, because CLPBAs aim to embed the distribution of the source class with the trigger into the feature space of the target class, the classification region for trigger samples is larger.
+The model relies on a combination of the trigger and natural features of the source class for misclassification, as evidenced by the Grad-CAM images in Figure 9 (Appendix A).
+It is also important to note that these defenses are built for dirty-label all-to-one attacks instead of clean-label one-to-one attacks as CLPBA.
+Therefore, the assumption of a consistent backdoor correlation for STRIP and IBD-PSC may not hold true.
+Appendix G Limitations and future works
 High variance.
@@
 A natural direction from this work would be to study clean-label physical backdoor attacks for different tasks, such as Object Detection and Image Segmentation.
-Appendix K Societal impact
-The challenges highlighted by Clean-label Physical Backdoor Attack (CLPBA) can push the AI community towards developing more trustworthy AI models.
+Appendix H Societal impact
+The challenges highlighted by the Clean-label Physical Backdoor Attack (CLPBA) can push the AI community towards developing more trustworthy AI models.
 By identifying and demonstrating the vulnerabilities associated with CLPBA, this research can drive the development of more advanced and effective defense mechanisms against CLPBA.
@@
 Also, policymakers can draft more effective regulations to safeguard against such threats.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The main claims made in the abstract and introduction accurately and adequately reflect the paper’s contributions and scope.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We address the limitations of our work in Appendix J.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: The paper does not have theoretical results.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: In Appendix C, we describe our dataset in details. In Appendix D, we disclose all the necessary experimental settings.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [No]
-  24.
-Justification: Due to ethical issues related to data privacy, we cannot release our facial recognition dataset.
-However, upon acceptance, we will release our animal classification dataset (see Appendix 18) and the code to ensure the reproducibility of our results.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: In Appendix D, we disclose all the necessary experimental settings.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: For each of the experiments, we run 10 times and report the mean with standard deviation.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We provide details regarding compute resources in Appendix D.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We have reviewed the NeurIPS code of conduct and confirm that, to the best of our knowledge, our submission adheres to all its guidelines.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: In Appendix K, we address potential positive and negative societal impacts of our work.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [Yes]
-  54. Justification: We describe the safeguards for our facial recognition dataset in Appendix C.1.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: In Appendix D.1, we have properly credited the creators of the original code framework that we use for the experiments in the paper.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: Our facial recognition dataset is documented in Appendix C, and our animal classification dataset is documented in Appendix 18.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [Yes]
-  69.
-Justification: We have included the instruction and consent forms that we send to our volunteers in the supplementary material.
-The volunteers were asked to read and sign the consent form if they agreed to participate.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [Yes]
-  74.
-Justification: The risks are fully described in the consent form we prepare for the volunteers in our dataset.
-Our study also got IRB approval prior to dataset collection.
-See Appendix C.1 for details.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+Figure 15: Example images of the 10 volunteers, representing the 10 classes in the dataset.
+Figure 16: Physical triggers of the animal classification dataset
+Figure 17: Left column: Main trigger. Right columns: Three false triggers with matching shapes, sizes, colors, and locations on the face to test the false positive rate of the model.

### v3

@@
 Physical triggers, natural objects within a physical scene, have emerged as a desirable alternative since they enable real-time backdoor activations without digital manipulation.
-However, current physical backdoor attacks require poisoned inputs to have incorrect labels, making them easily detectable by human inspection.
-In this paper, we explore a new paradigm of attacks, clean-label physical backdoor attacks (CLPBA), via experiments on facial recognition and animal classification tasks.
+However, current physical backdoor attacks are dirty-label, meaning that the injected poisoned inputs have incorrect labels, making them susceptible to filtering by human inspection.
+Therefore, in this paper, we explore a new paradigm of backdoor attacks, Clean-Label Physical Backdoor Attacks (CLPBA), via experiments on two collected datasets of facial recognition and animal classification.
 Our study reveals that CLPBA could be a serious threat with the right poisoning algorithm and physical trigger.
-A key finding is that different from digital backdoor attacks which exploit memorization to plant backdoors in deep nets, CLPBA works by embedding the feature of the trigger distribution (i.e., the distribution of trigger samples) to the poisoned images through the perturbations.
-We also find that representative defenses cannot defend against CLPBA easily since CLPBA fundamentally breaks the core assumptions behind these defenses.
-Our study highlights accidental backdoor activations as a limitation of CLPBA, happening when unintended objects or classes cause the model to misclassify as the target class.
-The code and dataset can be found in https://github.com/21thinh/Clean-Label-Physical-Backdoor-Attacks.
+Notably, unlike digital backdoor attacks that exploit memorization to plant backdoors in models, CLPBA works by embedding the feature of the trigger distribution (i.e., the distribution of trigger samples) into the poisoned images through the perturbations.
+We also find that representative filtering and firewall defenses are not effective against CLPBA since CLPBA fundamentally breaks the core assumptions behind these defenses.
 1 Introduction
@@
 However, existing physical backdoor attacks are dirty-label, meaning that the labels of trigger samples (i.e., images containing the trigger) are converted to the target-class label.
-The inconsistency between image contents and labels allows data inspectors to easily remove trigger samples from the dataset, especially when such samples contain a visible trigger (e.g., sunglasses) as a common feature.
-Thus, this paper addresses the research question: ”Is it feasible to execute a successful physical backdoor attack by injecting a minimal number of clean-label, trigger-free instances with invisible perturbations into the training set?
+The inconsistency between image contents and labels allows data inspectors to easily remove trigger samples from the dataset, especially when such samples contain a visible trigger (e.g., sunglasses) as a distinctive feature.
+Thus, this paper addresses the research question: ”Is it feasible to execute a successful physical backdoor attack by injecting a small number of poisoned instances that are clean-label and triggerless into the training set?
 Figure 1: General pipeline of CLPBA.
@@
 At inference time, the victim model trained on these perturbed samples will incorrectly classify the source-class samples with the trigger as the target class.
-This paper introduces Clean-Label Physical Backdoor Attacks (CLPBA), as characterized by:
-  • Clean-label: The injected poisoned samples preserve the ground-truth labels.
-  • Hidden-trigger: The poisoned samples are perturbed with constrained noise but they do not explicitly contain the trigger.
+This paper introduces Clean-Label Physical Backdoor Attacks (CLPBA), characterized by:
+  • Clean-label: The injected poisoned samples preserve the ground-truth labels instead of having the target labels.
+  • Hidden-trigger: Unlike standard backdoor attacks, the poisoned samples of CLPBA do not explicitly contain the trigger but are perturbed with constrained noise.
 Thus, they are harder to detect since there is no visible difference between poisoned and clean samples.
   • Real-time: These attacks do not require digital manipulation at test time, allowing for applications in real-time environments, such as facial recognition systems or traffic sign detection for autonomous vehicles.
-Since there are currently no public benchmark datasets to study physical backdoor attacks, we collected a large-scale facial classification dataset that contains 21,238 images of 10 volunteers and 7 triggers representing common facial accessories, following the design of [55].
+Since there are currently no public datasets to study physical backdoor attacks, we collected a large-scale facial classification dataset that contains 21,238 images of 10 volunteers and 7 triggers representing common facial accessories, following the design of [52].
 We also collected 2000 cat images (500 clean images and 1500 trigger images) containing 3 physical triggers (tennis ball, phone, book) and then combined these images with an existing animal classification dataset [1].
@@
   1. 1.
-    Our study reveals that different from digital backdoor attacks which exploit memorization to plant backdoors in deep nets, CLPBA attacks the model by embedding the features of the source-class trigger distribution (i.e., the distribution of trigger images from the source class that are to be misclassified) to the poisoned images.
+    We introduce CLPBA, a new threat of physical backdoor attacks that optimize perturbations on clean-label samples to embed the backdoor, instead of injecting dirty-label trigger samples.
+Thus, the attacks can be both stealthy (clean-label, hidden-trigger) and applicable in real-time via physical triggers.
   2. 2.
+    Our study reveals that, different from digital backdoor attack,s which exploit memorization to plant backdoors in deep nets, CLPBA attacks the model by embedding the features of the source-class trigger distribution (i.e., the distribution of trigger images from the source class that are to be misclassified) to the poisoned images.
+  3. 3.
     We identify a trade-off between perceptual stealthiness and the effectiveness of CLPBA.
@@
 Empirically, our regularization improves the convergence rate of poisoning algorithms, thereby enhancing the effectiveness and stealthiness of the poisoned samples.
-  3. 3.
-    Our study identifies accidental backdoor activations as a key limitation of CLPBA, occurring when unintended objects or classes cause the model to misclassify as the target class.
-To address this, we introduce repel terms in the adversarial loss to reduce the victim model’s sensitivity to trigger instances from unintended classes.
-  4. 4.
-    To facilitate research into physical backdoor attacks, we publicly release the trigger images of the animal classification dataset and the code for reproducibility.
 2 Related Works
@@
 Thus, we are the first to comprehensively study the performance of clean-label backdoor attacks with physical triggers on two tasks, facial recognition and animal classification.
-3 Methodology
-3.1 Threat Model
-Attacker’s capability.
-We follow the same threat model as hidden-trigger backdoor attacks [44, 49].
-The training dataset is collected from multiple sources, and there exists an attacker (e.g., data contributors, data collectors) who can access and perturb a small portion of the data to manipulate the behavior of the victim model.
-The attack is one-to-one: any instances from the source class are misclassified as belonging to the target class with the presence of the trigger.
-For example, the source class is an employee in a company who can wear a special pair of sunglasses to fool a facial recognition system into classifying him as the CEO so that he could get an illegal privilege, such as access to confidential documents or high-ranking facilities.
-Due to the nature of the physical trigger, we argue that all-to-one attacks have low utility as this scenario could raise suspicion among model users; for example, benign participants wearing the trigger object could accidentally activate the backdoor behavior and report the issue to authority.
-Figure 2: Physical triggers in the collected datasets
-We further assume the attacker can select a trigger from a limited pool of physical objects and has access to trigger samples from the source class.
-This assumption is reasonable, as the source class could be the attacker or collaborators.
-In some cases, physical triggers, being natural objects, may already exist in the training dataset of the target class [56].
-In this paper, we consider a realistic scenario where the target class dataset lacks natural triggers, and the attacker cannot directly inject trigger instances into it.
-3.2 Problem Formulation
-For the rest of the paper, we define trigger instances as the clean samples with the injected trigger, and the poisoned instances as the injected samples to the training dataset that backdoors the model.
-For dirty-label backdoor attacks, poisoned instances are trigger instances with incorrect labels (i.e., mislabeled as the target class).
-For hidden-trigger backdoor attacks, however, the poisoned instances are images that are perturbed adversarially without the trigger.
-Notations.
-Denote $y_{t}$ as the label of target class.
-Let $D=\bigcup^{k}_{i=1}D_{i}$ be the training dataset with $k$ classes.
-Let $\mathcal{D}_{i}$, $\widetilde{\mathcal{D}}_{i}$ be the distribution of instances in class $i$ without and with the physical trigger.
-Let $F\colon\mathcal{X}\rightarrow R^{k}$ be the classifier model, with parameters $\theta$, that gives prediction scores (i.e., logits) for each of $k$ classes given an input $x\in\mathcal{X}$.
-$F$ can be decomposed as $F=f\circ g$ with $f$ as the feature extractor and $g$ be the classifier.
-The poison pair $s-t$ is the pair of source class and target class with index labels.
-The attacker aims to inject a poison set ${P}$ to the training set ${D}$ such that the victim model trained on ${D}^{p}:={D}\cup{P}$ will misclassify trigger samples from the source class as the target class.
-To achieve this goal, the attacker needs to solve the following bilevel optimization problem:
-$\underset{\delta\in C}{\text{min}}\;L_{adv}(\delta)=\mathbb{E}_{(\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s})}\Bigl{[}\mathcal{L}\bigl{(}F(\tilde{x}_{s};\theta(\delta)),y_{t}\bigr{)}\Bigr{]}$ (1)
-$\theta(\delta)\in\operatorname*{arg\,min}_{\theta}\;L_{train}(\theta)=\sum_{x,y\in D^{p}}\mathcal{L}\bigl{(}F(x;\theta),y\bigr{)}$ (2)
-where $C$ is the constraint set of the perturbations, $\mathcal{L}$ is the objective function and $\widetilde{\mathcal{D}}_{s}$ is the distribution of trigger instances from the source class.
-Following previous works [44, 49, 60], the poison set $P$ is a small subset of $D_{t}$.
-Formally, ${P}=\{x_{t}^{(i)}+\delta^{(i)},y_{t}\}^{K}_{i=1}$ contains $K$ poisoned target-class samples.
-The constraint set is $\ell_{\infty}$-norm ball: $C=\{\delta\;|\>||\delta||_{\infty}<\epsilon\}$.
-The value $\epsilon$ controls the magnitude of the perturbations and thus the stealthiness of poisoned samples.
-We define the poison ratio $\alpha$ as the percentage of poisoned samples in the target-class data.
-Note that this may differ from previous works on backdoor attacks which define poison ratio over the entire training set.
-With the above formulation, we introduce three poisoning algorithms for CLPBA based on hidden-trigger backdoor attacks [44, 49] and clean-label data poisoning [21]:
-  • Feature Matching (FM) [44]: This method targets transfer-learning models by finding perturbations that shift poisoned samples closer to $\widetilde{\mathcal{D}}_{s}$ in the feature space:
-    $\mathcal{A}_{FM}=\min_{\delta\in C}\|f(x_{t}+\delta)-f(\tilde{x}_{s})\|_{2}^{2}\quad$ (3)
-    $\text{where }x_{t}\in P,\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s}$ (4)
-  • Meta Poison (MP) [21]: This method is based on meta-learning [13], where it approximates the inner objective $\theta(\delta)$ by unrolling $K$ (default to 2) SGD steps before computing outer objective:
-    $\theta_{1}$ $=\theta_{0}-\alpha\nabla_{\theta}L_{train}(\theta_{0}\;|\;\delta_{i})$ (5)
-    $\theta_{2}$ $=\theta_{1}-\alpha\nabla_{\theta}L_{train}(\theta_{1}\;|\;\delta_{i})$ (6)
-    $\delta_{i+1}$ $=\delta_{i}-\beta\nabla_{\delta}L_{adv}(\delta_{i}\;|\;\theta_{2})\text{ s.t }\delta\in C$ (7)
-  • Gradient Matching (GM) [49]: This method targets DNNs trained from scratch by aligning the gradient of the model’s parameters on the poisoned dataset $P$ with the gradient to optimize $L_{adv}$:
-    $\mathcal{A}_{GM}$ $=\min_{\delta\in C}\left(1-\frac{\nabla_{\theta}L_{poi}\cdot\nabla_{\theta}L_{adv}}{\|\nabla_{\theta}L_{poi}\|\cdot\|\nabla_{\theta}L_{adv}\|}\right)$ (8)
-    where $\text{where }L_{poi}$ $=\frac{1}{K}\sum_{i=1}^{K}L(F(x_{t}^{(i)}+\delta^{(i)};\theta),y_{t})$ (9)
-As a baseline, we also compare the poisoning algorithms with the Naive attack, in which the attacker injects trigger target-class samples into the target-class dataset so that the model learns the trigger as a feature of the target class.
-3.3 Data Collection
-Our facial dataset contains 21,238 images of 10 volunteers (9 males, 1 female, all Asian) and 7 triggers.
-It consists of 3,344 clean images, 9,790 trigger images of 7 physical triggers, and 8,104 additional images to evaluate accidental backdoor activations (see Section 7).
-Our animal dataset is created by combining a public Kaggle dataset [1], containing 45 classes of mammals, with a new cat class containing 500 cat images.
-For this dataset, we collected 1,500 cat images with 3 physical triggers (tennis ball, phone, book) to study CLPBA.
-The images are captured in various indoor and outdoor settings, under varying weather conditions, and with diverse shooting angles and distances.
-Details of our dataset can be found in the Appendix B.
-Table 1: ASR (%) of poisoning algorithms on 5 random source-target pairs, averaged over 10 runs with standard deviation in brackets.
-Poison pair Poison ratio $\alpha$ Naive MP FM GM
-$\epsilon=8/225$ $\epsilon=16/255$ $\epsilon=8/225$ $\epsilon=16/255$ $\epsilon=8/225$ $\epsilon=16/255$
-(a) Facial Recognition (sunglasses trigger)
-1 – 7 5% 35.0 (18.9) 14.2 (2.6) 10.5 (3.6) 84.0 (3.9) 80.9 (7.4) 97.2 (0.5) 99.4 (0.5)
-10% 81.5 (3.5) 13.7 (2.3) 13.4 (2.6) 82.3 (0.5) 86.9 (2.5) 98.0 (0.5) 100.0 (0.0)
-2 – 0 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 85.0 (4.4)
-10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 99.4 (1.0)
-4 – 3 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.4 (1.2) 15.8 (16.4) 96.4 (1.2)
-10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 12.8 (11.6) 48.9 (16.2) 98.6 (0.3)
-6 – 8 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.7 (2.9) 18.2 (12.3) 7.2 (11.0) 97.8 (2.4)
-10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 4.7 (8.1) 22.9 (8.3) 21.8 (19.2) 98.3 (0.8)
-9 – 5 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.6 (0.5) 5.8 (3.2) 22.7 (10.9) 95.6 (1.8)
-10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 3.4 (0.5) 30.4 (1.5) 51.4 (17.8) 97.8 (0.2)
-(b) Animal Classification (tennis trigger)
-cat – fox 5% N/A 0.0 (0.0) 0.0 (0.0) 0.3 (0.5) 0.2 (0.3) 8.2 (4.4) 9.6 (4.1)
-10% N/A 0.0 (0.0) 0.0 (0.0) 1.0 (0.8) 0.2 (0.4) 20.5 (8.9) 17.8 (7.7)
-cat – kangaroo 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.9 (0.1) 8.4 (0.8)
-10% N/A 0.0 (0.0) 0.0 (0.0) 0.1 (0.1) 0.1 (0.1) 9.4 (2.1) 18.0 (2.0)
-cat – mongoose 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 2.7 (0.3) 5.3 (0.4)
-10% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 6.7 (0.1) 10.3 (2.9)
-cat – sea-lion 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 7.4 (0.6) 13.4 (2.9)
-10% N/A 0.0 (0.0) 0.0 (0.0) 0.3 (0.1) 0.0 (0.0) 22.1 (5.3) 23.6 (5.8)
-cat – tapir 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.8 (0.8) 7.8 (5.0)
-10% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 29.8 (8.6) 36.3 (7.6)
-4 Experiments
-For Facial recognition, we choose the victim model to be ResNet50 [18] pre-trained on the VGGFace2 dataset [4].
-For Animal classification, the victim model is ResNet18 pre-trained on Imagenet [43].
-We assume a gray-box scenario where the attacker knows the architecture of the victim model but does not have access to its training process.
-To enable a fair comparison across different poisoning algorithms, we standardize the poisoning hyperparameters, including the number of iterations, optimizer, batch size, learning rate, and other relevant parameters.
-The performance of attacks is measured by attack success rate (%) (ASR), the percentage of trigger source-class samples that are misclassified as the target class.
-Note that the prediction accuracy (ACC) on clean samples consistently exceeds 99% for facial recognition and 93% for animal classification across different attacks, so it is not reported in the paper.
-Detailed experimental configurations can be found in Appendix C.
-Appendix D provides additional experiments on other architectures, hyperparameters, and physical triggers.
-4.1 Comparison of Poisoning Algorithms
-The results from Table 1 show that GM is the most effective poisoning algorithm.
-Higher values of $\epsilon$ and $\alpha$ could increase ASR, but $\epsilon$ has a greater impact as it allows the poisoning objective to converge to smaller values and the poisoned samples to contain more adversarial features of the source-class trigger distribution.
-Additionally, CLPBA is more effective for facial recognition than animal classification, probably because the ”sunglasses” is a very dominant feature that hides important classification features (i.e., the eyes and the nose) for face recognition, while the ”tennis ball” trigger is a relatively minor feature so it cannot outweigh the natural features of the cat.
-Dataset bias.
-Interestingly, we can observe that the poison pair 1-7 in Table 1a has an unusually high ASR compared to other pairs.
-Even with no trigger images injected, ASR reaches 26.21 % for the Naive method.
-We believe this is because certain samples in the target class contain similar features to the physical trigger, reinforcing the trigger as a dominant feature once trigger images are added.
-In this case, it may be the eyeglasses of the target-class that have a similar shape and texture to the sunglasses.
-While this correlation may be accidental due to our dataset collection and trained classifier, it still raises an important issue of dataset bias that a backdoor attacker can exploit to select an effective trigger aligning with the natural features of the target class.
-In Appendix A, we show that the removal of training samples with strong trigger features drastically reduces the ASR and thus prevents the dataset bias.
-4.2 Analysis of CLPBA
-Two unique attributes of CLPBA are that it is hidden-trigger and source-specific.
-The first attribute means the trigger is not present in the poisoned samples, while the second ensures only source-class samples can cause the model to misclassify with the trigger.
-CLPBA achieves these attributes by embedding features of the source-class trigger distribution, $\widetilde{\mathcal{D}}_{s}$, into the perturbations in the poisoned target samples.
-Specifically, the embedded features consist of adversarial features and trigger features.
-We define adversarial features as the non-robust features [25] that shift the representations of target instances toward the source-class distribution in the feature space, and trigger features as the high-level features of the physical trigger extracted by the neural network.
-The existence of adversarial features ensures the source-specific attribute, while reconstructed trigger features allow the model to indirectly learn the physical trigger as a feature of the target class.
-Figure 3: Experiment to show the existence of adversarial features. We set a high $\epsilon$ of 32, $\alpha=2$, poison pair is $9-5$.
-Existence of adversarial features.
-To justify the existence of adversarial features, we take inspiration from [25] to test whether the model trained on poisoned target samples could successfully predict source-class samples.
-We design the experiment as illustrated in Figure 3.
-First, we craft the GM-poisoned samples with the original dataset and the attacker model.
-Then, we filter out all source-class samples, before relabelling all GM-poisoned samples to the source class.
-We then train the victim model from scratch on this modified dataset before evaluating the recall (true positive rate) of predicting source-class samples that have been filtered previously.
-To avoid randomness, we set a prediction threshold of 85% after the softmax function.
-Notably, the model trained on this modified dataset has a prediction accuracy of 75% on source-class samples.
-This shows that the perturbations on poisoned target instances contain adversarial features representative of source-class distribution.
-To make sure that the trained model is not biased towards predicting out-of-distribution samples as the source class, we also evaluate the false positive rate of classifying 200 random face images from the CelebA dataset [35] as the source class, and the accuracy is 0%.
-Existence of trigger features.
-By scaling the perturbations on GM-poisoned samples, we observe that the perturbations appear to reconstruct the shape of the physical trigger.
-Interestingly, setting a high value of $\epsilon$ even reconstructs the samples in $\widetilde{\mathcal{D}}_{s}$, as can be observed in Figure 5.
-(a) Clean model
-(b) Poisoned model
-Figure 4: Feature space visualization.
-Feature space analysis.
-In Figure 4, we project source-class samples (green), target-class samples (blue), trigger source samples (yellow), and poisoned target samples (red) in 2-dimension feature space of both clean and poisoned models where the first axis is $\mu_{s}-\mu_{t}$ and the second axis is $\theta_{s}-\theta_{t}$, the normal vector to the decision boundary between the source and target class.
-In both feature spaces, we can see that some poisoned target instances are close to the samples of $\widetilde{\mathcal{D}}_{s}$.
-For the clean model, $\widetilde{\mathcal{D}}_{s}$ overlaps with ${\mathcal{D}}_{s}$ in the feature space.
-For the poisoned model, however, $\widetilde{\mathcal{D}}_{s}$ is shifted towards $\widetilde{\mathcal{D}}_{t}$ in the feature space.
-From this observation, we can say that since poisoned target instances contain representative features of $\widetilde{\mathcal{D}}_{s}$, especially the trigger feature which is not a part of ${\mathcal{D}}_{s}$, training the model on these poisoned instances cause $\widetilde{\mathcal{D}}_{s}$ to move closer to the classification region of the target class in the feature space.
-Figure 5: From left to right: tennis, fake-beard, and sunglasses triggers. First row: a sample in $\widetilde{D}_{s}$. Second row: GM perturbation.
-4.3 Defenses
-We evaluate the performance of CLPBA against representative filtering defenses (filter poisoned samples during training), and firewall defenses (filter trigger samples during inference).
-For the filtering defenses, we evaluate Elimination Rate (ER), the percentage of poisoned samples that are correctly filtered, and Sacrifice Rate (SR), the percentage of clean samples that are incorrectly filtered.
-For firewall defenses, we evaluate the True Positive Rate (TPR), the percentage of trigger source-class samples that are correctly filtered, and the False Positive Rate (FPR), the percentage of non-trigger samples that are incorrectly filtered.
-For details on these defenses, we refer readers to Appendix F.
-Table 2: Defense evaluation on GM attack ($\alpha=0.1$, $\epsilon=16/226$, poison pair $7-0$). ASR is 99.7% without defenses.
-Filtering Defenses
-Metrics AC[3] SS[52] SPECTRE[17] SCAN[51]
-ER (%) 62.16 97.30 72.97 0
-SR (%) 15.77 19.77 19.50 0
-ASR (%) 99.7 0.0 99.7 99.7
-Firewall Defenses
-Metrics STRIP[14] CD[20] IBD-PSC[19] Frequency[59]
-TPR (%) 0.0 79.6 0.0 0.0
-FPR (%) 6.7 65.3 16.7 0.9
-For filtering defenses, we evaluate Spectral Signature (SS) [52], Activation Clustering (AC) [3], SPECTRE [17], and SCAN [51].
-Table 2 shows that only SS is effective against CLPBA.
-Although we can see that AC and SPECTRE have high elimination rates (62.16% and 72.97%), the attack still achieves 99.7 % ASR.
-This shows that the GM poisoning objective (Equation 8) focuses on a small subset of effective poison instances to match the gradient of adversarial loss and distill the features from the trigger distribution.
-This is supported in Figure 4 where only a subset of poisoned samples overlap with the trigger distribution.
-Detecting such effective instances is generally difficult since they correlate strongly with each other and they transform the model’s decision boundary of the target class.
-Table 2 also shows that firewall defenses are not effective for CLPBAs because they either miss trigger samples (low TPR) or incorrectly filter out too many benign samples (high FPR).
-This behavior is expected since these defenses are built on the memorization property of backdoor attacks, which differs from the generalization property of CLPBA.
-5 Improving the Effectiveness of CLPBA
-Empirically, higher $\epsilon$ lead to a higher ASR since the poisoning objective can converge to a smaller value.
-However, the perceptual quality of poisoned samples is compromised since the poisoned images have high-frequency artifacts that are clearly visible (Figure 5).
-The drawbacks of the $\ell_{\infty}$ regularization on perturbations are that it is non-differentiable, and the projection on $\ell_{\infty}$-norm ball is only executed at the end of an optimization step, making backdoor and poison stealthiness two separate tasks.
-As a consequence, the perceptual quality of poisoned images is compromised because (1) the pixel values of perturbation are often pushed to the boundary of the norm ball, (2) the noise can appear in unnecessary regions of the image, and (3) the perturbations contain high-frequency artifacts.
-To improve the visual quality and convergence of the poisoning objective, we propose to replace the current $\ell_{\infty}$ regularization with a Pixel regularization containing a soft pixel-wise $\ell_{\infty}$ loss and discretized Total Variation Upwind (TVU) [6]:
-$\mathcal{R}_{pix}(\delta)$ $=||\delta^{+}-M\odot\epsilon||_{1}+\alpha TVU(\delta;\beta=1)$ (10)
-$TVU(\delta;\beta)$ $=\sum_{p}^{\text{all pixels}}\biggl{(}\sum_{q\in\mathcal{N}(p)}\big{(}\delta^{(p)}-\delta^{(q)}\big{)}^{2}\biggr{)}^{\frac{\beta}{2}}$ (11)
-where $\delta^{+}$ takes the absolute value of every element in $\delta$, $M$ is a mask of pixel-wise embedding ratio: $M_{i,j,k}=\frac{x_{i,j,k}}{\max_{k}{x_{i,j,k}}}$, assuming $x_{i,j,k}\in[0,255]$.
-$M$ encourages the perturbation to align with the underlying image content.
-In Equation 11, $\mathcal{N}(p)$ are neighbouring pixels of pixel $p$.
-The weighting factor $\alpha$ is set to 5.
-Both terms in Equation 10 improve the smoothness of perturbations, with $TVU(\delta)$ minimizing the difference between adjacent pixels.
-The side-by-side comparison between standard $\ell_{\infty}$ regularization and our pixel regularization is given in Figure 6.
-Our pixel regularization leads to a better convergence of the poisoning objective (Equation 8) while simultaneously reducing the $\ell_{2}$ norm of the perturbations to a smaller value, indicating an improved perceptual quality.
-Consequently, the GM attack with pixel regularization achieves an ASR of $97.8(1.1)$, compared to $51.9(18.7)$ for the original GM attack under the same settings.
-We also evaluate the perceptual quality of the poisoned samples statistically using Peak Signal-to-Noise Ratio (PSNR) [24], following previous works [29, 31].
-Our pixel regularization results in higher values of PSNR (36.9dB vs 29.8dB), corresponding to a better perceptual quality.
-(a) Standard $\ell_{\infty}$ regularization
-(b) Pixel Regularization
-(c) Poisoning loss
-(d) Perturbation norm
-Figure 6: Comparison between standard $\ell_{\infty}$-regularization and our Pixel Regularization.
-Top row: visual comparison; Bottom row: convergence rate.
-Poisoned samples are crafted with GM attack ($\alpha=0.1$, $\epsilon=8/255$) on poison pair 77773333.
-6 Analysis of Attack Factors
-The results in Table 1 show that ASR varies between different poison pairs for a given trigger, while the results in Table 4 show that different triggers may have different performances for a given poison pair.
-This section examines the impact of poison pairs and the physical trigger on CLPBA.
-Poison pair Clean model loss Backdoored model ASR (%)
-3-6 0.6855 99.0 (0.8)
-4-6 0.0038 86.8 (2.1)
-Table 3: Performance of GM attack on source class 3 and source class 4, for a fixed target class 6. Clean model loss is measured over inference-time trigger samples of the source class.
-Source class.
-A source class is more susceptible to physical attacks at inference if the trigger obscures the class’s discriminative classification regions, enabling the trigger features to outweigh salient features and cause misclassification.
-In Table 3, the classification loss of the clean ResNet50 model on inference-time trigger samples is much higher for source class 3 than for source class 4, resulting in a higher ASR for class 3.
-In Appendix A, we use Grad-CAM [47] to visualize discriminative regions detected by a clean and poisoned ResNet50 model (i.e., trained on clean and poisoned data).
-We observe that in trigger instances of source class 3, the clean model struggles to extract meaningful facial features and relies on less prominent features for classification.
-Conversely, the poisoned model focuses on the physical trigger for classification.
-Poison pair N/A Earrings Fake Beard Headband Hat Sticker Sunglasses Face Mask
-7-3 GM 10.8 (5.3) 71.1 (11.0) 8.3 (4.8) 62.8 (8.6) 0.0 (0.0) 46.6 (12.2) 93.7 (2.5)
-FM 0.0 (0.0) 28.5 (5.0) 0.3 (0.3) 21.5 (7.7) 0.0 (0.0) 0.3 (0.4) 86.6 (10.3)
-9-5 GM 1.6 (2.3) 97.8 (0.8) 0.0 (0.0) 19.7 (28.3) 65.1 (14.2) 25.3 (11.3) 58.5 (16.1)
-FM 0.0 (0.0) 90.3 (7.0) 0.0 (0.0) 0.0 (0.0) 14.8 (0.2) 5.8 (5.9) 20.1 (5.4)
-Table 4: Performance of GM attack using various physical triggers. Poisoned samples are crafted with $\epsilon=8$ and $\alpha=0.1$.
-Target class.
-A target class can be more susceptible to CLPBA due to the similarity between the class’s natural features and the trigger features.
-For example, in Appendix A, we show that the eyeglasses in some samples of the target class could represent weaker backdoor features for the attacker’s trigger.
-We note that there is also an unexplored relationship between the distribution of samples in the target class $\mathcal{D}_{t}$ and the distribution of trigger samples in the source class $\widetilde{\mathcal{D}}_{s}$ that allows for feature alignment of poisoned samples in high dimensional space.
-Physical trigger.
-We evaluate GM and FM attacks on poison pairs $7-3$ and $9-5$ using triggers of different sizes and levels of stealthiness (see Appendix B for details on our trigger selection).
-We observe that the effectiveness of a trigger depends on the poison pair.
-Surprisingly, in poison pair $9-5$, the ”face mask”, a large trigger, is not as effective as the ”fake beard” trigger, a small and stealthy trigger.
-From our previous analysis, there are two factors impacting trigger effectiveness: whether the trigger aligns with the natural features of the target class, and whether the trigger hides important features of the source class.
-7 Accidental Backdoor Activations
-For CLPBA to be stealthy, we argue that it needs to be source-specific and trigger-specific.
-That is, (1) only the source class can use the trigger to manipulate the victim model’s prediction at inference time; (2) the trigger should be discriminative so that the victim model does not incorrectly recognize similar-looking objects as triggers.
-We introduce two new metrics, False Positive Rate, and Suspicion Rate, to test accidental backdoor activations:
-  • Suspicion rate (SR): The measurement of whether the victim model incorrectly misclassifies trigger samples from non-source classes as belonging to the target class.
-  • False positive rate (FPR): The measurement of whether samples from the source class can fool the victim model with false triggers (i.e., objects that resemble the attacker’s physical trigger).
-Poison pair Trigger Loss type ASR ACC FPR SR
-9-5 Fake beard Original 98.7 99.9 50.8 12.5
-With repel 98.7 99.9 49.7 3.9
-7-3 Sunglasses Original 97.5 99.3 0.7 13.9
-With repel 97.5 99.4 0.9 2.5
-Table 5: FPR and SR for GM attack. The set-up of our dataset to evaluate SR and FPR is given in Appendix B.
-Between the two metrics, SR is more important because it directly impacts attack stealthiness if a benign client is misclassified as the target class with the trigger or similar-looking items.
-The issue of false triggers, measured by FPR, is challenging to solve as the attacker cannot anticipate all possible false triggers in the real world.
-However, the attacker can partially mitigate the issue of false source classes, as measured by SR, by adding repel terms to the adversarial loss $L_{adv}$ in Equation 2:
-$L_{adv}=N\cdot\overbrace{\sum_{x\in\widetilde{D}_{s}}\mathcal{L}\bigl{(}F_{\theta}(x),y_{t}\bigr{)}}^{\text{adversarial loss}}$
-[…]
-Here, $N$ is the total number of classes, and $\widetilde{D}_{i}$ is the dataset of trigger instances in class $i$.
-The repel terms ensure trigger instances from non-source classes are repelled from the target class classification, enhancing trigger specificity to the source class.
-As shown in Table 5, incorporating the repel term in the GM objective (Equation 8) significantly reduces SR while maintaining high ASR.
-8 Conclusion
-Through extensive experiments, our research underscores the vulnerability of DNNs to clean-label physical backdoor attacks.
-We then show that CLPBA works by embedding adversarial features and trigger features in the poisoned instances.
-Additionally, we propose a novel pixel regularization to improve the effectiveness and stealthiness of poisoned samples.
-A significant limitation identified is accidental backdoor activations of the backdoored model due to false triggers or non-source classes, which could compromise attack stealthiness.
-To address this, we propose incorporating repel terms into the adversarial objective to enhance trigger specificity to the source class.
-Our findings highlight the critical need for robust defenses against CLPBA.
-By releasing the dataset and code, we hope to facilitate future research on improving the resilience of DNNs against such threats.
-Supplementary Material
-Appendix A Natural Physical Backdoors
+3 Natural Physical Backdoors
 For a backdoored DNN, the backdoor trigger can be understood as the strong feature for classification of the target class [38, 25, 49].
@@
 The average distance of sample from $\widetilde{\mathcal{D}}_{t}$ to $\widetilde{\mathcal{D}}_{s}$ is even smaller than the average distance of sample from $\mathcal{D}_{s}$ to $\widetilde{\mathcal{D}}_{s}$.
-This means that when the model overfits to the sunglasses feature of the target class dataset, inference-time samples from $\widetilde{\mathcal{D}}_{s}$ will be mapped to a region closer to $\widetilde{\mathcal{D}}_{t}$ than $\mathcal{D}_{s}$ in the feature space (Figure 8), resulting in misclassification.
+This means that when the model overfits to the sunglasses feature of the target class dataset, inference-time samples from $\widetilde{\mathcal{D}}_{s}$ will be mapped to a region closer to $\widetilde{\mathcal{D}}_{t}$ than $\mathcal{D}_{s}$ in the feature space (Figure 3), resulting in mis-classification.
 (a) Clean samples (w/o trigger)
@@
 The attention is targeted at ground truth labels (i.e, the first row is target 3 and the second row is target 4).
+Table 2: GM attack performance for source classes 3 and 4 (target: class 6). Clean model loss is measured on inference-time trigger samples from the source class.
+Source $\to$ Target Clean Loss ASR (%)
+3 $\to$ 6 0.6855 99.0 (0.8)
+4 $\to$ 6 0.0038 86.8 (2.1)
 […]
-Appendix B Dataset
-B.1 Ethics & Data Protection
+4 Clean-Label Physical Backdoor Attack
+4.1 Problem Formulation
+The effectiveness of the Naive attack, as introduced in Section 3, is highly dependent on the combination of source class, target class, and physical trigger and thus may not generalize to realistic settings.
+Furthermore, the existence of a physical trigger in inserted poisoned images may raise suspicion.
+In this section, we introduce the problem formulation and poisoning algorithms of CLPBA, which crafts triggerless poisoned samples that indirectly plant the backdoor (activated by a physical trigger) into the models.
+Notations.
+Denote $y_{t}$ as the label of target class.
+Let $D=\bigcup^{k}_{i=1}D_{i}$ be the training dataset with $k$ classes.
+Let $\mathcal{D}_{i}$, $\widetilde{\mathcal{D}}_{i}$ be the distribution of instances in class $i$ without and with the physical trigger.
+Let $F\colon\mathcal{X}\rightarrow R^{k}$ be the classifier model, with parameters $\theta$, that gives prediction scores (i.e., logits) for each of $k$ classes given an input $x\in\mathcal{X}$.
+$F$ can be decomposed as $F=f\circ g$ with $f$ as the feature extractor and $g$ be the classifier.
+The poison pair $s-t$ is the pair of source class and target class with index labels.
+The attacker aims to inject a poison set ${P}$ to the training set ${D}$ such that the victim model trained on ${D}^{p}:={D}\cup{P}$ will misclassify trigger samples from the source class as the target class.
+To achieve this goal, the attacker needs to solve the following bilevel optimization problem:
+$\underset{\delta\in C}{\text{min}}\;L_{adv}(\delta)=\mathbb{E}_{(\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s})}\Bigl{[}\mathcal{L}\bigl{(}F(\tilde{x}_{s};\theta(\delta)),y_{t}\bigr{)}\Bigr{]}$ (2)
+$\theta(\delta)\in\operatorname*{arg\,min}_{\theta}\;L_{train}(\theta)=\sum_{x,y\in D^{p}}\mathcal{L}\bigl{(}F(x;\theta),y\bigr{)}$ (3)
+where $C$ is the constraint set of the perturbations: $C=\{\delta\;|\>||\delta||_{\infty}<\epsilon\}$.
+$\mathcal{L}$ is the objective function and $\widetilde{\mathcal{D}}_{s}$ is the distribution of trigger instances from the source class.
+The poison set $P$ is a small subset of $D_{t}$.
+Formally, ${P}=\{x_{t}^{(i)}+\delta^{(i)},y_{t}\}^{K}_{i=1}$ contains $K$ poisoned target-class samples.
+The value $\epsilon$ controls the magnitude of the perturbations and thus the stealthiness of poisoned samples.
+We define the poison ratio $\alpha$ as the percentage of poisoned samples in the target-class data.
+Note that this may differ from previous works on backdoor attacks which define poison ratio over the entire training set.
+With the above formulation, we introduce three poisoning algorithms for CLPBA based on existing clean-label data-poisoning methods:
+  • Feature Matching (FM) [41]: This method crafts perturbations that shift poisoned samples closer to source trigger distribution $\widetilde{\mathcal{D}}_{s}$ in the feature space, thus representing them as dirty-label samples with trigger:
+    $\mathcal{A}_{FM}=\min_{\delta\in C}\|f(x_{t}+\delta)-f(\tilde{x}_{s})\|_{2}^{2}\quad$ (4)
+    $\text{where }x_{t}\in P,\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s}$ (5)
+  • Meta Poison (MP) [21]: This method is based on meta-learning [13], where it approximates the inner objective $\theta(\delta)$ by unrolling $K$ (default to 2) SGD steps before computing outer objective:
+    $\theta_{1}$ $=\theta_{0}-\alpha\nabla_{\theta}L_{train}(\theta_{0}\;|\;\delta_{i})$ (6)
+    $\theta_{2}$ $=\theta_{1}-\alpha\nabla_{\theta}L_{train}(\theta_{1}\;|\;\delta_{i})$ (7)
+    $\delta_{i+1}$ $=\delta_{i}-\beta\nabla_{\delta}L_{adv}(\delta_{i}\;|\;\theta_{2})\text{ s.t }\delta\in C$ (8)
+  • Gradient Matching (GM) [46]: This method crafts perturbations that align the gradient of the model on the poisoned dataset $P$ with the gradient to optimize $L_{adv}$:
+    $\mathcal{A}_{GM}$ $=\min_{\delta\in C}\left(1-\frac{\nabla_{\theta}L_{poi}\cdot\nabla_{\theta}L_{adv}}{\|\nabla_{\theta}L_{poi}\|\cdot\|\nabla_{\theta}L_{adv}\|}\right)$ (9)
+    where $\text{where }L_{poi}$ $=\frac{1}{K}\sum_{i=1}^{K}L(F(x_{t}^{(i)}+\delta^{(i)};\theta),y_{t})$ (10)
+4.2 Threat Model
+Attacker’s capability.
+In our threat model, the training dataset is collected from multiple sources, and there exists an attacker (e.g., data contributors, data collectors) who can access and perturb a small portion of the data to manipulate the behavior of the victim model.
+Following previous works [41, 46], we assume a gray-box scenario where the attacker knows the architecture of the victim model but does not have access to its training process.
+The attack is one-to-one: any instances from the source class are misclassified as belonging to the target class with the presence of the trigger.
+For example, the source class is an employee in a company who can wear a special pair of sunglasses to fool a facial recognition system into classifying him as the CEO so that he could get an illegal privilege, such as access to confidential documents or high-ranking facilities.
+Due to the nature of the physical trigger, we argue that all-to-one attacks have low utility as this scenario could raise suspicion among model users; for example, benign participants wearing the trigger object could accidentally activate the backdoor behavior and report the issue to authority.
+Figure 5: Physical triggers in the collected datasets
+We further assume the attacker can select a trigger from a limited pool of physical objects and has access to trigger samples from the source class.
+This assumption is reasonable, as the source class could be the attacker or collaborators.
+In some cases, physical triggers, being natural objects, may already exist in the training dataset of the target class [53].
+In this paper, we consider a realistic scenario where the target class dataset lacks natural triggers, and the attacker cannot directly inject trigger instances into it.
+4.3 Data Collection
+Our facial dataset contains 21,238 images of 10 volunteers (9 males, 1 female, all Asian) and 7 triggers.
+It consists of 3,344 clean images, 9,790 trigger images of 7 physical triggers, and 8,104 additional images to evaluate accidental backdoor activations (see Section 8).
+Our animal dataset is created by combining a public Kaggle dataset [1], containing 45 classes of mammals, with a new cat class containing 500 cat images.
+For this dataset, we collected 1,500 cat images with 3 physical triggers (tennis ball, phone, book) to study CLPBA.
+The images are captured in various indoor and outdoor settings, under varying weather conditions, and with diverse shooting angles and distances.
+Details of our dataset can be found in the Appendix B.
+Table 3: ASR (%) of poisoning algorithms on selected source-target pairs, averaged over 10 runs with standard deviation in brackets.
+Scenario Poison ratio $\alpha$ Naive MP FM GM
+$\epsilon=8/225$ $\epsilon=16/255$ $\epsilon=8/225$ $\epsilon=16/255$ $\epsilon=8/225$ $\epsilon=16/255$
+(a) Facial Recognition (sunglasses) trigger
+finetuning 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.6 (0.5) 5.8 (3.2) 22.7 (10.9) 95.6 (1.8)
+10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 3.4 (0.5) 30.4 (1.5) 51.4 (17.8) 97.8 (0.2)
+transfer 5% 0.0 (0.0) 2.2 (0.5) 1.8 (1.3) 81.8 (5.6) 77.8 (9.6) 91.2 (7.6) 92.5 (8.5)
+10% 0.0 (0.0) 2.4 (1.8) 3.6 (1.2) 79.3 (2.3) 64.5 (25.7) 93.6 (6.9) 97.2 (2.8)
+(b) Animal Classification (tennis) trigger
+fine-tuning 5% N/A 0.0 (0.0) 0.0 (0.0) 0.3 (0.5) 0.2 (0.3) 8.2 (4.4) 9.6 (4.1)
+10% N/A 0.0 (0.0) 0.0 (0.0) 1.0 (0.8) 0.2 (0.4) 20.5 (8.9) 17.8 (7.7)
+transfer 5% N/A 1.1 (0.7) 3.2 (2.5) 34.5 (1.9) 29.7 (3.0) 62.0 (2.1) 75.6 (5.5)
+10% N/A 1.5 (1.2) 2.8 (1.2) 33.6 (2.8) 34.3 (2.3) 92.8 (1.2) 88.4 (3.7)
+5 Experiments
+5.1 Experiment Settings
+We split our datasets with a ratio of 8:2 for train and test data.
+We use 50% of source-class trigger images to craft poisoned samples, and the remaining 50% is used to evaluate ASR.
+Note that the prediction accuracy (ACC) on clean samples consistently exceeds 99% for facial recognition and 93% for animal classification across different attacks, so it is not reported in the paper.
+We choose ResNet50 [18], pre-trained on the VGGFace2 dataset [4], to be the victim model for Facial Classification, while ResNet18 pre-trained on Imagenet [40] is used For Animal classification.
+We evaluate the attacks under transfer learning and fine-tuning settings, two common settings in clean-label data poisoning [41, 46].
+For transfer learning, the feature extractor is frozen and only the last linear layer is trained, while for fine-tuning, we set a low learning rate of 0.001 to train the victim model on the poisoned dataset.
+To enable a fair comparison across different poisoning algorithms, we standardize the poisoning hyperparameters, including the number of iterations, optimizer, batch size, learning rate, and other relevant parameters.
+For the Facial Classification task, we select the poison-pair $9-5$ (as class labels in our dataset) for our experiment.
+This poison pair is manually selected to avoid dataset bias, as explained in Section 3.
+For the Animal Classification task, we select the poison-pair cat-fox.
+5.2 Baseline evaluation
+From the experiment results demonstrated in Table 3, we have the following observations:
+  • GM is the most effective poisoning algorithm.
+GM is especially effective for the fine-tuning scenario, which is a harder scenario compared to transfer learning since the weights of all layers of the victim model are updated.
+Matching the gradient update of $L_{adv}$ allows the GM perturbations to utilize the backward information of all layers, instead of only the last linear layer as in FM.
+MP is not effective because the poisoned samples overfit the updated parameters of the unrolled surrogate model in the inner loop, which greatly differ from the victim model.
+  • Impact of $\epsilon$ and $\alpha$.
+While increasing either $\epsilon$ or $\alpha$ would improve the performance of attacks, we observe that $\epsilon$ plays a more important role than $\alpha$, especially for fine-tuning scenario.
+A high value of $\epsilon$ is critical for the poisoning objective to converge, while a high value of $\alpha$ does not guarantee that every poisoned sample is effective.
+  • Effectiveness of CLPBA varies across datasets.
+As shown in Table 3, CLPBA is significantly less effective in Animal Classification than in Facial Recognition.
+This reduced performance is due to the tennis trigger being as dominant as the sunglasses trigger while also appearing in varying locations across the trigger samples, making it harder for the model to learn.
+5.3 Analysis of CLPBA
+Two unique attributes of CLPBA are that it is hidden-trigger and source-specific.
+The first attribute means the trigger is not present in the poisoned samples, while the second ensures only source-class samples can cause the model to misclassify with the trigger.
+CLPBA achieves these attributes by embedding features of the source-class trigger distribution, $\widetilde{\mathcal{D}}_{s}$, into the perturbations in the poisoned target samples.
+Specifically, the embedded features consist of adversarial features and trigger features.
+We define adversarial features as the non-robust features [24] that shift the representations of target instances toward the source-class distribution in the feature space, and trigger features as the high-level features of the physical trigger extracted by the neural network.
+The existence of adversarial features ensures the source-specific attribute, while reconstructed trigger features allow the model to indirectly learn the physical trigger as a feature of the target class.
+Figure 6: Experiment to show the existence of adversarial features. We set a high $\epsilon$ of 32, $\alpha=2$, poison pair is $9-5$.
+Existence of adversarial features.
+To justify the existence of adversarial features, we take inspiration from [24] to test whether the model trained on poisoned target samples could successfully predict source-class samples.
+We design the experiment as illustrated in Figure 6.
+Then, we filter out all source-class samples before relabelling all GM-poisoned samples to the source class.
+We then train the victim model from scratch on this modified dataset before evaluating the recall (true positive rate) of predicting source-class samples that have been filtered previously.
+To avoid randomness, we set a prediction threshold of 85% after the softmax function.
+Notably, the model trained on this modified dataset has a prediction accuracy of 75% on source-class samples.
+This shows that the perturbations on poisoned target instances contain adversarial features representative of the source-class distribution.
+To make sure that the trained model is not biased towards predicting out-of-distribution samples as the source class, we also evaluate the false positive rate of classifying 200 random face images from the CelebA dataset [32] as the source class, and the accuracy is 0%.
+The setup of this experiment can be found in Appendix D.
+Existence of trigger features.
+By scaling the perturbations on GM-poisoned samples, we observe that the perturbations appear to reconstruct the shape of the physical trigger.
+Interestingly, setting a high value of $\epsilon$ even reconstructs the samples in $\widetilde{\mathcal{D}}_{s}$, as can be observed in Figure 8.
+(a) Clean model
+(b) Poisoned model
+Figure 7: Feature space visualization.
+Feature space analysis.
+In Figure 7, we project source-class samples (green), target-class samples (blue), trigger source samples (yellow), and poisoned target samples (red) in 2-dimension feature space of both clean and poisoned models where the first axis is $\mu_{s}-\mu_{t}$ and the second axis is $\theta_{s}-\theta_{t}$, the normal vector to the decision boundary between the source and target class.
+In both feature spaces, we can see that some poisoned target instances are close to the samples of $\widetilde{\mathcal{D}}_{s}$.
+For the clean model, $\widetilde{\mathcal{D}}_{s}$ overlaps with ${\mathcal{D}}_{s}$ in the feature space.
+For the poisoned model, however, $\widetilde{\mathcal{D}}_{s}$ is shifted towards $\widetilde{\mathcal{D}}_{t}$ in the feature space.
+From this observation, we can say that since poisoned target instances contain representative features of $\widetilde{\mathcal{D}}_{s}$, especially the trigger feature which is not a part of ${\mathcal{D}}_{s}$, training the model on these poisoned instances cause $\widetilde{\mathcal{D}}_{s}$ to move closer to the classification region of the target class in the feature space.
+Figure 8: From left to right: tennis, fake-beard, and sunglasses triggers. First row: a sample in $\widetilde{D}_{s}$. Second row: GM perturbation.
+For filtering defenses, we evaluate Spectral Signature (SS) [49], Activation Clustering (AC) [3], SPECTRE [17], and SCAN [48].
+Table 6 shows that only SS is effective against CLPBA.
+Although we can see that AC and SPECTRE have high elimination rates (62.16% and 72.97%), the attack still achieves 99.7 % ASR.
+This shows that the GM poisoning objective (Equation 9) focuses on a small subset of effective poison instances to match the gradient of adversarial loss and distill the features from the trigger distribution.
+This is supported in Figure 7 where only a subset of poisoned samples overlap with the trigger distribution.
+Detecting such effective instances is generally difficult since they correlate strongly with each other and they transform the model’s decision boundary of the target class.
+Table 6 also shows that firewall defenses are not effective for CLPBAs because they either miss trigger samples (low TPR) or incorrectly filter out too many benign samples (high FPR).
+This behavior is expected since these defenses are built on the memorization property of backdoor attacks, which differs from the generalization property of CLPBA.
+6 Improving the Effectiveness of CLPBA
+6.1 Pixel regularization
+Empirically, higher $\epsilon$ lead to a higher ASR since the poisoning objective can converge to a smaller value.
+However, the perceptual quality of poisoned samples is compromised since the poisoned images have high-frequency artifacts that are clearly visible (Figure 8).
+The drawbacks of the $\ell_{\infty}$ regularization on perturbations are that it is non-differentiable, and the projection on $\ell_{\infty}$-norm ball is only executed at the end of an optimization step, making backdoor and poison stealthiness two separate tasks.
+As a consequence, the perceptual quality of poisoned images is compromised because (1) the pixel values of perturbation are often pushed to the boundary of the norm ball, (2) the noise can appear in unnecessary regions of the image, and (3) the perturbations contain high-frequency artifacts.
+To improve the visual quality and convergence of the poisoning objective, we propose to replace the current $\ell_{\infty}$ regularization with a Pixel regularization containing a soft pixel-wise $\ell_{\infty}$ loss and discretized Total Variation Upwind (TVU) [6]:
+$\mathcal{R}_{pix}(\delta)$ $=||\delta^{+}-M\odot\epsilon||_{1}+\alpha TVU(\delta;\beta=1)$ (11)
+$TVU(\delta;\beta)$ $=\sum_{p}^{\text{all pixels}}\biggl{(}\sum_{q\in\mathcal{N}(p)}\big{(}\delta^{(p)}-\delta^{(q)}\big{)}^{2}\biggr{)}^{\frac{\beta}{2}}$ (12)
+where $\delta^{+}$ takes the absolute value of every element in $\delta$, $M$ is a mask of pixel-wise embedding ratio: $M_{i,j,k}=\frac{x_{i,j,k}}{\max_{k}{x_{i,j,k}}}$, assuming $x_{i,j,k}\in[0,255]$.
+$M$ encourages the perturbation to align with the underlying image content.
+In Equation 12, $\mathcal{N}(p)$ are neighbouring pixels of pixel $p$.
+The weighting factor $\alpha$ is set to 5.
+Both terms in Equation 11 improve the smoothness of perturbations, with $TVU(\delta)$ minimizing the difference between adjacent pixels.
+Side-by-side comparison between standard $\ell_{\infty}$ regularization and our pixel regularization is given in Figure 9 and Figure 10.
+Our pixel regularization leads to a high ASR even for small values of $\epsilon=[4,8]$, while the ASR for standard $\ell_{\infty}$ regularization is below 20%.
+At the same time, our pixel regularization also has higher PSNR values for the poisoned images at every $\epsilon$, indicating better perceptual stealthiness.
+Table 4 demonstrates the improved effectiveness of pixel regularization across different poison pairs.
+(a) Standard $\ell_{\infty}$ regularization
+(b) Pixel Regularization
+Figure 9: Visualization between standard $\ell_{\infty}$-regularization and our Pixel Regularization. Poisoned samples are crafted with GM attack ($\alpha=0.1$, $\epsilon=16/255$)
+Table 4: ASR (%) of GM attack + tennis with two regularizers.
+Regularizer Fox Kangaroo Mongoose Sea-lion Tapir
+$\ell_{\infty}$ regu 9.6 (4.1) 8.4 (0.8) 5.3 (0.4) 13.4 (2.9) 7.8 (5.0)
+pix. regu 30.9 (10.4) 12.5 (1.6) 4.4 (1.0) 24.2 (2.3) 26.6 (5.2)
+(a) Standard $\ell_{\infty}$ regularization
+(b) Pixel Regularization
+Figure 10: Comparison between standard $\ell_{\infty}$-regularization and our Pixel Regularization on different values of $\epsilon$.
+Higher PSNR indicates better perceptual quality of poisoned images.
+6.2 Ensembling and Retraining
+Following [46], we also apply ensembling and retraining for CLPBA methods in the fine-tuning setting of the Animal Classification task.
+In our experiment, we apply an ensemble of 3 ResNet18 models; each of these models will retrain on the poisoned dataset after every 200 attack iterations.
+The attacker optimizes poisoned samples with GM for 1000 iterations.
+Table 5 shows that applying these techniques significantly improves the attack success rate.
+Table 5: ASR (%) of GM ($\alpha=0.1,\epsilon=16/255$) on the Animal Classification task with tennis trigger.
+Method ASR (%)
+Baseline 17.8 (7.7)
+Retraining 25.5 (4.2)
+Ensemble+Retraining 29.7 (3.1)
+7 Defending against CLPBA
+We evaluate the performance of CLPBA against representative filtering defenses (filter poisoned samples during training), and firewall defenses (filter trigger samples during inference).
+For the filtering defenses, we evaluate Elimination Rate (ER), the percentage of poisoned samples that are correctly filtered, and Sacrifice Rate (SR), the percentage of clean samples that are incorrectly filtered.
+For firewall defenses, we evaluate the True Positive Rate (TPR), the percentage of trigger source-class samples that are correctly filtered, and the False Positive Rate (FPR), the percentage of non-trigger samples that are incorrectly filtered.
+For details on these defenses, we refer readers to Appendix F.
+Table 6: Defense evaluation on GM attack ($\alpha=0.1$, $\epsilon=16/226$, poison pair $7-0$). ASR is 99.7% without defenses.
+Filtering Defenses
+Metrics AC[3] SS[49] SPECTRE[17] SCAN[48]
+ER (%) 62.16 97.30 72.97 0
+SR (%) 15.77 19.77 19.50 0
+ASR (%) 99.7 0.0 99.7 99.7
+Firewall Defenses
+Metrics STRIP[14] CD[20] IBD-PSC[19] Frequency[56]
+TPR (%) 0.0 79.6 0.0 0.0
+FPR (%) 6.7 65.3 16.7 0.9
+8 Accidental Backdoor Activations
+For CLPBA to be stealthy, we argue that it needs to be source-specific and trigger-specific.
+That is, (1) only the source class can use the trigger to manipulate the victim model’s prediction at inference time; (2) the trigger should be discriminative so that the victim model does not incorrectly recognize similar-looking objects as triggers.
+We introduce two new metrics, False Positive Rate, and Suspicion Rate, to test accidental backdoor activations:
+  • Suspicion rate (SR): The measurement of whether the victim model incorrectly misclassifies trigger samples from non-source classes as belonging to the target class.
+  • False positive rate (FPR): The measurement of whether samples from the source class can fool the victim model with false triggers (i.e., objects that resemble the attacker’s physical trigger).
+Poison pair Trigger Loss type ASR ACC FPR SR
+9-5 Fake beard Original 98.7 99.9 50.8 12.5
+With repel 98.7 99.9 49.7 3.9
+7-3 Sunglasses Original 97.5 99.3 0.7 13.9
+With repel 97.5 99.4 0.9 2.5
+Table 7: FPR and SR for GM attack. The set-up of our dataset to evaluate SR and FPR is given in Appendix B.
+Between the two metrics, SR is more important because it directly impacts attack stealthiness if a benign client is misclassified as the target class with the trigger or similar-looking items.
+The issue of false triggers, measured by FPR, is challenging to solve as the attacker cannot anticipate all possible false triggers in the real world.
+However, the attacker can partially mitigate the issue of false source classes, as measured by SR, by adding repel terms to the adversarial loss $L_{adv}$ in Equation 3:
+$L_{adv}=N\cdot\overbrace{\sum_{x\in\widetilde{D}_{s}}\mathcal{L}\bigl{(}F_{\theta}(x),y_{t}\bigr{)}}^{\text{adversarial loss}}$
+[…]
+Here, $N$ is the total number of classes, and $\widetilde{D}_{i}$ is the dataset of trigger instances in class $i$.
+The repel terms ensure trigger instances from non-source classes are repelled from the target class classification, enhancing trigger specificity to the source class.
+As shown in Table 7, incorporating the repel term in the GM objective (Equation 9) significantly reduces SR while maintaining high ASR.
+One limitation of using the repel terms is that the attacker must possess trigger samples from non-source classes.
+To circumvent this, the attacker can patch clean images with digital triggers, mapping them to appropriate facial areas.
+9 Conclusion
+Through extensive experiments, our research underscores the vulnerability of DNNs to clean-label physical backdoor attacks.
+We then show that CLPBA works by embedding adversarial features and trigger features in the poisoned instances.
+Additionally, we propose a novel pixel regularization to improve the effectiveness and stealthiness of poisoned samples.
+A significant limitation identified is accidental backdoor activations of the backdoored model due to false triggers or non-source classes, which could compromise attack stealthiness.
+To address this, we propose incorporating repel terms into the adversarial objective to enhance trigger specificity to the source class.
+Supplementary Material
+Appendix A Dataset
+A.1 Ethics & Data Protection
 IRB approval.
@@
 Additionally, all images in the paper are partially obscured to ensure that the identities of our volunteers are not exposed.
-B.2 Dataset Collection
+A.2 Dataset Collection
 Due to the lack of publicly available datasets to study physical backdoor attacks, we collect a facial recognition dataset following the design consisting of approximately 21,238 images of 10 volunteers (9 males, 1 female, all Asian) and 7 triggers.
@@
 With these criteria, we select 7 triggers with evaluations given in Table 8.
-B.3 Dataset Preprocessing
+A.3 Dataset Preprocessing
 After collecting the images, we utilize a pre-trained MTCNN [58] (Multi-task Cascaded Convolutional Networks) model to detect and crop the face area.
@@
 Figure 11: Visualization of the face detection process. Left: Original image. Right: Processed image with face area cropped and resized to $224\times 224$ pixels.
-B.4 Dataset Structure
+A.4 Dataset Structure
 The train and test data are trigger-free images (the participants do not wear any trigger items) of 10 classes corresponding to 10 volunteers. The train and test data are split in a $80:20$ ratio.
 False triggers.
-To evaluate the sensitivity of poisoned models to false triggers, as measured by false positive rate (FPR) (see Appendix C.4), we select three additional false triggers for each of the main triggers.
+To evaluate the sensitivity of poisoned models to false triggers, as measured by false positive rate (FPR) (see Appendix B.4), we select three additional false triggers for each of the main triggers.
 Because it is hard to find a representative set of false triggers for each of the main triggers, we select the false triggers by changing the sizes, colors, and shapes of the objects (e.g, we change the shape of earrings and hats, change the colors of the headband, change the size of sticker).
@@
 Therefore, a higher FPR of a trigger does not necessarily mean that that trigger is less specific and stealthy than other triggers.
-B.5 Animal Classification Dataset
+A.5 Animal Classification Dataset
 Besides facial recognition, we also conduct an extended evaluation of CLPBA on animal classification.
@@
 Visualization of trigger images for this dataset is given in Figure 19  [^1] https://www.kaggle.com/datasets/asaniczka/mammals-image-classification-dataset-45-animals.
-Appendix C Experiment Settings
-C.1 Computational Resources.
+Appendix B Experiment Settings
+B.1 Computational Resources.
 All the experiments are conducted on the two servers.
-The first one with 7 RTX 3090 Ti 24 GB GPUS, and the second one contains 6 RTX A5000 24 GB GPUS.
+The first one with 7 RTX 3090 Ti 24 GB GPUS, and the second one has 6 RTX A5000 24 GB GPUS.
 The code is implemented in Pytorch.
@@
 We use Pytorch as the framework for our model training.
-C.2 Training Settings.
+B.2 Training Settings.
 We evaluate the attacks on four Convolutional Neural Networks for facial recognition: ResNet50 [18], VGG16 [45], MobileNetV4 [42], Inception-ResNet [47].
@@
 Inner SGD steps (MP) $2$
-C.3 Attack Settings.
+B.3 Attack Settings.
 Our experiments follow the gray-box scenario in which the attacker model has the same architecture and is trained on the same dataset as the victim model but has different weight initialization, and the attacker cannot access and modify the weights and training procedure of the victim model.
@@
 In our experiment, the attacker selects the poison samples randomly.
-C.4 Evaluation Metrics.
+B.4 Evaluation Metrics.
 To evaluate the performance of CLPBA, we adopt two standard metrics for backdoor attacks:
@@
   • Accuracy (ACC): The model’s prediction accuracy on clean, ordinary test-time data when the model is poisoned.
-Appendix D Additional Experiments
-D.1 Blended Label-Consistent Attack
+Appendix C Additional Experiments
+C.1 Blended Label-Consistent Attack
 For completeness, in our evaluation, we also evaluate Blended Label-Consistent attack [50, 8] which works by perturbing the poisoned samples with adversarial noise to make the salient features of the samples harder to learn before injecting the digital trigger, forcing the model’s classification to depend on the trigger.
@@
 Then, using facial landmarks, we map the trigger to the appropriate facial area with perspective transformation.
-D.2 Experiments on Different Models
+C.2 Experiments on Different Models
 Table 11: CLPBA ($\alpha=0.1$, $\epsilon=16/226$) on VGG16
@@
 We further evaluate the performance of CLPBA on 6 random poison pairs, on VGG16 and Inception-ResNet model architecture for the Facial Recognition task in Table 11 and Table 12.
-D.3 Scaling the Dataset Size
+C.3 Scaling the Dataset Size
 To test the effect of dataset size and the number of classes on the performance of CLPBA, we scale our dataset by merging with two benchmark datasets, PubFig [27] and CelebA [32], with the following configurations:
@@
 Interestingly, as shown in Table 14 and Table 13, scaling dataset size only minimally impacts the attack performance.
-D.4 Evaluating Digital Trigger for CLPBA
+C.4 Evaluating Digital Trigger for CLPBA
 In CLPBA, we assume that the attackers hold trigger samples from the source class to perform poisoning.
@@
 GM-physical 100.0 (0.0) 99.4 (1.0) 98.3 (0.8) 97.8 (0.2)
-D.5 Evaluating transferability of CLPBA
+C.5 Evaluating transferability of CLPBA
 We evaluate the black-box settings of CLPBA for two poison pairs $7-1$ and $3-6$ in Table 16 and Table 17 Generally, we find that CLPBA under black-box settings have a worse performance, which means that the semantic features extracted by different model architectures may differ greatly.
@@
 VGG16 72.9 49.3 71.5 22.2 100.0 100.0
-Appendix E Additional Visualization
-E.1 Poisoned samples of different poisoning algorithms
+Appendix D Additional Visualization
 Figure 13: Additional visualization of poisoned samples
+Visualization of poisoned samples.
 As shown in Figure 13, the perturbations of GM appear to reconstruct the sunglasses trigger.
@@
 GM is more successful than FM because it successfully reconstructs the trigger in the perturbation in most of the samples.
-E.2 Additional imperceptibility comparison of regularizations
 Figure 14: Additional visualizations of GM attack with $\ell_{\infty}$ regularization and Pixel Regularization.
 Left image generated by standard $\ell_{\infty}$ regularization, Right Image generated by Pixel Regularization.
-Appendix F Defenses
+Figure 15: Visualization of poisoned squirrel images with the standard $\ell_{\infty}$ regularization
+Figure 16: Visualization of poisoned squirrel images with pixel regularization
+Appendix E Defenses
 To defend against backdoor attacks in DNNs, a multitude of strategies have been developed, each addressing different vulnerabilities.
@@
 Backdoor samples consistently yield unusually small CPs, making them easy to identify.
-F.1 Evaluation on filtering defenses.
+E.1 Evaluation on filtering defenses.
 Backdoor poisoning attacks have been observed to leave discernible traces in the feature space of poisoned models.
@@
 Figure 17: 2D-PCA view of target-class representation produced by the poisoned model
-F.2 Evaluation on firewall defenses.
+E.2 Evaluation on firewall defenses.
 As demonstrated in Table 6 in the main paper, input-level detection methods are not effective for CLPBAs because they either miss trigger samples or incorrectly filter out too many benign samples.
@@
 Therefore, the assumption of a consistent backdoor correlation for STRIP and IBD-PSC may not hold true.
-Appendix G Limitations and future works
+Appendix F Limitations and future works
 High variance.
@@
 A natural direction from this work would be to study clean-label physical backdoor attacks for different tasks, such as Object Detection and Image Segmentation.
-Appendix H Societal impact
+Appendix G Societal impact
 The challenges highlighted by the Clean-label Physical Backdoor Attack (CLPBA) can push the AI community towards developing more trustworthy AI models.